A Bibliography of Project Gutenberg

Steve Thomas Barr Smith Library, Room 210

The University of Adelaide Adelaide, South Australia 5005

Australia

Tel: +61 8 303 5190 FAX: +61 8 303 4369

E-mail: sthomas@library.adelaide.edu.au
WWW URL 1: http://www.library.adelaide.edu.au/ual/staff/sthomas.html
WWW URL 2: http://www.library.adelaide.edu.au/~sthomas/

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.15

Title word cross-reference

#1 [188].

$30 [143].

1.5M [784]. 104th [257]. 10a [1203]. 1945-1995 [259]. 19th [1651]. 1st [379, 1025, 105, 1027].

30th [894]. 32nd [70].

465th [256].

7-bit [960, 961, 962, 963]. 7th [1611].

8-bit [975].

A. [518]. A.L. [239]. Abbey [122]. ABC [1285]. Abnormal [1180]. Abounding [631]. Abraham [872, 1757]. Abroad [92, 120]. Absentee [1425]. Abyss [532, 1637]. Abyssinia [629, 1388]. Accession [1420]. Accoramboni [774]. account [1431]. Accused [469]. Acquaintance [851]. Acres [349]. Across [1190, 591]. Act [233]. Acton [743]. Actor [1651]. Adam [488, 291, 379]. Adams [946]. Address [232, 1, 5, 6, 10, 11, 105]. Adieu [1504]. Admiral [1440]. Ado [1072, 1469, 1470, 1727]. Adonis [1008]. Adventure [293, 1117]. Adventures [1767, 471, 481, 1592, 1609, 386, 875, 684, 1326, 1172, 930, 1148, 77, 75]. Adversary [1109]. Aeneid [210, 209]. Aerial [1235]. Aeronautics [845]. Aeropagitica [585]. Aeroplane [789]. Aeroplanes [765, 1397]. Affair [344, 834]. Affairs [424]. Africa [1002, 1404]. African [1393]. After [1574, 1212]. against [998]. Age [272, 520]. agency [169]. Agent [940]. Ages [1563]. Agesilaus [1123]. Agnes [739]. Agony [1756]. Ahead [541]. Air [918, 1238, 832, 749, 752]. Airships [734]. al [1166, 1239, 1581, 1263, 437]. Aladdin [58]. Alamo [268]. Alcibiades [1625, 1626]. Alcove [1793]. Alexander [95]. Alexandria [1229]. Alice [13, 894, 946, 115]. Aline [308]. Alkahest [1405]. Allan [688]. Almayer [697]. Altar [619]. Altruist [1295]. Amantis [247]. Amateur [1764, 683, 67]. Amazing [1539]. Ambassadors [413]. America [586, 787, 1807]. American [1175, 652, 168, 178, 943, 1806]. Among [1236, 1317]. Amours [1345]. Amulet [808]. Amy [476]. Anabasis [1124]. Anatomy [1176]. Ancien [1289]. Ancient [626, 152, 818]. Ancients [616]. Andes [525]. Andronicus [1060, 1457, 1715]. Angler [660]. Anglo [595, 634]. Anglo-Saxon [595, 634]. Animals [1181]. Ann [503]. Anna [1351]. Annals [1264]. Anne [48, 46, 52, 523]. Annexes [90]. Anniversary [256]. Anomalies [719]. Anthem [1204, 1203]. anthm10 [1203]. Anthology [188, 1234, 1153]. Antiquities [1357]. Antonia [224]. Antony [1084, 1484, 1738]. Apes [79]. Apocalypse [1436]. Apocrypha [125, 1682, 1714, 1497, 1492]. Apollo [117]. Apology [1604, 1125]. Appreciations [387]. Apprentices [859]. April [1201]. Arabian [129, 810]. Arcadia [1431]. Arcis [1812]. Arcturus [1283]. Arden [1312]. Ardmore [794]. Argonautica [801]. Arizona [728, 1012, 725]. Arms [1152]. Arriere [1797]. Arrow [819]. Art [36, 169, 508, 473, 133, 1403]. Arthur [745].

Arthurian [802]. Articles [100, 254]. Ascents [865]. Aspern [193]. Asse [1614]. Astor [393]. Astoria [1325]. Atheist [1174]. Athenians [1132]. Athens [1086, 1486, 1740]. Atomic [260, 662]. Attic [222]. Aucassin [1527]. Augsburg [256, 584]. Aunt [1615]. Australian [1153]. Author [328, 301, 722, 754]. Authors [1443]. Autobiographies [1651]. Autobiography [149, 1269, 1227, 670, 457, 1324]. Autocrat [723]. aux [1557]. Avonlea [48, 1308]. Awakening [161]. Ayrshire [1336].

B. [1685, 693]. Baartock [1041]. Bab [729, 897, 899, 347]. Babbit [1110]. Babies [981]. Babington [1685]. Baby [558, 831]. Bachelor [1146]. Back [836, 207, 116]. Backward [601]. Badge [74, 444]. Baldwin [1102]. Ball [1259, 338]. Ballad [1665, 282, 1020]. Ballads [626, 729, 897, 899, 767, 1017, 435, 993, 240, 958, 394]. Ballantrae [835]. Balloon [865]. Baltimore [1338]. Ban [1797]. Banner [704]. Barbarian [1705]. Barber [809]. Bargain [621]. Barlaam [721]. Barleycorn [299]. Barnaby [883]. Barnum [1525]. Baron [526]. Basket [470]. Bassington [534]. Battle [653, 600]. Baxter [1650]. BC [950]. bClassic [1773]. Be [890]. Beacon [1449]. Beagle [910]. Bearded [1705]. Beast [1047, 273]. Beasts [86, 250]. Beautiful [1382]. Beauty [252, 273]. Bede [488]. Before [291]. Begins [1354]. Behavior [1398]. Behind [453]. Being [815, 1781, 1431]. Beliefs [1224]. Believe [1765]. Bell [743]. Bellew [1545]. Belt [127]. Ben [149]. Bente [1668]. Benton [713]. Beowulf [947]. Bequest [143]. Bergerac [1207, 1208, 1209]. Bergson [1301]. Berne [234]. Best [1313]. Betty [1695, 1214]. Between [759]. Beyond [337]. Bibber [315]. Bible [12, 32, 125, 796, 1530, 1531, 396, 1541, 1558]. Biblical [595]. Bibliomaniac [424]. Bickerstaff [1044]. Bickerstaff-Partridge [1044]. Bicycling [1217]. Big [919, 1710, 119, 549]. Bill [4, 1271]. Billie [343, 342]. Billy [1710]. Biloquist [813]. Biog [518]. Biographical [743]. Biography [139, 1178, 385]. Bird [1330, 698]. Birds [1542]. Bishop [462, 1222]. bit [960, 961, 962, 963, 975]. Black [1536, 68, 389, 931, 1685, 1705, 1412, 252, 819]. Black-Bearded [1705]. Blades [983]. Blanc [359]. Bleak [986]. Blind [1394]. Blix [382]. Blood [1777]. Blue [1552, 206, 484, 1354, 374]. Blunders [352]. Boat [915, 289]. Bobbsey [709, 691]. Bohemia [1754]. Bohemian [958]. Bomb [260, 258]. Bombings [662]. Bonaparte [1300]. Bondage [184, 332]. Bonneville [1326]. Book [19, 232, 796, 1258, 427, 218, 484, 519, 618, 617, 948, 741, 379, 939, 1119, 1120, 1580, 1511]. Books [125, 200, 1256, 1563, 406, 600]. Bookshop [173]. Boone [875]. Booth [405]. Border [1193, 1751]. Born [992]. Boss [1342]. Both [12, 1530]. Bourgeoisie [1589]. Box [1534]. Boy [277, 443, 507, 501, 1757, 850]. Boyhood [756, 318]. Boys [541]. Boz [853]. Breakers [380]. Breakfast [723]. Breaking [1550]. Breitmann [435]. Breteche [1656]. Brethren [806]. Brick [1581]. Bride [452]. Bridge [356, 95]. Brief [231]. Brinker [736]. Brinker/Silver [736]. Britain [848, 1746]. British [734]. Brontë [1649, 1769]. Brother [761, 1676, 720].

Brothers [1334]. Brown [186, 205, 1432]. Browning [241, 632, 633]. Brumaire [1300]. Bruno [597]. Bucky [1751]. Bucolics [211, 212]. Bulba [1151]. Bullet [1778]. Bungay [695]. Bunker [1561]. Bunner [292]. Bunyan [1000]. Bureau [31]. Bureaucracy [1297]. Burial [313]. Buried [1707]. Burning [718]. Burnt [574]. Bush [1565]. Business [1755, 701, 1273]. Buttered [333]. Bygone [1224]. Byways [360].

c [1312, 638, 712]. Cabin [1158, 185]. Cachalot [1310]. Cactus [1175]. Cadignan [1298]. Caesar [200, 1074, 1472, 1729]. Cage [1098]. Cairo [1805]. Calamity [1509]. Calhoun [712]. Call [197]. Called [1440]. Camera [1237]. Camilias [1557]. Camille [1557]. Camp [1522]. Campaign [1522]. Canaan [464]. Cane [1683]. Cannon [1315]. Canoe [1036]. Canongate [1770]. Cape [857]. Captain [275, 1189, 1326, 1007]. Captives [850]. Captivity [822]. Car [339]. Cardinal [512]. Carmina [219]. Carol [47, 698]. Carolina [1780]. Carpet [622, 807]. Carriage [462]. Carrie [215]. Carwin [813]. Cary [968, 969, 970, 971]. Case [1778, 1774, 1777, 1776, 1775]. Cash [277]. Cast [380]. Casterbridge [144]. Castle [1376, 673]. Castro [769]. Cat [1629, 1662]. Catalan [659, 784]. Catechism [1668, 1619]. Cathechism [398]. Cathedral [1501]. Catherine [1796]. Catriona [566]. Cavalier [993]. Cavalry [1126]. Cave [230]. cby [1388]. Cedars [453]. Celestina [1567]. Cenci [773]. Centre [936]. Centuries [1435]. Century [1651]. Cérè [1249]. Certain [269, 1358]. Cevennes [514]. Chamber [1617]. Chance [1428, 1217]. Chancellor [1600, 1647]. Charge [274]. Charlotte [1649, 1769, 172]. Charmed [1763]. Charmides [1529, 994]. Chartreuse [768]. Cheechako [240]. Chemical [1225]. Chessman [1107]. Chief [631]. Chieftains [317]. Chignecto [1305]. Child [676, 1657, 954, 216, 137]. Children [455, 454, 881, 750, 1815, 294, 874]. Chimes [630]. China [1624, 502]. Chinese [371, 501, 569]. Ching [198]. Chita [694]. Christ [1601]. Christendom [486]. Christian [1510]. Christmas [47, 1419, 1792, 698]. Christopher [1440, 1444, 216]. Chronicle [634]. Chronicles [1308, 1770, 1329]. Chuzzlewit [934]. CIA [16, 27, 88, 49, 181, 548, 1610]. Cinderella [1548]. Circular [415]. Circus [830]. Cities [99, 555]. City [755, 496, 337, 1396, 860, 1192, 1309]. Civil [579, 72]. Civilization [1688, 1638]. Class [804]. Classes [1141]. Claus [498]. Clay [711]. Cleopatra [1084, 1484, 1738]. Cleves [448]. Clock [565]. Clocks [826]. Cloister [1320]. Clotelle [223]. Club [1645]. Co. [434]. Codex [595]. Col [875]. Collected [100, 1692]. Collection [1357, 329, 1017, 1025, 559, 1563, 1027]. Colored [223]. Colour [1159]. Coloured [242]. Columbus [1440, 1444]. Column [1756]. Comedians [1196]. Comedy [971, 967, 1173, 1058, 1454, 1713]. Comet [1307]. Comic [1173]. Commedia [960, 961, 962, 963, 972, 973, 974, 975]. Commentaries [200]. Commentary [1499, 1439]. Commission [1362]. Common [148]. Communist [62]. Compact [9]. Company [869]. Comparison [1203]. Complaint [1091, 1493]. Compleat [660]. Complete [780, 1319, 668, 669, 101]. Compromise [80]. Computer [40, 706]. Computers [649]. Concentration [1519]. Conduct [901]. Confessio [247]. Confession [256, 584]. Confidence [179]. Conflict [414]. Confutatio [824]. Congo [984]. Congress [54]. Congressional [232]. Connecticut [87]. Conquer [364]. Conquest [1163, 1277, 464]. Constant [659, 784]. Constants [194, 611]. Constitution [7, 589, 590]. Consul [1708]. Continent [1190, 150, 266]. Contract [1506]. Contrast [533]. Contributions [1416]. Controversy [1197]. Convention [234]. Cook [896]. Copperfield [738]. Copy [1662]. Copy-Cat [1662]. Copyright [234, 233]. Coral [623]. Core [124, 524]. Coriolanus [1085, 1485, 1739]. Cormac [246]. Cornelius [1406]. Corner [1144]. Cornhill [1805]. Corrupted [1167]. Cost [837, 459]. Could [1706]. Count [1561, 1138, 286]. Counter [703]. Counter-Current [703]. Counties [949]. Country [1304, 1669, 1370, 691, 348]. Couples [882]. Coupon [225]. Courage [74, 444]. Course [1446]. Court [502]. Courtesan [1608]. Courtship [1501]. Cousin [1695, 1798]. Cow [1669]. Cow-Country [1669]. Coxon [1147]. Crackdown [102]. Cracksman [683]. Cranford [375]. Cratylus [1564]. Crayon [1791]. Creatures [658, 1418]. Creeds [1510]. Creek [356]. Cressida [1078, 1478, 1732]. Crewe [138]. Cricket [655]. Crime [745]. Criminal [458, 1274]. Criminals [427]. Crisis [1280]. Cristo [1138]. Crisyde [238]. Critias [1520]. Crito [1605]. Crock [1554]. Cromwell [1682]. Cross [1679, 296]. Crossing [369]. Crowd [29, 108, 426]. Cruise [1703, 1310, 693]. Crusoe [500, 540]. Crystal [1512]. Culprit [298]. Cumberland [305]. Cupid [409]. Curdie [686]. Curiosities [719]. Curiosity [677]. Current [703]. Cyberpunk [895]. Cymbeline [1087, 1488, 1741]. Cyrano [1207, 1208, 1209].

Daddy [158]. Daddy-Long-Legs [158]. Daguerrotype [168]. Daisy [190]. Dale [1172]. Damaged [1111]. Dame [1557]. Damien [262]. Damnation [134]. Dancing [1685]. Dangerous [1642]. Daniel [799, 875]. Danielis [799]. Danish [1104]. Danny [564]. Dante [960, 961, 962, 963, 968, 969, 970, 971, 964, 965, 966, 967]. Daphne [720]. Dark [1574, 1013]. Darkest [456]. Darkness [201, 505]. d’Arthur [1205]. Data [31]. Dau [1668]. Daughter [1433, 1790, 870]. David [1587, 738, 421]. Davis [387]. Daw [1704]. Dawn [441, 1551, 67]. Day [188, 1094, 1762, 34, 1199, 494]. Daylight [718]. Days [320, 251, 1242, 1432, 196, 1514, 1642, 104]. DC [730]. Dead [1035, 1037, 1652, 619, 1443]. Deal [1333]. Dealer [1145]. Dear [220]. Death [246, 1685, 8, 620, 1616]. Deb [347]. Dec [953]. Decameron [896]. Decision [343]. Declan [794]. Declaration [3, 281]. Decline [1170, 708, 861, 866]. Deep [1573]. Defense [1223]. Degree [1302]. Delivered [373]. Delusions [613, 690, 855]. Democracy [823, 787]. Denmark [1076]. Deputy [1812]. Derrick [1613]. Descarte [60]. Description [1046]. Desert [1505, 1122, 483, 1215]. Deserted [1675]. Destinies [1572]. Destiny [1594].

Detective [1773, 94]. Deutercanonical [125]. Devil [938]. Dialogues [1157]. Diamond [1236]. Diamonds [349]. Diary [1776, 724, 989, 1808]. Dick [17]. Dickens [710]. Dictionary [40, 229, 637, 638, 639, 640, 641, 642, 643, 644, 645, 646, 647, 650, 938]. Digits [716, 717, 53, 130, 605, 606, 607, 608, 609, 612]. Dim [977]. Diplomat [586]. Dirigibles [765]. Disappearance [1121]. Disasters [753]. Discourse [60]. Discoverer [1179]. Disobedience [72]. Divided [717]. Divina [960, 961, 962, 963, 972, 973, 974, 975]. Dobson [1787]. Doctor [1304, 1367, 482, 160]. Dolittle [482]. Dolly [1157]. Dolphin [1703]. Dombey [792]. Dome [242]. Domestic [1363]. Dominion [832]. Don [959]. Doni [1753]. Donkey [514]. Doolittle [1108]. Doone [811]. Door [437]. Dope [1136]. Dorado [1698]. Dore [1350]. Dorian [175]. Dorothy [401]. Dorrit [929]. Douay [1530, 1531, 1558]. Douay-Rheims [1558]. Double [1145]. Double-Dealer [1145]. Down [779, 333, 555]. Dr [1108, 751, 783, 174, 1137, 43, 44]. Dracula [326]. Drama [1379]. Dreadful [1192]. Dream [251, 1242, 1640, 338, 1410, 1067, 1464, 1722, 1689]. Dreams [827, 439, 892, 523, 1391]. Drift [1618]. Driven [509]. Drood [543]. Duchess [846]. Duchesse [450, 775]. Duke [1228]. Dummy [119]. d’Urbervilles [111]. During [1563]. Dusk [890]. Dust [1200, 911, 439, 411]. Dwarf [1412]. Dynamiter [624].

Early [276, 287]. Earnest [815]. Earth [237, 124, 524, 84]. Earthquake [1509]. East [1361]. Eastern [949]. Ebb [1553]. Ebb-Tide [1553]. Economic [1508]. Economist [1127]. Eden [1019]. Edinburgh [363]. Edison [791]. Edition [49, 256, 894, 203]. Editorials [714]. Education [823, 954]. Edward [1714]. Edwin [543]. Effects [661]. Efficiency [1273]. Egg [410]. Egoist [1633]. Egypt [1598]. Eighteenth [1300]. Electric [916, 1318]. Elephant [288]. Elixir [1169]. Ellis [743]. Email [76]. Emblems [744]. Emerald [496]. Emile [705, 1027]. Emma [159, 434]. Emotion [1181]. Emotions [1187]. Empire [708, 861, 866, 1422]. Enchanted [497]. Encyclopedia [182]. End [506, 1103, 170, 1079]. Ends [1479, 1733]. Endureth [692]. Enemies [1256]. Enemy [220]. Engineer [457]. Engineers [261]. England [1170, 626, 456, 949, 676, 993, 1420]. English [1258, 1358, 1773, 255, 1563, 1232]. Enoch [732, 1312]. Entertainments [310]. Entire [971, 967]. Eothen [263]. Episode [1408, 192]. Equal [1781]. Errand [443]. Errors [1058, 1454, 1713]. Eryxias [1630]. Essay [1173, 1508, 1155]. Essays [552, 1438, 1269, 1314, 1585, 1543, 1386, 988, 604, 746, 1511]. Esther [797, 798]. Etc [611, 430, 285, 288, 273, 715, 1186, 338, 903, 1811, 769]. Eternity [793]. Etext [32, 236]. Etexts [54]. Ethics [885, 886, 914, 937, 941]. Eugenie [1661]. European [663]. Europeans [180]. Euthydemus [1547]. Euthyphro [1590]. Eve [1433, 1587, 379]. Evergreens [828]. Everybody [671]. Evil [1575]. Evolution [1006, 1515]. Exmoor [811]. Expectations [1352]. Experience [68, 598]. Experiences [1232, 1171, 1343]. Explanation [782]. Expose [416]. Expression [1181]. Eyes [1607, 206]. Eyre [1213].

F [968, 969, 970, 971, 1287, 5]. Fables [23, 30, 164, 355, 324]. Face [1043]. Facino [1683]. Factbook [16, 27, 88, 49, 181, 548, 1610]. Factorial [194]. Faerie [306]. Fair [576]. Fairies [535]. Fairy [1546, 484, 519, 618, 617]. Faith [1050]. Fakebook [895]. Faker [1709]. Falk [474]. Fall [708, 861, 866, 1422, 898, 635]. Familiar [406]. Family [261]. Fan [762]. Fang [876]. Fanny [331]. Fantasies [407]. Fantastic [355]. Far [29, 108, 1361]. Faraday [1179]. Farm [1522, 1393, 479]. Farmer [1628]. Father [1191, 186, 205, 1388, 262, 870, 951]. Fauntleroy [460]. Faustus [751, 783]. Fay [298]. Fear [1817]. Federalist [20, 1356]. Fellow [820]. Fer [391]. Ferragus [1597]. Fever [1158]. FH [640]. Fibonacci [283]. Fiction [276, 287, 703]. Fiddler [648]. Fields [334]. Fife [1579]. Fifth [118, 157, 485]. Fighters [1317]. Figure [622]. Figures [116]. File [788]. Financier [1782]. Findelkind [1321]. Finished [1670]. Finn [77]. Fire [1317, 1509, 1113]. Fire-Tongue [1113]. Firm [1248]. Firmiani [1311]. Firs [348]. First [66, 1441, 1190, 11, 1305, 976]. Fish [877]. Fisherman [1093]. Five [750]. Flame [568]. Flanders [351, 334]. Flatland [98, 183]. Fleece [1562]. Fleeming [675]. Flirt [278]. Floor [836]. Flower [164, 1552]. Flowers [1542]. Flying [1160, 1686, 873]. Folk [389]. Folks [907]. Folly [697]. Foolish [1582]. Foot [511]. Footnote [515]. Footprints [1332]. Forbidden [860]. Forces [1624]. Forged [225]. Forgot [530]. Formula [659]. Fortune [384, 412]. Forty [1435]. Foster [476]. Found [1776]. Foundations [1524]. Four [1436, 802, 378]. Fourth [1442]. Fox [1767]. Fracasse [1189]. Frame [1748]. France [767, 1746]. Frances [316]. Francis [889, 814, 817]. Francisco [1509]. Frank [1522]. Frankenstein [85]. Freckles [112]. Free [1022]. Freedom [1015, 562, 184, 667]. French [230, 1255]. Freud [724]. Friend [563, 854]. Friendship [1166]. Frivolous [409]. Frozen [1573]. Fu [174, 1137]. Fu-Manchu [1137]. Fugitive [1797]. Fund [1147]. Furioso [592]. Further [684]. Future [1746].

G [1685]. Gables [78, 46]. Galatians [1499]. Gambara [1814]. Game [1114]. Gaming [510, 447]. Gaol [282, 1020]. Garden [114, 327, 1281, 851, 1381, 137]. Gardens [1286]. Gargantua [1154]. Gathering [761]. Gaudissart [1427, 1426]. Gen [817]. General [405, 1126]. Genius [1575]. Gentle [1747]. Gentlemen [884, 1062, 1459, 1717]. George [782, 139]. Georgia [547]. Georgics [214, 213]. Gerard [1592]. German [1281, 1424]. Gettysburg [6]. Ghost [621, 1243]. Ghosts [812]. Giant [1315]. Gibson [217]. Gift [2]. Ginx [558]. Girl [1607, 921, 1560, 428, 1679, 1551, 724, 442, 501, 126, 1808, 679]. Girls [1523]. Give [8]. Glass [1548, 14, 242]. Glasses [1149]. Glaucus [672]. Glen [1566]. Glider [918]. Glimpses [1216]. Glinda [927]. Go [541]. Goats [1239]. Goblin [685]. Gobseck [1341]. God [616, 370, 1009]. God-Idea [616]. Gods [65]. Going [1374]. Gold [483, 1554]. Golden [1614, 716, 610, 1607, 1220, 1778, 1417, 842, 442, 272, 1562, 297, 657, 678]. Golf [1446]. Good [904, 399, 340]. Goodness [665]. Goods [1111]. Goose [1808]. Gorgias [1621]. Goriot [1191]. Government [527]. Graal [722]. Grace [631]. Grafter [1747]. Grain [411]. Grand [1805]. Grande [1656, 285]. Grandet [1661]. Grant [1026]. Grass [1023, 1276]. Grassou [1184]. Gray [806, 175]. Great [1440, 1352, 317, 370, 1048, 1538, 753, 549, 408, 1202]. Green [908, 46]. Greenmantle [538]. Grenadiere [1380]. Grettir [328]. Grey [739, 1526]. Grogan [821]. Ground [1365]. Group [748]. Grub [1655]. Guide [295, 119, 41, 649]. Gull [1700]. Gulliver [581, 800]. Guns [313]. Gutenberg [182, 637, 638, 639, 640, 641, 642, 643, 644, 645, 646, 647, 650, 553, 1258]. Gypsies [544, 849].

H. [968, 969, 970, 971]. Hacker [40, 102]. Hackers [706]. Hadleyburg [1167]. Halket [1383]. Hall [935]. Hamlet [1076, 1474, 1731]. Handbook [168]. Hans [736]. Happy [868]. Hard [758, 436]. Harding [387]. Hart [713]. Harvester [330]. Hated [1407]. Haunted [171, 621, 173]. Heads [888]. Heap [309]. Hearst [714]. Heart [201, 505, 1671, 1680]. Hearth [655, 1320]. Hearts [1197]. Heaven [1007]. Heel [1118]. Heidi [1400]. Height [743]. Heights [740]. Heimskringla [575]. Helen [381]. Hell [968, 391]. Hellenica [1128]. Hells [1272]. Help [901]. Hemispheres [280]. Henri [1301]. Henry [711, 1415, 1069, 1071, 1054, 1055, 1056, 1073, 1090, 1466, 1468, 1450, 1451, 1452, 1471, 1491, 1711, 1724, 1726, 1728, 1744, 700]. Hentzau [1099]. Her [838, 870, 1281, 1229]. Heretics [451]. Heritage [1253, 1215, 1643]. Herland [34]. Hermione [748]. Hermiston [361]. Hero [1045, 879, 1806]. Herodias [1245]. Heroes [1045, 317, 654, 706, 945]. Heroine [223]. Herself [331]. Hesiod [329]. Hiawatha [21]. Hidden [1503]. Hiero [1129]. High [722]. Hilarius [761]. Hill [536]. Hills [1800]. Himself [388]. Hindustan [1422]. Hippias [1622]. Hiram [1628]. Hiroshima [662]. Historic [1523, 259]. Historical [1627, 1314, 1313]. History [722, 790, 676, 708, 861, 866, 1104, 231, 1806, 1449, 1420, 1163, 1277, 545, 1078, 169, 515, 1430, 845, 486, 1360, 1654]. Hitchhiker [41]. Holiday [781]. Holidays [340]. Hollow [42]. Holly [1346]. Holly-Tree [1346]. Holmes [109, 203, 805, 1609]. Holy [722, 1530, 1531, 376, 1558]. Home [509, 1752]. HomeBrew [284]. HomePages [284]. Homerica [329]. Honorine [1632]. Hood [803, 930]. Horatio [913]. Horse [1665, 1040]. Horseman [1252]. Horsemanship [1130]. Horsemen [1436]. Hotel [507, 171]. Hour [269]. Hours [1220]. House [1279, 1200, 711, 453, 1749, 986, 78, 1813, 523, 898, 265, 844]. HTML [281, 894]. Huckleberry [77]. Hudson [177]. Huell [852]. Hull [1279]. Human [598, 1618, 332]. Humphrey [565]. Hunted [779]. Hunter [412]. Hunting [15, 786]. Husband [856]. Hyde [1228, 43, 44]. Hydriotaphia [563]. Hymns [398].

Icelandic [328, 574]. Iconoclast [546]. Idea [616]. Ideal [856]. Idle [859, 820]. idleness [1511]. Idyll [1217]. Idylls [587]. If [1265]. Ignacio [1340]. II [1427, 1598, 1420, 1626, 1065, 1462, 1720]. III [1714, 1057, 1453, 1712]. Illustrated [183]. Illustrations [115]. Illustrious [1426, 1399]. Image [492]. Imitation [1601]. Importance [815, 825]. Improvement [979]. Inaugural [891, 1, 5, 10, 11, 105]. Increasing [1273]. Independence [3, 281]. Indian [904, 318, 317, 320, 321, 583, 319]. Industrial [385]. Industry [702]. Inferno [972, 960, 964]. Inhabitants [542]. Initials [1799]. Ink [1435]. Inland [513]. Inn [1385]. Innocence [186, 520]. Insidious [174]. Intentions [858]. Interesting [1502]. Interlude [1539]. International [192]. Internet [119, 231, 36, 41, 50]. Interpretation [1508]. Interpreted [892]. Interviewed [217]. Intro [1699]. Introduction [241]. Invention [702]. Inventions [791]. Invisible [1009]. Ioasaph [721]. Ion [1583]. Irish [1343]. Iron [847, 1251, 1118, 1507]. Iron-Mills [847]. Island [623, 497, 694, 52, 121, 310, 1221, 160]. Islands [615, 1115]. Isthmus [1305]. Italian [975, 960, 961, 962, 963]. Italy [627, 1746]. Itinerary [1102]. IV [200, 1069, 1071, 1466, 1468, 1724, 1726]. Ivanhoe [83]. Ivanoff [1701]. Ives [303].

J.P. [1271]. Jackal [392]. Jacket [1116]. Jacob [393]. Jade [371]. James [12, 731, 700, 1420, 1541, 699]. Jane [1213]. Japan [589, 590]. Jargon [40, 788]. Jasmin [809]. Jasper [693]. Jean [517]. Jefferson [689]. Jekyll [43, 44]. Jenkin [675]. Jerome [1388]. Jerry [1115, 1676]. Jerusalem [373]. Jesuits [1431]. Jew [867]. Jewels [93]. Jews [1445]. Jimmie [1172]. Jimmyjohn [1342]. Jock [1616]. Joe [507, 999]. John [712, 754, 1039, 393, 299, 1064, 1461, 1719, 1000, 338]. Johnson [1513]. Joke [839]. Jolly [1750, 1144]. Josh [936]. Journal [1180, 357, 1100]. Journals [1444]. Journey [1556, 776]. Jours [772]. Jr [188]. Jr. [700]. Juana [1389]. Jude [154]. Juliet [1066, 1463, 1721]. Julius [1074, 1472, 1729]. Junction [1371]. Jungle [107, 1047, 218, 141]. Junius [595]. Jury [1155]. Just [907, 421]. Justice [153].

Kai [1030, 1220]. Kansas [358]. Karenina [1351]. Katy [465]. Kelver [1288]. Kenilworth [1555]. Kensington [1286]. Key [417, 1773]. Khayam [228]. Kidnapped [498, 402]. Kidnapped2 [566]. Kindness [837]. King [12, 188, 1714, 345, 392, 388, 1640, 1541, 678, 1069, 1071, 1054, 1055, 1056, 1073, 1090, 1064, 1082, 1057, 1065, 1466, 1468, 1450, 1451, 1452, 1471, 1491, 1461, 1482, 1453, 1462, 1724, 1726, 1728, 1744, 1719, 1736, 1712, 1720, 587, 1009]. Kingdom [1394]. Kinglake [518]. Kinsmen [1492]. Klondyke [1603]. Knew [1595, 1666]. Knight [305, 359]. Knights [508]. Knocks [436]. Kong [1031]. Korner [838]. Kreutzer [666]. Kwaidan [1164].

L’Abbesse [769]. Labour [1392, 1063, 1460, 1718]. Lace [1219]. Lacedaemonians [1132]. Laches [1533]. Laddie [267]. Lady [912, 727, 1500, 1570, 1576, 1646, 302, 1013, 377, 762, 1338]. lagoon [374]. Laider [733]. Lair [1142]. Laird [1616]. Lame [477]. Lammermoor [452]. Lamp [58, 1774, 404]. Lamplighter [893]. Land [346, 55, 530, 1018, 1103, 1275, 829, 1697, 1332]. Landing [117]. Landmark [259]. LandSat [730]. Langeais [450]. Lanier [1178]. Lapham [155]. Large [1668]. Lark [45]. Last [1686, 906, 694, 710, 1514, 285, 1202, 1348]. Later [1211]. Latin [200, 796, 799, 797, 798, 208, 255, 209, 213]. Latitude [1759]. Latter [1094]. Latter-Day [1094]. Laughed [1551]. Lavender [1219]. Lavengro [433]. Law [1570, 871]. Lawn [1403]. Laws [1696]. Lay [354]. Lays [818]. Lazarillo [301, 418]. Lazy [517, 859]. Leagues [165]. Lear [1082, 1482, 1736]. Leaves [1276]. Lectures [1314, 1585, 1350, 746]. Legacy [1373]. Legal [1643]. Legatees [1336]. Legend [42, 1413]. Legends [1445, 322, 319]. Legs [158]. Leisure [804]. Lemorne [852]. Lenox [431]. Lescaut [449]. Lesser [1589, 1622]. Lesson [864]. Letter [563, 1775, 724, 35, 253]. Letters [580, 271, 1097, 386, 857, 1444, 701, 1443, 1347, 632, 599, 614, 368]. Letty [302]. Liber [797]. Liberty [8]. Libraries [1563]. Library [1773, 881]. Life [295, 246, 794, 1169, 1355, 1608, 499, 1525, 727, 1513, 471, 1039, 1537, 1097, 315, 847, 1763, 653, 25, 791, 1649, 1769, 1444, 502, 889, 521, 687, 1005, 1159, 1230, 1437, 1757, 632, 1404, 1410, 1086, 633, 814, 905, 913, 1659, 1033, 227, 1000, 817]. Life/Adventures [471]. Light [1598, 1049, 674, 169]. Lights [1449]. Like [1075, 1473, 1730]. Lilies [1247]. Lilith [1588]. Lily [1518, 1161]. Limberlost [126]. Lin [1337]. Lincoln [872, 1757, 1659]. Line [432]. Lion [1568, 620]. Lirriper [1368, 1373]. Lisbon [1100]. List [194]. Literary [795, 352]. Literature [295, 358, 1347]. Little [493, 1548, 346, 1330, 147, 460, 1578, 1560, 929, 1511, 848, 1543, 477, 1119, 679]. Lives [651]. Livin [309]. LOC [54]. Lock [1773]. Locomotive [1318]. Locrine [1498]. Lodgings [1368]. Log [64, 1750, 128]. London [1103, 1372]. Lone [990]. Long [158]. Longfellow [964, 965, 966, 967]. Looking [601, 14]. Lord [460, 913, 745]. Lorna [811]. Lose [1706]. Lost [755, 925, 365, 150, 140, 1749, 266, 22, 28, 1063, 1460, 1718]. Louis [696, 567, 1300, 314]. Love [1166, 1198, 424, 841, 687, 1063, 1460, 1718, 423]. Lover [1091, 1493]. Lovers [199]. Lower [1810]. Lucasta [680]. Lucile [1794]. Luck [1093]. Lucrece [1455, 1456]. Luggage [1366]. Lunacy [1362]. Lunar [117]. Lune [771]. Lung [1030, 1220]. Lure [977]. Lute [371]. Luther [188]. Lying [430]. Lyrical [1165]. Lyrics [767]. Lysis [1528].

Maastricht [663]. Macbeth [1083, 1483, 1737]. Machiavelli [1010]. Machine [873, 37]. Mad [345]. Madam [1646]. Madame [1311]. Madden [1771]. Madding [29, 108]. Made [388]. Mae [1771]. Magazine [671, 203]. Magdalen [1571]. Maggie [428]. Magi [2]. Magic [58, 1261, 400, 410]. Magnificent [1092]. Maid [1306, 73, 932]. Main [245, 522]. Maitre [1406]. Majorie [1704]. Make [1765]. Make-Believe [1765]. Makers [1236, 1024]. Making [1563]. Malbone [956]. Malefactor [1268]. Malta [867]. Man [1755, 759, 1595, 1666, 1644, 1535, 1181, 1765, 1706, 621, 715, 701, 195, 1810, 1620, 998, 296, 952, 71, 1167, 980]. Manalive [1664]. Manchu [174, 1137]. Mandeville [754]. Mania [1685]. Manifesto [62]. Mankind [726]. Manoeuvres [1766]. Manon [449]. Manse [491]. Mansfield [142]. Mansion [681]. Mansions [908]. Many [242]. Many-Coloured [242]. Map [221]. Marcas [1783]. March [1611]. Margaret [323, 1615]. Maria [135]. Marian [932]. Marie [1639]. Marigold [1367]. Mariner [152]. Marion [889, 814, 817]. Market [279]. Market-Place [279]. Marner [529]. Marriage [1506]. Married [342]. Marries [1620]. Mars [581, 65, 63, 73, 69, 1107, 542]. Martin [188, 934, 1019, 1228, 1640]. Marvelous [55]. Mary [1417]. Mass [1174]. Massacre [1447]. Massimilla [1753]. Master [417, 565, 864, 1302, 835, 952]. Masterman [1364]. Masterpiece [1503]. Mastery [749]. Mates [999]. Math [194]. Mathematical [611]. Mayflower [9]. Mayor [144]. McChesney [434]. McGuffey [1441, 1442]. McLean [1337]. McTeague [166]. Me [8, 521, 163]. Mean [716, 610]. Meaning [1510]. Measure [1080, 1480, 1734]. Medici [1796, 563]. Medicine [719, 1515]. Medium [1785]. Melmoth [1231]. Memoir [1287, 675]. Memoirs [1026, 805, 613, 690, 855]. Memorabilia [1131]. Memorial [567]. Memories [362]. Memory [694]. Men [1260, 97, 658, 1418, 300, 1652, 289, 1050, 812, 1507, 702, 325, 406, 976]. Menexenus [1631]. Meno [235, 1591]. Mental [1350]. Merchant [1068, 1465, 1723]. Mercies [838]. Merry [1789, 1070, 1467, 1725, 325]. Mersenne [70]. Message [1143, 420, 1359]. Messenger [1228, 1648]. Messengers [1761]. Michael [1676, 1784]. Middle [1563]. Middlemarch [146]. Midsummer [1067, 1464, 1722]. Mignon [1434]. Miles [562]. Miller [190]. Million [716, 717, 51, 130, 605, 606, 607, 608, 609, 612]. Millions [286]. Mills [847]. Mind [426]. Mine [831]. Miracle [416]. Mirror [1031, 1021, 1615]. Mirth [265]. Misalliance [909]. Miscellaneous [611, 1387, 1262]. Misérables [136]. Misery [425]. Miss [1569, 1688, 343, 342]. Missionary [1002]. Mississippi [227]. Mistress [1409]. MO [642]. Moby [17]. Mode [1593]. Modern [1515, 1119, 1120]. Modest [1034]. Modeste [1434]. Moghul [1422]. Mohicans [906]. Moll [351, 1687]. Mond [772]. Money [1576, 1694]. Mongers [416]. Monk [1579, 578]. Monster [97]. Monte [1138]. Montezuma [1790]. Monthly [953]. Montrose [1413]. Moon [692, 1581, 1043, 1401, 204, 737, 84, 976, 1216]. Moon-Face [1043]. Moonstone [156]. Moor [1081]. Moral [744, 390]. Morals [354]. Moran [302]. Moravians [547]. Moreau [160]. Mormon [19]. Morn [593]. morrow [441, 475]. Mort [1205]. Mosses [491]. Most [1502]. Mother [34, 1339]. Motion [117, 237]. Motorboat [541]. Mountain [1818]. Mountains [727, 446]. MPG [237]. Mr [539, 43, 44]. Mrs [1569, 1368, 1373, 838, 822, 1051]. Much [1595, 1666, 1072, 1469, 1470, 1727]. Mucker [312]. Mudfog [878]. Mugby [1371]. Munro [271]. Murray [1287]. Musketeers [1210]. Mutual [854]. My [224, 1576, 1707, 184, 851, 1615].

Mysterious [834, 1221]. Mystery [1627, 516, 543, 1599, 1773, 1634, 1813, 1446]. Myth [1024]. Myth-Makers [1024]. Myths [1024, 322].

Nada [1161]. NAFTA [90]. Nagasaki [662]. Name [1390]. Names [796]. Narrative [1623, 25, 822]. National [259]. Nations [1502]. Native [123, 1404]. Natural [64, 1360, 128]. Nature [1187, 1709, 715, 598, 757]. Nebendahl [841]. Need [1437]. Nelson [913]. Newspapers [714]. Next [1101]. Nibelungenlied [1105]. Nicholas [933, 840]. Nickleby [933]. Nicolete [1527]. Nig [561]. Night [1635, 992, 294, 1067, 1077, 1464, 1476, 1477, 1722, 310, 1192, 1199]. Night-Born [992]. Nightingale [569]. Nights [129, 810, 725]. Nine [537]. Ninth [1795]. Njal [574]. No [1390, 1577, 1375, 1652, 825]. Noble [1492]. Nobody [989]. Noir [770]. Nonsense [948]. North [1175, 207, 1807]. Northanger [122]. Norway [756]. Notes [1097, 652, 577, 943, 363]. Nothing [1072, 1469, 1470, 1727]. Notice [743]. Nov [953]. Novel [461, 1052]. Novelist [1613]. Novels [1578, 1313]. NREN [57]. Nucingen [1248]. Nuclear [661]. Number [64, 128, 283]. Numbers [66].

O [26, 309, 996]. O.T. [1558]. Oakdale [344]. Obscure [154]. Observations [1411]. Occurrence [356]. O’Connor [1751]. Oct [953]. Octopus [249]. Odyssey [1674, 1673]. Oedipus [33]. Off [1307]. Official [527]. Ogilvy [323]. O’Hara [1551]. Ohio [397]. Old [1803, 1306, 1624, 1146, 677, 320, 222, 491, 1773, 1792, 767, 1219, 1563, 319]. Oldport [956]. Oliver [707]. Omar [228]. Once [658, 1418]. One [717, 470, 760, 276, 51, 605, 606, 607, 608, 609, 612]. Online [81]. Only [1799]. Opar [93]. Open [253]. Opera [176]. Options [1532]. Orations [862, 208]. Orcival [1599]. Oregon [978]. Organic [1188]. Oriental [586]. Origin [1510, 1187, 1182]. Orlando [592]. Orthodoxy [131]. Othello [1081, 1481, 1735]. Other [1556, 628, 161, 878, 298, 1411, 1438, 391, 1662, 1694, 1151, 245, 244, 569, 405, 984, 1043, 1005, 110, 410, 600, 994, 868, 1565, 1342, 1781, 1548, 1795, 1678, 665, 1418, 687, 1029, 1618, 666, 1020, 745]. Otis [699]. Otranto [673]. Otto [353]. Our [854, 879, 1643, 561]. Outcast [615]. Outdoor [1745]. Outfield [366]. Outlaw [350]. Outline [1430]. Owl [356]. Own [564, 1444]. Oz [55, 56, 401, 496, 400, 467, 466, 927, 925, 921, 924, 923, 922, 926]. Ozma [467].

P1 [955]. Padre [1340]. Padua [846]. Pagan [1510]. Pageant [395]. Pages [553]. Paid [438]. Painted [1816]. Paintings [230]. Pair [206]. Palliano [775]. Pamphlets [1094]. Pan [18, 1286, 370]. Pantagruel [1154]. Paper [1036]. Papers [602, 557, 1387, 1415, 20, 1356, 193, 490, 1044, 1414]. Paradise [777, 22, 28, 59, 970, 966]. Paradiso [974, 962]. Paraguay [1431]. Parasite [336]. Parents [1612, 874]. Paris [1447]. Parish [1264]. Park [142]. Parme [768]. Parmenides [1636]. Parnassus [1023]. Parody [1430].

Part [1069, 1071, 1054, 1055, 1056, 1466, 1468, 1450, 1451, 1452, 1711, 1724, 1726, 276, 1538]. Partridge [1044]. Party [1381]. Passing [836]. Passion [1505, 1011]. Passionate [1494, 1495]. Past [734]. Past/Present [734]. Pastor [1777]. Patchwork [921]. Pathfinder [1820]. Pathology [430]. Patriotic [903]. Patrol [877]. Paul [274, 636, 1288]. Pauper [1779]. Paz [1323]. Peace [1363, 1766, 1225, 1429]. Peasantry [626]. Peddler [636]. Pellucidar [582]. Penelope [1232, 1171, 1343, 1809]. Penrod [383, 1112]. Pension [1424]. People [440, 531, 1694, 1637]. Perfect [1439, 1398]. Pericles [1487]. Perils [1358]. Perseverance [901]. Personal [1026, 664, 1659]. Personality [1524]. Perspectives [661]. Persuasion [106]. Peru [1163, 1277]. Peter [18, 1286, 1383]. Pethel [731]. PG [1025, 1027]. Phaedo [1606]. Phaedrus [1584]. Phantasmagoria [628]. Phantastes [306]. Phantom [176]. Pharsalia [579]. Phil [648]. Philanthropist [809]. Philebus [1690]. Philobiblon [603]. Philoktetes [778]. Philosopher [839]. Philosophy [425, 1301, 833]. Phineas [1525]. Phoenix [807, 1475]. Photographs [260]. Photography [169]. Pi [51, 717]. Pickwick [557]. Picture [237, 730, 175]. Pictures [117, 627, 169]. Picturesque [363]. Pieces [743, 843, 600]. Pierre [1184]. Pierrette [1653]. Pilgrim [132, 1494, 1495]. Pimpernel [61]. Pinocchio [481]. Pioneer [335]. Pioneers [26]. Pirates [939]. Pivot [1638]. Place [279, 1263]. Places [784, 1680, 64, 51, 128]. Plague [357]. Plain [1800]. Plains [591, 1252]. Planet [542]. Playboy [1194]. Plays [780, 696]. Plot [710]. Plutarch [651]. Pocket [1776]. Poems [239, 626, 551, 982, 243, 1233, 628, 1560, 298, 1516, 1241, 656, 1517, 1165, 928, 245, 244, 767, 1287, 556, 1183, 569, 405, 984, 1319, 680, 1005, 1335, 1140, 110, 378, 997, 903, 594, 422, 1423, 1421, 816, 390, 994, 1095, 1020]. Poet [809]. Poetical [1691]. poetry [371]. Poets [1395]. Point [1550]. Pointed [348]. Poison [127]. Polestar [275]. Political [955]. Politics [1295]. Polity [1132]. Polly [1750, 830]. Pollyanna [1402]. Pomegranates [844]. Pompeii [1514]. Ponkapog [602]. Pons [1798]. Pontificia [824]. Pook [536]. Pool [1777, 1122, 737]. Poor [1344, 465]. Poppy [983]. Popular [953, 426, 613, 690, 855]. Portraits [362]. Postscripts [1809]. Power [1519, 288]. PQ [643]. Practice [169]. Prayers [593]. Pre [699]. Pre-Revolutionist [699]. Prediction [70]. Preface [743]. Prejudice [1296]. Pres [1]. Prescott [274]. Present [734]. President [891]. Prester [588]. Previous [590]. Price [89, 438]. Pride [1296, 1596, 1667, 944]. Prime [66, 70]. Prince [1754, 365, 629, 1186, 477, 1399, 1076, 353, 1779, 868]. Princess [925, 147, 63, 308, 448, 674, 686, 685, 763]. Princesse [1298]. Prisoner [96]. Prisoners [1358]. Private [1415]. production [169]. Profession [1051]. Professor [991]. profitable [1511]. Profits [1508]. Profundis [887]. Progress [132]. Project [182, 553, 1258, 527]. Properti [219]. Property [341]. Prophetia [799]. Proposal [1034]. Proposed [667]. Pros [447]. Prose [1292]. Protagoras [1540]. Proud [465]. Provost [1250]. Prufrock [1411]. Psychological [703]. Psychology [1180, 1274, 429]. PT [1048]. Puck [536]. Puddler [1251]. Pudd’n’head [103]. Puerisque [367].

Punkin [936]. Pupil [995]. Purcell [490]. Purgatorio [973, 961]. Purgatory [969, 965]. Purple [1254]. Purse [1150]. Put [284]. Puzzle [710].

QREAD [236]. Quack [670]. Quality [1500]. Quatermain [688]. Queen [1218]. Queer [1645]. Quest [442]. Question [1759]. Quixote [959].

R [644, 1619]. R. [1287]. Race [285]. Racer [917]. Racing [541]. Racket [1629]. Rackrent [1376]. Radar [221]. Radio [67]. Raffles [684]. Railway [1815]. Rain [346]. Rainbow [511]. Rally [1797]. Ramanujan [659]. Ranch [1160]. Ranger [990]. Rape [1455, 1456]. Rasselas [629]. Ratio [610]. Raveloe [529]. Read [890]. Reader [1441, 1442, 236]. Reading [1005, 282, 1020]. Ready [1364]. Real [1410]. Reason [60]. Rebecca [479, 1329]. Recipes [1038]. Reconciled [1231]. Record [664]. Records [261]. Recruit [1378]. Red [1385, 404, 74, 444, 1679, 519, 1693, 760, 1813, 296]. Redburn [465]. Reddy [1767]. Redheaded [366]. Reef [264]. Regained [59]. Regime [1289]. Reginald [1811]. Registered [1775]. Religio [563]. Religion [1508]. Religious [598, 903, 390]. Remarkable [427]. Remarks [711, 712, 713]. Remember [268]. Reminiscences [1270, 785]. Renascence [110]. Report [527]. Reporter [388]. Reprinted [843]. Republic [151, 1448]. Resartus [1014]. Rescue [1658]. Research [1092]. Restoration [822]. Restored [1819]. Return [82, 109, 203, 123, 1137]. Revelation [1586]. Revenues [1133]. Revolution [1255, 429, 706]. Revolutionist [699]. Rewards [535]. Rezanov [472]. Rheims [1530, 1531, 1558]. Rhine [1226]. Rhymes [1797, 1692, 1593, 296, 290]. Rhythm [1230]. Richard [387, 1057, 1065, 1453, 1462, 1712, 1720]. Riddle [1226]. Ride [163]. Riders [1254, 957]. Rienzi [1348]. Rights [4]. Rime [152]. Ring [1439, 863]. Rinkitink [924]. Rio [285]. Riots [883]. Rise [155, 431]. River [1234, 195, 678, 996]. Riverman [1053]. Rivers [1511, 573]. Road [466, 297]. Roadmender [682]. Roads [1594, 667]. Robbery [1152]. Robe [1536]. Robert [696, 567, 632, 314, 633]. Robin [803, 930]. Robinson [500, 540]. Rocky [727]. Rocque [665]. Roderick [177]. Roget [24]. Rogue [1537]. Roland [372]. Rolf [1042]. Rolling [478, 290]. Roman [708, 861, 866, 1348]. Romance [811, 781, 956, 306, 374]. Romances [802]. Romany [403]. Rome [818]. Romeo [1066, 1463, 1721]. Room [1634, 1526]. Root [53, 130, 605, 606, 607, 608, 609, 612]. Rose [863, 996, 1565]. Rose-Bush [1565]. Rotating [237]. Rouge [770]. Round [404, 1416, 735]. Round-Up [735]. Roundabout [1414]. Rover [1116]. Royalty [1819]. Rubaiyat [228]. Rudge [883]. Ruins [1349]. Ruling [1011]. Runabout [916]. Runaway [1249]. Running [562]. Rupert [1099]. Russia [1280, 1278, 1303, 1811]. Rye [403]. Ryecroft [1415].

S [1026, 645]. Sabbath [593]. Sad [1789]. Saga [328, 574]. Sage [1254]. Saints [1195]. Salammbo [1244]. Salome [1293, 1294]. Saltbush [1271].

Sam [1112]. San [1509]. Sand [139]. Sanitary [1585]. Santa [498]. Sara [138]. Sarrasine [1768]. Sartain [391]. Sartor [1014]. Satire [558]. Savile [745]. Savoyard [900]. Sawyer [75, 92, 94]. Saxon [595, 634]. Say [1577]. Sayings [842]. Scapegoat [1257]. Scarecrow [923]. Scarlet [226, 339, 35, 61]. Sceaux [1259]. Scenes [1608]. Scholemaster [1786]. School [709, 1432]. Schoolmistress [1678]. Schools [1229]. Science [953, 1170, 1598, 1285, 1350, 486, 1654]. Scotland [1171]. Scoundrels [1580]. Scout [1238]. Scouts [1032, 1202]. Screw [191]. Sea [1803, 1700, 1018, 1021, 1359, 1028, 1162, 753, 957, 573, 165]. Sea-Gull [1700]. Seal [1693]. Search [1316]. Seas [445]. Seashore [1379]. Second [1752, 10, 1201, 235, 1120]. Secret [114, 1109, 202, 940, 1635, 649, 1680]. Secrets [1298]. Seed [983]. Seekers [742]. Seen [586]. Selborne [1360]. Select [1183]. Selected [208, 1266, 1269, 570, 1095, 1292, 327]. Self [901]. Senate [711, 712, 713]. Sense [162, 148]. Sensibility [162]. Sentiment [1370]. Sentimental [776]. Seraphita [1384]. Sergius [951]. Series [283]. Sesame [1247]. Set [1022]. Settlers [1305]. Seven [1260, 1344, 78, 1168]. Seventeen [1559]. Sexti [219]. Shadow [432, 1566, 568]. Shakespeare [1682, 1714, 550, 1139, 1240, 1382]. Shakespearian [1059]. Shandy [1033]. Shape [1817]. Sharer [202]. Shelley [1290, 1291]. Sherlock [109, 203, 805, 1609]. Shook [1322]. Shop [677]. Short [743, 161, 600, 276, 287]. Shrew [1061, 1458, 1716]. Shuttle [487]. Side [333, 777]. Sidney [1178]. Sign [1629]. Silas [529, 155]. Silver [736]. Silverado [495]. Silverman [782]. Simple [1206]. Sinking [753]. Sinners [631]. Sins [838]. Sioux [1253, 322]. Sir [754, 1497]. Sister [572, 215, 1677]. Sisters [982, 292]. Site [260, 258, 259]. Sixpence [204]. Skald [246]. Skates [736]. Sketches [878, 853, 882, 884, 728, 786]. Skin [1261]. Sky [917, 998]. Slander [1227]. Slave [100]. Sleeper [747]. Sleepy [42]. Slipper [1548]. Sliprails [1267]. Smalcald [254]. Small [398, 1619]. Smith [1619]. Smoke [1545]. Snark [15]. Snow [1776, 492]. Snowy [195]. Snyders [840]. Soames [732]. Social [795, 1585]. Socialist [1602]. Society [1374]. Sociology [458]. Soil [1369]. Sojourner [1623]. Soldiers [384]. Some [1270, 1419, 1414, 661, 1431]. Somebody [1366, 679]. Son [1407, 91, 792]. Sonata [666]. Song [21, 45, 1699, 372, 512]. Songs [626, 1233, 1612, 900, 993, 1789, 468, 423, 1677]. Sonnets [1004, 1059, 1496, 1013]. Sons [1369, 199]. Sophist [1681]. Sophocle [33]. Soul [1598, 321, 1206, 840, 1361, 1222, 980]. Souls [389, 1035, 1037]. South [1002, 1162, 1404, 445]. Southwest [295]. Spain [396, 544]. Spangled [704]. Spanish [301]. Speaks [789]. Species [1175, 1182]. Speech [1611]. Speeches [891, 795]. Spell [189]. Spirit [1193, 1173, 1263]. Splendid [1788]. Spoon [1234]. Sportsman [1134]. Spy [1760]. Square [53, 130, 605, 606, 607, 608, 609, 612]. Squatters [495]. St [794, 665, 303]. Stage [829]. Stage-Land [829]. Staircase [415]. Stand [1686]. Standfast [539]. Stands [469]. Star [704, 990, 1116]. Star-Rover [1116]. Star-Spangled [704]. Stark [271]. Stars [1598, 1049, 1168]. Start [1355].

State [4, 7]. States [221, 232, 233, 281, 891, 116]. Statesman [1684]. Steep [307]. Steets [428]. Step [1015]. Steps [537]. Stereo [157]. Sterling [1039]. Stevenson [696, 567, 314]. Stick [1710]. Stone [478, 290]. Stoops [364]. Stopper [1512]. Stories [671, 571, 1284, 455, 454, 161, 1419, 1243, 391, 1662, 222, 1266, 1502, 1773, 583, 1043, 1382, 340, 559, 936, 410, 1689, 1565, 1342, 1548, 1795, 327, 1678, 665, 1418, 687, 1618, 666, 745]. Storm [1657]. Stormfield [1007]. Story [574, 1106, 1440, 494, 482, 726, 564, 1006, 808, 742, 235, 1745, 1393, 335, 1650]. Straight [1333]. Strange [1121]. Strategy [1225]. Street [1655, 245, 522, 1168]. Strength [1029]. String [485]. Strogoff [1784]. Strong [328, 1029]. Studies [406]. Study [1327, 1660, 1777, 226, 598, 426, 1541, 518]. Style [1001]. Styles [834]. Sub [347]. Sub-Deb [347]. Submarine [915]. Sullivan [780]. Summer [395, 167]. Sunday [888]. Sunnybrook [479]. Super [250]. Super-Beasts [250]. Supernatural [1565]. Surfing [50]. Survivors [1600, 1647]. Susan [912, 431]. Swan [1699]. Swift [480, 915, 916, 917, 918, 919, 920, 1236, 1317, 1235, 1237, 1238, 1315, 1316, 1318]. Sword [983]. Sylvie [597]. Symphony [118, 157, 257]. Symposium [1549, 1135]. Syndicate [408]. Syntheses [1188].

T [1525]. Table [723, 510, 447]. Take [163]. Tale [99, 1751, 1088, 1489, 1742, 1040]. Tales [1546, 280, 692, 107, 516, 1156, 1151, 489, 942, 1652, 1800, 550, 1139, 1240, 1806, 877, 1162, 1603, 1313, 407, 868, 1029]. Talisman [1331]. Tamburlaine [1048, 1538]. Taming [1061, 1458, 1716]. Tanglewood [942]. Tank [920]. Tao [198]. Tapestried [1617]. Taras [1151]. Tarascon [1804]. Tariffs [90]. Tarn [793]. Tartarin [1804]. Tarzan [86, 82, 91, 79, 93, 107, 1353]. Tattine [1758]. Taverns [1003]. Te [198]. Tea [741]. Telegraph [945]. Telephone [790]. Telford [905]. Tell [455, 454, 1652, 1542]. Tempest [1089, 1490, 1743]. Temple [172]. Ten [1211, 1810]. Tenant [935]. Tennis [1403]. Terminal [80]. Terre [771]. Terror [1408, 516]. Tess [111]. Test [260]. Testament [1531]. Testaments [12, 1530]. Tether [506]. text [963, 975]. Theaetetus [1672]. Their [1510]. Theologico [955]. Theologico-Political [955]. Theology [486]. Theory [804]. There [311]. Theron [134]. Thesaurus [24]. Theses [255]. Things [757, 1781]. Thinkers [748]. Third [836]. Thirty [537, 545]. Thirty-Nine [537]. Thomas [713, 689, 1497, 905]. Thoroughfare [1375]. Thoughts [715, 820]. Thousand [562]. Three [888, 1243, 1210, 289, 288, 1003]. Threshold [657]. thru [200]. Thursday [1644]. Thuvia [73]. Tiddler [1365]. Tide [1553]. Tides [1016]. Tidy [1015]. Tiger [377]. Tik [922]. Tik-Tok [922]. Timaeus [1521]. Time [530, 531, 532, 1773, 879, 37]. Times [758]. Timon [1086, 1486, 1740]. Tin [926]. Tinker [1282]. Tirant [359]. Titanic [753]. Titles [796]. Titus [1060, 1457, 1715]. To-day [494]. To-morrow [441, 475]. Tok [922]. Told [489]. Tolstoy [785]. Tom [480, 915, 916, 917, 918, 919, 920, 1236, 1317, 1235, 1237, 1238, 1315, 1316, 1318, 1365, 1432, 821, 185, 75, 92, 94]. Tongue [1113]. Tono [695]. Too [1595, 1666]. Tormes [418]. Torn [350]. Touchstone [248]. Tour [859, 772]. Tours [1299]. Tower [902]. Towers [1772]. Town [1803]. Toys [1429]. Track [1185]. Tracks [478]. Trades [1645]. Traditional [58]. Tragedy [572, 1085, 1457]. Trail [978]. Trails [977, 307]. Tramp [120]. Transformation [764]. translated [1685]. Translating [253]. Translation [968, 969, 970, 971, 964, 965, 966, 967]. Translations [595]. Travel [604, 468]. Traveller [880]. Travellers [1344]. Travels [754, 1002, 514, 800]. Treasure [1707, 742, 121]. Treasury [549]. Treatise [399, 874, 955]. Treaty [90, 663]. Tree [1346]. Trees [1596, 1667, 244]. Trial [1155]. tribunes [1348]. Tried [1038]. Trigault [526]. Trilogy [33]. Trinity [260, 258, 259, 527]. Tristram [1033]. Trns [1668, 1619]. Troilus [238, 1078, 1478, 1732]. Troll [327]. Trooper [1383]. Troy [635, 381]. True [1440, 1038]. Truth [1623]. Tulip [931]. Tunnel [919]. Turmoil [1052]. Turn [191]. Turtle [1475]. TW [646]. Twelfth [1077, 1476, 1477]. Twelve [1689]. Twenty [1279, 1212]. Twice [489]. Twilight [571, 1697]. Twin [1272]. Twins [709, 691]. Twist [707]. ’Twixt [1018]. Two [53, 130, 1334, 1395, 1284, 280, 743, 1572, 99, 859, 860, 1062, 1459, 1492, 1717]. Typhoon [1096].

U [1160, 1686]. U. [1026]. U.S. [527]. Ulrich [841]. Unabridged [229, 637, 638, 639, 640, 641, 642, 643, 644, 645, 646, 647, 650]. Unbearable [534]. Uncle [1702, 936, 185]. Uncommercial [880]. Unconscious [1196]. Underdogs [528]. Underground [577, 1309]. Undersea [1316]. Understanding [979]. Underwoods [419]. Underworld [1372]. Undo [461]. Unfenced [1751]. Unicorn [1568]. Union [663]. United [4, 7, 221, 232, 233, 281, 891, 116]. Universal [234]. University [436]. Unknown [328, 301, 722, 754]. Unrest [1156]. Unseen [1438]. Unsocial [1602]. Untamed [1353]. Upon [380]. Ursula [1177]. USA [730]. Use [1437, 1563]. Usher [898].

V [733, 570, 1073, 1471, 1728, 1654]. V.86 [953]. V1 [1769, 613]. V2 [1649, 690]. V3 [855]. Vailima [368, 593]. Valley [1032, 1518, 1401]. Vanished [1431, 1648, 1012]. Vanity [576]. Vanya [1702]. Varieties [598]. Vaughan [1613]. Vendetta [1328]. Vengeance [526]. Venice [1068, 1081, 1465, 1723]. Venus [1008]. Vera [1785]. Verdugo [1377]. Verona [1062, 1459, 1717]. Veronica [503]. Verse [1258, 1119, 1120, 1153]. Verses [304, 137]. Version [183, 12, 1530, 1531, 1558, 157, 894]. Versus [852]. VI [1054, 1055, 1056, 1450, 1451, 1452, 1711]. Vibration [1795]. Vicar [1299]. Victoria [1218]. Vida [301]. VIII [1090, 1491, 1744]. Village [902]. Violet [618]. Violists [113]. Virgin [1582]. Virginian [1252]. Virginibus [367]. Vision [1788]. Visit [1007]. Vital [420]. Vittoria [774]. Voice [1396]. Vol [1420, 668, 669, 599, 614]. Volsungs [1106]. Volume [1180, 1598, 554]. Voyage [1036, 1345, 910, 1100, 1780, 1283, 1388, 513, 145]. Voyages [1108].

Vulgate [796, 799, 797, 798].

W [518]. Wagnerite [1439]. Waitstill [1650]. Wakes [747]. Waldeaux [316]. Walden [187]. Wales [625, 1046, 1102]. Walking [985]. Wall [437]. Wallet [1030]. War [920, 376, 1225, 1141, 579, 545, 408, 133, 765, 661, 38, 752, 1746]. War-Some [661]. Warden [596]. Ware [134]. Warfare [486]. Warlord [69]. Warren [1051]. Warship [1235]. Was [1644, 950, 196]. Washington [730]. Waste [1275]. Wasted [1762]. Watch [902]. Watch-Tower [902]. Water [1239, 981]. Water-Babies [981]. Way [1015, 456, 1246]. Ways [300, 360]. Weaver [529]. Web [284, 553]. Webster [229, 637, 638, 639, 640, 641, 642, 643, 644, 645, 646, 647, 650, 1611]. Wedding [1282]. Weir [361]. Well [170, 1079, 1479, 1733, 1195]. Went [1774]. Were [658, 1418]. West [1671, 1745]. Western [1049, 1194]. Westward [1802]. Wheels [1217]. Where [1354, 311]. Whirligigs [1544]. White [1330, 440, 1665, 560, 869, 876, 1687, 359, 1142]. Who [1595, 1666, 1644, 388, 1706, 1551, 950]. Wide [284, 196]. Wieland [764]. Wife [1535]. Wild [625, 197, 153]. Wildfell [935]. Wilding [1409]. Will [311, 1077]. William [217]. Willows [270]. Wilson [999, 103]. Wind [270, 207, 1565]. Windermere [762]. Windows [1816, 236]. Windsor [1070, 1467, 1725]. Winesburg [397]. Winter [1088, 1489, 1742]. Wisdom [205]. Within [1016, 871]. Wives [1070, 1467, 1725]. Wizard [1237, 56, 401]. Wolf [1028]. Woman [1327, 1660, 1675, 560, 1793, 1818, 1392, 825, 135]. Women [493, 358, 812, 1223]. Wonderful [56, 865]. Wonderland [13, 894, 115]. Woodlanders [463]. Woodman [926]. Woods [1042]. Work [881]. Works [1801, 554, 399, 1691, 668, 669, 101]. Workshop [54]. World [16, 27, 88, 49, 181, 284, 548, 1610, 1663, 140, 1246, 1101, 1438, 1322, 196, 170, 81, 1194, 104, 1022]. Worldly [360]. Worlds [1556, 38]. Worldwide [661]. Worm [1142]. Worship [1045]. Wouldbegoods [766]. Wreck [1417]. Wrecker [987]. Writing [473]. Writings [570]. Wrong [1663, 1534]. Wrongs [135]. ’Wuthering [743, 740]. Wyndham [1772]. Wyoming [1745].

X [669]. XZ [647].

Yankee [87, 849]. Yates [944]. Year [357, 1416, 545]. Years [1279, 1212, 1211, 860]. Yellow [617, 1634, 1791]. Yew [497]. Young [293, 882, 884, 724, 1628]. Youth [504]. Yukon [189].

Z [1783]. Zen [36]. Zenda [96]. Zincali [544]. Zola [705, 1027]. Zuleika

References

Clinton:1993:CIA

[1] Bill (William Jefferson) Clinton. Clinton’s Inaugural Address, US Pres Bill Clinton, volume na of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1993. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext93/clinton3.
zip.

Henry:1992:GM

[2] O. Henry. The Gift of the Magi, volume na of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext92/magi10.txt.

Anonymous:1971:DI

[3] Anonymous. Declaration of Independence, volume 1 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1971. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext90/when11.txt.

Anonymous:1972:USB

[4] Anonymous. The United States’ Bill of Rights, volume 2 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1972. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext90/bill11.txt.

Kennedy:1973:JFK

[5] John F. (John Fitzgerald) Kennedy. John F. Kennedy’s Inaugural Address, volume 3 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1973. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext90/jfk11.txt.

Lincoln:1973:GA

[6] Abraham Lincoln. Gettysburg Address, volume 4 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1973. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext90/getty11.txt.

Anonymous:1975:USC

[7] Anonymous. The United States’ Constitution, volume 5 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1975. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext90/const11.txt.

Henry:1976:GML

[8] Patrick Henry. Give Me Liberty Or Give Me Death, volume 6 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1976. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext90/liber11.txt.

Anonymous:1977:MC

[9] Anonymous. The Mayflower Compact, volume 7 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1977. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext90/mayfl11.zip.

Lincoln:1978:ALS

[10] Abraham Lincoln. Abraham Lincoln’s Second Inaugural Address, volume 8 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1978. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext90/linc211.txt.

Lincoln:1979:ALF

[11] Abraham Lincoln. Abraham Lincoln’s First Inaugural Address, volume 9 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1979. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext90/linc111.txt.

Anonymous:1989:BBB

[12] Anonymous. The Bible The Bible, Both Testaments, King James Version, volume 10 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1989. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext90/kjv10.zip.

Carroll:1991:AW

[13] Lewis Carroll. Alice in Wonderland, volume 11 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext91/alice30.zip.

Carroll:1991:TLG

[14] Lewis Carroll. Through the Looking Glass, volume 12 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext91/lglass18.zip.

Carroll:1991:HS

[15] Lewis Carroll. The Hunting of the Snark, volume 13 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext91/snark12.zip.

Anonymous:1991:CWF

[16] Anonymous. The 1990 CIA World Factbook, volume 14 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext91/world12.zip.

Melville:1991:MD

[17] Herman Melville. Moby Dick, volume 15 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/

etext91/moby.zip. Missing Chapter 72.

Barrie:1991:PP

[18] J. M. (James Matthew) Barrie. Peter Pan, volume 16 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext91/peter15a.zip;   ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/etext91/peter16.zip.

Anonymous:1991:BM

[19] Anonymous. The Book of Mormon, volume 17 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext91/mormon13.zip.

Hamilton:1991:FP

[20] Alexander Hamilton. The Federalist Papers, volume 18 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext91/feder16.zip.

Anonymous:1991:SH

[21] Anonymous. The Song of Hiawatha, volume 19 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext91/hisong12.zip.

Milton:1991:PL

[22] John Milton. Paradise Lost, volume 20 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext91/plboss10.zip.

Aesop:1991:AF

[23] Aesop. Aesop’s Fables, volume 21 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1991. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext91/aesop11.

zip. includes a ’life’ of Aesop.

Anonymous:1991:RT

[24] Anonymous. Roget’s Thesaurus, volume 22 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1991. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext91/roget13.zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext91/roget14a.zip.

Douglass:1992:NLF

[25] Frederick Douglass. Narrative of the Life of Frederick Douglass, volume 23 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext92/duglas11.zip.

Cather:1992:P

[26] Willa Cather. O Pioneers!, volume 24 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext92/opion11.zip.

Anonymous:1992:CWF

[27] Anonymous. The 1991 CIA World Factbook, volume 25 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/world91a.zip.

Milton:1992:PL

[28] John Milton. Paradise Lost, volume 26 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext92/plrabn12.zip.

Hardy:1992:FMC

[29] Thomas Hardy. Far From the Madding Crowd, volume 27 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/crowd13.zip.

Aesop:1992:FA

[30] Aesop. Fables of Aesop, volume 28 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1992. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext92/aesop10a.
zip.

US:1992:DCU

[31] US Census Bureau. Data From the 1990 Census, US Census Bureau, volume 29 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1992. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext92/uscen901.zip.

Anonymous:1992:BNE

[32] Anonymous. The Bible New Etext of Bible, volume 30 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/bible11.zip.

Sophocles:1992:SOT

[33] Sophocles. Sophocles’ Oedipus Trilogy, volume 31 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext92/oedip10.zip.

Gilman:1992:HMD

[34] Charlotte Perkins Gilman. Herland [for Mother’s Day], volume 32 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext92/hrlnd10.zip.

Hawthorne:1992:SL

[35] Nathaniel Hawthorne. The Scarlet Letter, volume 33 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext92/scrlt11.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext92/scrlt12.zip.

Kehoe:1992:ZAI

[36] Brendan P. Kehoe. Zen and the Art of Internet, volume 34 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/zen10.zip.

Wells:1992:TM

[37] H. G. (Herbert George) Wells. The Time Machine, volume 35 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/timem11.zip.

Wells:1992:WW

[38] H. G. (Herbert George) Wells. The War of the Worlds, volume 36 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext92/warw11.zip.

US:1992:UC

[39] US Census Bureau. The 1990 US Census, volume 37 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/uscen903.zip.

Anonymous:1992:HDC

[40] Anonymous. The Hackers’ Dictionary of Computer Jargon, volume 38 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext92/jargn10.zip.

Krol:1992:HGI

[41] Ed Krol. Hitchhiker’s Guide to the Internet, volume 39 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/hhgi10.zip.

Irving:1992:LSH

[42] Washington Irving. The Legend of Sleepy Hollow, volume 41 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/sleep10.zip.

Stevenson:1992:DJMa

[43] Robert Louis Stevenson. Dr. Jekyll and Mr. Hyde, volume 42 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/hyde10.zip.

Stevenson:1992:DJMb

[44] Robert Louis Stevenson. Dr. Jekyll and Mr. Hyde, volume 43 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/hydea10.zip.

Cather:1992:SL

[45] Willa Cather. Song of the Lark, volume 44 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext92/song10.zip.

Montgomery:1992:AGG

[46] L. M. (Lucy Maud) Montgomery. Anne of Green Gables, volume 45 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext92/anne11.zip.

Dickens:1992:CC

[47] Charles Dickens. A Christmas Carol, volume 46 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext92/carol12.zip.

Montgomery:1992:AA

[48] L. M. (Lucy Maud) Montgomery. Anne of Avonlea, volume 47 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1992. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext92/avon10.zip.

Anonymous:1993:WFU

[49] Anonymous. The World Factbook, US CIA, 1992 Edition, volume 48 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/world192.zip.

Polly:1993:SI

[50] Jean Armour Polly. Surfing the Internet, volume 49 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/surf10.zip.

Anonymous:1993:POM

[51] Anonymous. Pi to one million places, volume 50 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/pimil10.zip.

Montgomery:1993:AI

[52] L. M. (Lucy Maud) Montgomery. Anne of the Island, volume 51 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/iland10.zip.

Anonymous:1993:SRT

[53] Anonymous. The Square Root of Two to 100,000 digits, volume 52 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/2sqrt10.zip.

US:1993:LWE

[54] US Library of Congress. LOC Workshop on Etexts, US Library of Congress, volume 53 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1993. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext93/locet10.zip.

Baum:1993:MLO

[55] L. Frank (Lyman Frank) Baum. The Marvelous Land of Oz, volume 54 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/ozland10.zip.

Baum:1993:WWO

[56] L. Frank (Lyman Frank) Baum. The Wonderful Wizard of Oz, volume 55 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/wizoz10.zip.

Polly:1993:N

[57] Jean Armour Polly. NREN, volume 56 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/nren210.zip.

Anonymous:1993:AAM

[58] Anonymous. Aladdin and the Magic Lamp, Traditional, volume 57 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/alad10.zip.

Milton:1993:PR

[59] John Milton. Paradise Regained, volume 58 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/rgain10.zip.

Descartes:1993:DRD

[60] Rene Descartes. Descartes’ Reason Discourse, volume 59 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/dcart10.zip.

Orczy:1993:SP

[61] Emmuska Orczy Orczy, Baroness. The Scarlet Pimpernel, volume 60 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/scarp10.zip.

Marx:1993:CM

[62] Karl Marx. The Communist Manifesto, volume 61 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/manif11.zip.

Burroughs:1993:PM

[63] Edgar Rice Burroughs. A Princess of Mars, volume 62 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/pmars11.zip.

Anonymous:1993:NEN

[64] Anonymous. The Number “e” [Natural Log to 100,000 places], volume 63 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/ee610.zip.

Burroughs:1993:GM

[65] Edgar Rice Burroughs. Gods of Mars, volume 64 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/gmars11.zip.

Anonymous:1993:FPN

[66] Anonymous. The First 100,000 Prime Numbers, volume 65 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/prime12.zip.

Joly:1993:DAR

[67] Norman F. Joly. The Dawn of Amateur Radio, volume 66 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/radio10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext93/radio10a.zip.

Coombs:1993:BE

[68] Norman Coombs. Black Experience, volume 67 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/blexp10.zip.

Burroughs:1993:WM

[69] Edgar Rice Burroughs. Warlord of Mars, volume 68 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/wmars12.zip.

Anonymous:1993:MPM

[70] Anonymous. The 32nd Mersenne Prime, Mersenne Prediction, volume 69 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext93/32pri10.zip.

Twain:1993:WM

[71] Mark Twain. What Is Man?, volume 70 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/wman10.zip.

Thoreau:1993:CD

[72] Henry David Thoreau. Civil Disobedience, volume 71 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/civil10.zip.

Burroughs:1993:TMM

[73] Edgar Rice Burroughs. Thuvia, Maid of Mars, volume 72 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/mmars10.zip.

Crane:1993:RBC

[74] Stephen Crane. Red Badge of Courage, volume 73 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/badge10.zip.

Twain:1993:ATS

[75] Mark Twain. Adventures of Tom Sawyer, volume 74 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/sawyr10.zip.

Goodwin:1993:E

[76] John Goodwin. Email 101, volume 75 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/email025.zip.

Twain:1993:AHF

[77] Mark Twain. Adventures of Huckleberry Finn, volume 76 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/hfinn10.zip.

Hawthorne:1993:HSG

[78] Nathaniel Hawthorne. House of the Seven Gables, volume 77 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/7gabl10.zip.

Burroughs:1993:TA

[79] Edgar Rice Burroughs. Tarzan of the Apes, volume 78 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/tarzn10.zip.

Schartau:1993:TC

[80] Win Schartau. Terminal Compromise, volume 79 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/termc10.zip.

Presno:1993:OW

[81] Odd de Presno. The Online World, volume 80 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/online11.zip.

Burroughs:1993:RT

[82] Edgar Rice Burroughs. Return of Tarzan, volume 81 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/tarz210.zip.

Scott:1993:I

[83] Sir Walter Scott. Ivanhoe, volume 82 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1993. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext93/ivnho11.
zip.

Verne:1993:EM

[84] Jules Verne. From the Earth to the Moon, volume 83 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/moon10.zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/moon10a.zip.

Shelley:1993:F

[85] Mary Wollstonecraft Shelley. Frankenstein, volume 84 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext93/frank10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext93/frank10a.zip.

Burroughs:1993:BT

[86] Edgar Rice Burroughs. Beasts of Tarzan, volume 85 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/tarz310.zip.

Twain:1993:CY

[87] Mark Twain. A Connecticut Yankee, volume 86 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1993.

Anonymous:1993:CWF

[88] Anonymous. The 1993 CIA World Factbook, volume 87 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/world93.zip.

Anonymous:1993:P

[89] Anonymous. Price, volume 88 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1993. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext93/price10.
zip.

Anonymous:1993:NTA

[90] Anonymous. NAFTA, Treaty, Annexes, Tariffs, volume 89 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/naftannx.zip;   ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/etext93/naftarif.zip;   ftp://uiarchive.
cso.uiuc.edu/pub/etext/gutenberg/etext93/naftchap.zip.

Burroughs:1993:ST

[91] Edgar Rice Burroughs. Son of Tarzan, volume 90 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/tarz410.zip.

Twain:1993:TSA

[92] Mark Twain. Tom Sawyer Abroad, volume 91 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/sawy210.zip.

Burroughs:1993:TJO

[93] Edgar Rice Burroughs. Tarzan, Jewels of Opar, volume 92 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext93/tarz510.zip.

Twain:1993:TSD

[94] Mark Twain. Tom Sawyer Detective, volume 93 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/sawy310.zip.

Cather:1993:AB

[95] Willa Cather. Alexander’s Bridge, volume 94 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/alexb10.zip.

Hope:1993:PZ

[96] Anthony Hope. The Prisoner of Zenda, volume 95 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1993. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext93/zenda10.zip.

Burroughs:1994:MM

[97] Edgar Rice Burroughs. The Monster Men, volume 96 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/monst10.zip.

Abbott:1994:F

[98] Edwin A. Abbott. Flatland, volume 97 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/flat11.zip.

Dickens:1994:TTC

[99] Charles Dickens. A Tale of Two Cities, volume 98 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/2city11.zip.

Douglass:1994:CAF

[100] Frederick Douglass. Collected Articles of Frederick Douglass, a Slave, volume 99 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext94/dugl210.zip.

Shakespeare:1994:CWW

[101] William Shakespeare. The Complete Works of William Shakespeare, volume 100 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext94/shaks12.zip.

Sterling:1994:HC

[102] Bruce Sterling. Hacker Crackdown, volume 101 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/hack11a.zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/hack12.zip.

Twain:1994:PW

[103] Mark Twain. Pudd’n’head Wilson, volume 102 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/puddn10.zip.

Verne:1994:AWD

[104] Jules Verne. Around the World in 80 Days, volume 103 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/80day11.zip.

Roosevelt:1994:FDR

[105] Franklin Delano Roosevelt. Franklin Delano Roosevelt’s 1st Inaugural Address, volume 104 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext94/fdr10.zip.

Austen:1994:P

[106] Jane Austen. Persuasion, volume 105 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/persu11.
zip.

Burroughs:1994:JTT

[107] Edgar Rice Burroughs. Jungle Tales of Tarzan, volume 106 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/tarz610.zip.

Hardy:1994:FMC

[108] Thomas Hardy. Far From the Madding Crowd, volume 107 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/crowd10a.zip.

ConanDoyle:1994:RSH

[109] Sir Arthur Conan Doyle. The Return of Sherlock Holmes, volume 108 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/rholm10.zip.

Millay:1994:ROP

[110] Edna St. Vincent Millay. Renascence and Other Poems, volume 109 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/ednam10.zip.

Hardy:1994:TD

[111] Thomas Hardy. Tess of the d’Urbervilles, volume 110 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/tess10.zip.

Stratton-Porter:1994:F

[112] Gene Stratton-Porter. Freckles, volume 111 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/freck10.zip.

McGowan:1994:V

[113] Richard McGowan. Violists, volume 112 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/viols10.zip.

Burnett:1994:SG

[114] Frances Hodgson Burnett. The Secret Garden, volume 113 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/gardn11.zip.

Tenniel:1994:TIA

[115] John Tenniel. Tenniel’s Illustrations for Alice in Wonderland, volume 114 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/algif10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext94/algif10.

US:1994:USC

[116] US Census Bureau. United States Census Figures Back To 1630, volume 115 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/uscen10b.zip.

Anonymous:1994:MPA

[117] Anonymous. Motion Pictures of the Apollo 11 Lunar Landing, volume 116 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext94/landiavi.zip;   ftp://
uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/landimpg.
zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/
landiqt.zip.

Beethoven:1994:BFSa

[118] Ludwig van Beethoven. Beethoven’s Fifth Symphony, volume 117 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/lvb5s11.zip.

EFF:1994:BDG

[119] EFF. Big Dummy’s Guide To The Internet, volume 118 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/bigd22.zip.

Twain:1994:TA

[120] Mark Twain. A Tramp Abroad, volume 119 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/tramp11.zip.

Stevenson:1994:TI

[121] Robert Louis Stevenson. Treasure Island, volume 120 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/treas10.zip.

Austen:1994:NA

[122] Jane Austen. Northanger Abbey, volume 121 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/nabby10.zip.

Hardy:1994:RN

[123] Thomas Hardy. Return of the Native, volume 122 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/nativ10.zip.

Burroughs:1994:AEC

[124] Edgar Rice Burroughs. At the Earth’s Core, volume 123 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/ecore10.zip.

Anonymous:1994:ADB

[125] Anonymous. Apocrypha Deutercanonical Books of the Bible, volume 124 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/apoc10.zip.

Stratton-Porter:1994:GL

[126] Gene Stratton-Porter. A Girl of the Limberlost, volume 125 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/limbr10.zip.

ConanDoyle:1994:PB

[127] Sir Arthur Conan Doyle. The Poison Belt, volume 126 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/poisn10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext94/poisn10h.zip.

Anonymous:1994:NEN

[128] Anonymous. Number “e” (“natural log”) to 1,000,000 places, volume 127 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/ee710.zip.

Lang:1994:AN

[129] Andrew Lang. The Arabian Nights, volume 128 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/arabn11.zip.

Anonymous:1994:SRT

[130] Anonymous. The Square Root of Two to 5 million digits, volume 129 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/2sqrt10a.zip.

Chesterton:1994:O

[131] G. K. (Gilbert Keith) Chesterton. Orthodoxy, volume 130 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/ortho10.zip.

Bunyan:1994:PP

[132] John Bunyan. The Pilgrim’s Progress, volume 131 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/plgrm10.zip.

SunTzu:1994:AW

[133] Sun Tzu. The Art of War, volume 132 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/sunzu10.zip.

Frederic:1994:DTW

[134] Harold Frederic. The Damnation of Theron Ware, volume 133 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/dware10.zip.

Wollstonecraft:1994:MWW

[135] Mary Wollstonecraft. Maria or the Wrongs of Woman, volume 134 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/maria10.zip.

Hugo:1994:M

[136] Victor Marie Hugo. Les Misérables, volume 135 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/lesms10.zip.

Stevenson:1994:CGV

[137] Robert Louis Stevenson. A Child’s Garden of Verses, volume 136 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/child10.zip.

Burnett:1994:SC

[138] Frances Hodgson Burnett. Sara Crewe, volume 137 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/sarac10.zip.

Doumic:1994:BGS

[139] Rene Doumic. Biography of George Sand, volume 138 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/sandb10.zip.

ConanDoyle:1994:LW

[140] Sir Arthur Conan Doyle. The Lost World, volume 139 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/lostw10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext94/lostw10h.zip.

Sinclair:1994:J

[141] Upton Sinclair. The Jungle, volume 140 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/jungl10.zip.

Austen:1994:MP

[142] Jane Austen. Mansfield Park, volume 141 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/mansf10.zip.

Twain:1994:B

[143] Mark Twain. The $30,000 Bequest, volume 142 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994.

Hardy:1994:MC

[144] Thomas Hardy. The Mayor of Casterbridge, volume 143 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/mayrc10.zip.

Woolf:1994:V

[145] Virginia Woolf. The Voyage Out, volume 144 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/voout10.zip.

Eliot:1994:M

[146] George Eliot. Middlemarch, volume 145 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/mdmar10.zip.

Burnett:1994:LP

[147] Frances Hodgson Burnett. A Little Princess, volume 146 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994.

Paine:1994:CS

[148] Thomas Paine. Common Sense, volume 147 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/comsn10.zip.

Franklin:1994:ABF

[149] Benjamin Franklin. The Autobiography of Ben Franklin, volume 148 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/bfaut10.zip.

Burroughs:1994:LC

[150] Edgar Rice Burroughs. The Lost Continent, volume 149 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/lcont10.zip.

Plato:1994:R

[151] Plato. The Republic, volume 150 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/repub13.
zip.

Coleridge:1994:RAM

[152] Samuel Taylor Coleridge. The Rime of the Ancient Mariner, volume 151 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/rime10.zip.

Sprague:1994:WJ

[153] Ruth M. Sprague. Wild Justice, volume 152 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/wildj10.zip.

Hardy:1994:JO

[154] Thomas Hardy. Jude the Obscure, volume 153 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/jude11.zip.

Howells:1994:RSL

[155] William Dean Howells. Rise of Silas Lapham, volume 154 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/silap10.zip.

Collins:1994:M

[156] Wilkie Collins. The Moonstone, volume 155 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/mston10.zip.

Beethoven:1994:BFSb

[157] Ludwig van Beethoven. Beethoven’s Fifth Symphony, New Version in Stereo, volume 156 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext94/lvb5s10a.zip.

Webster:1994:D

[158] Jean Webster. Daddy-Long-Legs, volume 157 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/dlleg10.zip.

Austen:1994:E

[159] Jane Austen. Emma, volume 158 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/emma10.
zip.

Wells:1994:IDM

[160] H. G. (Herbert George) Wells. The Island of Doctor Moreau, volume 159 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/dmoro11.zip.

Chopin:1994:AOS

[161] Kate O’Flaherty Chopin. The Awakening and Other Short Stories, volume 160 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext94/awakn10.zip.

Austen:1994:SS

[162] Jane Austen. Sense and Sensibility, volume 161 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/sense11.zip.

Laxer:1994:TMR

[163] Mark Eliot Laxer. Take Me For A Ride, volume 162 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/tride10.zip.

Alcott:1994:FF

[164] Louisa May Alcott. Flower Fables, volume 163 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/ffabl10.zip.

Verne:1994:LUS

[165] Jules Verne. 20,000 Leagues Under the Sea, volume 164 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/2000010.zip.

Norris:1994:M

[166] Frank Norris. McTeague, volume 165 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/mcteg10.
zip.

Wharton:1994:S

[167] Edith Wharton. Summer, volume 166 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/summr10.
zip.

Humphrey:1994:HAD

[168] S. D. (Samuel Dwight) Humphrey. Handbook of American Daguerrotype, volume 167 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/amdag10.
zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/
amdgf10.zip.

Snelling:1994:HPA

[169] Henry Hunt Snelling. History and Practice of the Art of Photography or the production of pictures through the agency of light, volume 168 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext94/hipho10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext94/hiphogif.zip.

Morris:1994:WAW

[170] William Morris. The Well At The World’s End, volume 169 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/wwend10.zip.

Collins:1994:HH

[171] Wilkie Collins. The Haunted Hotel, volume 170 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/hhotl10.zip.

Rowson:1994:CT

[172] Susanna Rowson. Charlotte Temple, volume 171 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/chtem10.zip.

Morley:1994:HB

[173] Christopher Morley. The Haunted Bookshop, volume 172 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/hbook10.zip.

Rohmer:1994:IDF

[174] Sax Rohmer. The Insidious Dr. Fu Manchu, volume 173 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/fuman12.zip.

Wilde:1994:PDG

[175] Oscar Wilde. The Picture of Dorian Gray, volume 174 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/dgray10.zip.

Leroux:1994:PO

[176] Gaston Leroux. The Phantom of the Opera, volume 175 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/phant11.zip.

James:1994:RH

[177] Henry James. Roderick Hudson, volume 176 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/rhuds10.zip.

James:1994:A

[178] Henry James. The American, volume 177 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/theam10.zip.

James:1994:C

[179] Henry James. Confidence, volume 178 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1994. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext94/confi10.
zip.

James:1994:E

[180] Henry James. The Europeans, volume 179 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext94/theeu10.zip.

Anonymous:1994:CWF

[181] Anonymous. The 1994 CIA World Factbook, volume 180 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1994. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext94/world94.zip.

Anonymous:1995:PGE

[182] Anonymous. The Project Gutenberg Encyclopedia, volume 181–200 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/etext95/pge0112.zip   (Volume 1 of 28);
 ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext95/pge0110r.
zip (Volume   1 of 28).

Abbott:1995:FIV

[183] Edwin A. Abbott. Flatland [Illustrated Version], volume 201 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/flat10a.zip.

Douglass:1995:MBM

[184] Frederick Douglass. My Bondage and My Freedom, volume 202 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/bfree10.zip.

Stowe:1995:UTC

[185] Harriet Beecher Stowe. Uncle Tom’s Cabin, volume 203 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/utomc10.zip.

Chesterton:1995:IFB

[186] G. K. (Gilbert Keith) Chesterton. The Innocence of Father Brown, volume 204 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext95/infrb10.zip.

Thoreau:1995:W

[187] Henry David Thoreau. Walden, volume 205 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/waldn10.zip.

Anonymous:1995:MLK

[188] Anonymous. Martin Luther King Jr. Day Anthology #1, volume 206 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/1mlkd11.zip.

Service:1995:SY

[189] Robert William Service. The Spell of the Yukon, volume 207 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/spyuk11.zip.

James:1995:DM

[190] Henry James. Daisy Miller, volume 208 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/dasym10.zip.

James:1995:TS

[191] Henry James. Turn of the Screw, volume 209 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/tturn10.zip.

James:1995:IE

[192] Henry James. An International Episode, volume 210 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/intep10.zip.

James:1995:AP

[193] Henry James. The Aspern Papers, volume 211 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/asprn10.zip.

Anonymous:1995:FMC

[194] Anonymous. Factorial Math Constants A List of Factorial Math Constants, volume 212 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext95/factr10.zip.

Paterson:1995:MSR

[195] A. B. (Andrew Barton) Paterson. The Man From Snowy River, volume 213 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/snowy11.zip.

Lawson:1995:DWW

[196] Henry Lawson. In the Days When the World Was Wide, volume 214 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/dwwww11.zip.

London:1995:CW

[197] Jack London. The Call of the Wild, volume 215 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/callw10.zip.

Lao-tzu:1995:TTC

[198] Lao-tzu. Tao Te Ching, volume 216 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext95/taote10.
zip.

Lawrence:1995:SL

[199] D. H. (David Herbert) Lawrence. Sons and Lovers, volume 217 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/sonsa11.zip.

Anonymous:1995:CCL

[200] Anonymous. Caesar’s Commentaries in Latin [Books I thru IV], volume 218 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/galli10.zip.

Conrad:1995:HD

[201] Joseph Conrad. Heart of Darkness, volume 219 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/hdark11.zip.

Conrad:1995:SS

[202] Joseph Conrad. The Secret Sharer, volume 220 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/sshar11.zip.

ConanDoyle:1995:RSH

[203] Sir Arthur Conan Doyle. The Return of Sherlock Holmes [Magazine Edition], volume 221 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext95/rholm11b.zip.

Maugham:1995:MS

[204] W. Somerset (William Somerset) Maugham. Moon and Sixpence, volume 222 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/moona10.zip.

Chesterton:1995:WFB

[205] G. K. (Gilbert Keith) Chesterton. The Wisdom of Father Brown, volume 223 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/wifrb10.zip.

Hardy:1995:PBE

[206] Thomas Hardy. A Pair of Blue Eyes, volume 224 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/pblue10.zip.

MacDonald:1995:ABN

[207] George MacDonald. At the Back of the North Wind, volume 225 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/nwind10.zip.

Cicero:1995:COL

[208] Marcus Tullius Cicero. Cicero’s Orations [in Latin, Selected Orations], volume 226 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext95/ccero10.zip.

Virgil:1995:AL

[209] 70-19 BC Virgil. The Aeneid [in Latin], volume 227 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/anidl10.zip.

Virgil:1995:A

[210] 70-19 BC Virgil. The Aeneid, volume 228 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/anide10.zip.

Virgil:1995:Ba

[211] 70-19 BC Virgil. The Bucolics, volume 229 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/bucol10.zip.

Virgil:1995:Bb

[212] 70-19 BC Virgil. The Bucolics, volume 230 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/bucoe10.zip.

Virgil:1995:GL

[213] 70-19 BC Virgil. The Georgics [in Latin], volume 231 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/georl10.zip.

Virgil:1995:G

[214] 70-19 BC Virgil. The Georgics, volume 232 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/geore10.zip.

Dreiser:1995:SC

[215] Theodore Dreiser. Sister Carrie, volume 233 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/scarr10.zip.

Morris:1995:CC

[216] William Morris. Child Christopher, volume 234 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/chilc10.zip.

Salza:1995:WGI

[217] Giuseppe Salza. William Gibson Interviewed, volume 235 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/wmgib10.zip.

Kipling:1995:JB

[218] Rudyard Kipling. The Jungle Book, volume 236 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/jnglb10.zip.

Propertius:1995:SPC

[219] Sextus Propertius. Sexti Properti Carmina, volume 237 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/prpti10.zip.

Webster:1995:DE

[220] Jean Webster. Dear Enemy, volume 238 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/drnmy10.zip.

Anonymous:1995:RMU

[221] Anonymous. Radar Map of the United States, volume 239 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/usmpr10.zip.

Harris:1995:SOA

[222] Robert A. Harris. Stories From The Old Attic, volume 240 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/sftoa10.zip.

Brown:1995:CCH

[223] William Wells Brown. Clotelle; or The Colored Heroine, volume 241 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/clotl11.zip.

Cather:1995:MA

[224] Willa Cather. My Antonia, volume 242 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/myant10.zip.

Tolstoy:1995:FC

[225] Count Leo Nikoleyevich Tolstoy. The Forged Coupon, volume 243 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/forgd10.zip.

ConanDoyle:1995:SS

[226] Sir Arthur Conan Doyle. A Study In Scarlet, volume 244 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/study10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext95/study10h.zip   ((HTML)
).

Twain:1995:LM

[227] Mark Twain. Life on the Mississippi, volume 245 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/lmiss11.zip.

Fitzgerald:1995:ROK

[228] Edward Fitzgerald. The Rubaiyat of Omar Khayam, volume 246 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/rubai10.zip.

Anonymous:1995:WUD

[229] Anonymous. Webster’s Unabridged Dictionary, volume 247–248 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/etext95/wbstr11a.zip   (1st 100 pages);
 ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext95/wbstr11b.

zip (2nd 100 pages). includes 8-bit characters.

Anonymous:1995:FCP

[230] Anonymous. French Cave Paintings, volume 249 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/cavep11.zip.

Hart:1995:BHI

[231] Michael S. Hart. A Brief History of the Internet, volume 250 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/bhoti01.zip.

Anonymous:1995:USCa

[232] Anonymous. United States Congressional Address Book, 1995, volume 251 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/uscon95.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext95/uscon95a.zip;   ftp://
uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext95/uscon95b.
zip.

Anonymous:1995:USCb

[233] Anonymous. The United States Copyright Act of 1976, volume 252 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/uscpy76.zip.

Anonymous:1995:BUC

[234] Anonymous. The Berne Universal Copyright Convention, volume 253 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/berne10.zip.

Plato:1995:SSM

[235] Plato. The Second Story of Meno, volume 254 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/meno210.zip.

Scavezze:1995:QER

[236] Dan Scavezze. QREAD, Etext Reader for Windows, volume 255 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/qread121.zip.

Anonymous:1995:MMP

[237] Anonymous. MPG Motion Picture of Rotating Earth, volume 256 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/earth10.zip.

Chaucer:1995:TC

[238] Geoffrey Chaucer. Troilus and Crisyde, volume 257 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/troic10.zip.

Anonymous:1995:PA

[239] Anonymous. The Poems of A.L., volume 258 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/agord10.zip.

Service:1995:BC

[240] Robert William Service. Ballads of a Cheechako, volume 259 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/bchee10.zip.

Corson:1995:IB

[241] Hiram Corson. Introduction to Browning, volume 260 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/inbro10.zip.

Lowell:1995:DMG

[242] Amy Lowell. Dome of Many-Coloured Glass, volume 261 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/domcg10.zip.

Brooke:1995:PRB

[243] Rupert Brooke. Poems of Rupert Brooke, volume 262 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/rupbr10.zip.

Kilmer:1995:TOP

[244] Joyce Kilmer. Trees and Other Poems, volume 263 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/trees10.zip.

Kilmer:1995:MSO

[245] Joyce Kilmer. Main Street and Other Poems, volume 264 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/jkmst10.zip.

Anonymous:1995:LDC

[246] Anonymous. The Life and Death of Cormac the Skald, volume 265 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/cormc10.zip.

Gower:1995:CA

[247] John Gower. “Confessio Amantis”, volume 266 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/conam10.zip.

Wharton:1995:T

[248] Edith Wharton. The Touchstone, volume 267 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/touch10.zip.

Norris:1995:O

[249] Frank Norris. The Octopus, volume 268 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/octop10.zip.

Saki:1995:BS

[250] Saki. Beasts and Super-Beasts, volume 269 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/beast10.zip.

Grahame:1995:DD

[251] Kenneth Grahame. Dream Days, volume 270 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/drday10.zip.

Sewell:1995:BB

[252] Anna Sewell. Black Beauty, volume 271 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/bbeau10.zip.

Luther:1995:OLT

[253] Martin Luther. An Open Letter on Translating, volume 272 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/ltran11.zip.

Luther:1995:SA

[254] Martin Luther. The Smalcald Articles, volume 273 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/smcal10.zip.

Luther:1995:MLT

[255] Martin Luther. Martin Luther’s 95 Theses, In English and Latin, volume 274 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/the9510.zip.

Anonymous:1995:ACA

[256] Anonymous. The Augsburg Confession The Augsburg Confession, 465th Anniversary Edition, volume 275 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext95/augsb10.
zip.

Haydn:1995:FHS

[257] Franz Haydn. Franz Haydn’s 104th Symphony, volume 276 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/fh104s10.zip.

National:1995:TAB

[258] National Atomic Museum. Trinity (Atomic Bomb) Site, volume 277 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/1trnt10.zip. The 50th Anniversary of the

Atomic Bomb.

White:1995:TSN

[259] White Sands Missile Range Public Affairs Office. Trinity Site: 1945-1995. A National Historic Landmark, volume 278 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/

etext95/2trnt10.zip. The 50th Anniversary of the Atomic Bomb.

Anonymous:1995:TAB

[260] Anonymous. Trinity Atomic Bomb Test Site Photographs, volume 279 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/3trnt10.zip. The 50th Anniversary of the

Atomic Bomb.

Stevenson:1995:RFE

[261] Robert Louis Stevenson. Records of a Family of Engineers, volume 280 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/rfeng10.zip.

Stevenson:1995:FD

[262] Robert Louis Stevenson. Father Damien, volume 281 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/frdam10.zip.

Kinglake:1995:E

[263] Alexander William Kinglake. Eothen, volume 282 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/eothn10.zip.

Wharton:1995:R

[264] Edith Wharton. The Reef, volume 283 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/treef10.zip.

Wharton:1995:HM

[265] Edith Wharton. House of Mirth, volume 284 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/hmirt10.zip.

Hyne:1995:LC

[266] Charles John Cutcliffe Wright Hyne. The Lost Continent, volume 285 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/lostc10.zip.

Stratton-Porter:1995:L

[267] Gene Stratton-Porter. Laddie, volume 286 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/laddi10.zip.

Barr:1995:RA

[268] Amelia E. Barr, Edith Huddleston. Remember the Alamo, volume 287 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/alamo10.zip.

Cabell:1995:CH

[269] James Branch Cabell. The Certain Hour, volume 288 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/chour10.zip.

Grahame:1995:WW

[270] Kenneth Grahame. The Wind in the Willows, volume 289 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/wwill10.zip.

ConanDoyle:1995:SML

[271] Sir Arthur Conan Doyle. The Stark Munro Letters, volume 290 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/strkm10.zip.

Grahame:1995:GA

[272] Kenneth Grahame. The Golden Age, volume 291 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/gldna10.zip.

Taylor:1995:BBE

[273] Bayard Taylor. Beauty and The Beast, Etc., volume 292 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/bbetc10.zip.

Alger:1995:PPC

[274] Horatio Alger. Paul Prescott’s Charge, volume 293 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/presc10.zip.

ConanDoyle:1995:CP

[275] Sir Arthur Conan Doyle. The Captain of the Polestar, volume 294 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/polst10.zip.

Wharton:1995:ESFa

[276] Edith Wharton. The Early Short Fiction. Part One, volume 295 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/whrt110.zip.

Alger:1995:CB

[277] Horatio Alger. The Cash Boy, volume 296 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/cashb10.zip.

Tarkington:1995:F

[278] Booth Tarkington. The Flirt, volume 297 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/flirt10.zip.

Frederic:1995:M

[279] Harold Frederic. The Market-Place, volume 298 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/markt10.zip.

Boyesen:1995:TTH

[280] Hjalmar Hjorth Boyesen. Tales From Two Hemispheres, volume 299 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/twohe10.zip.

Anonymous:1995:USD

[281] Anonymous. United States Declaration of Independence in HTML, volume 300 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext95/1when10a.zip.

Wilde:1995:BRG

[282] Oscar Wilde. Ballad of Reading Gaol, volume 301 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/rgaol10.zip.

Anonymous:1995:FNS

[283] Anonymous. The Fibonacci Number Series, volume 302 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/fibns10.zip.

Anonymous:1995:HHP

[284] Anonymous. HomeBrew HomePages Put YOU On The World Wide Web, volume 303 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext95/homeb10.zip.

Paterson:1995:RGL

[285] A. B. (Andrew Barton) Paterson. Rio Grande’s Last Race, Etc., volume 304 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/rlast10.zip.

Gaboriau:1995:CM

[286] Emile Gaboriau. The Count’s Millions, volume 305 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/cntmi10.zip.

Wharton:1995:ESFb

[287] Edith Wharton. The Early Short Fiction, volume 306 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/whrt210.zip.

Paterson:1995:TEP

[288] A. B. (Andrew Barton) Paterson. Three Elephant Power Etc., volume 307 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/3elph10.zip.

Jerome:1995:TMB

[289] Jerome K. Jerome, (Jerome Klapka). Three Men in a Boat, volume 308 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/3boat10.zip.

Service:1995:RRS

[290] Robert William Service. Rhymes of a Rolling Stone, volume 309 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/rolst10.zip.

London:1995:BA

[291] Jack London. Before Adam, volume 310 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/badam10.zip.

Wharton:1995:BS

[292] Edith Wharton. Bunner Sisters, volume 311 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/bunnr10.zip.

Benet:1995:YA

[293] Stephen Vincent Benet. Young Adventure, volume 312 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/yngad10.zip.

Robinson:1995:CN

[294] Edwin Arlington Robinson. Children of the Night, volume 313 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/chnit10.zip.

Anonymous:1995:GLL

[295] Anonymous. Guide to Life and Literature of the Southwest, volume 314 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/swest10.zip.

Service:1995:RRC

[296] Robert William Service. Rhymes of a Red Cross Man, volume 315 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/redcr10.zip.

Montgomery:1995:GR

[297] L. M. (Lucy Maud) Montgomery. The Golden Road, volume 316 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/goldr10.zip.

Drake:1995:CFO

[298] Joseph Rodman Drake. Culprit Fay and Other Poems, volume 317 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/cufay10.zip.

London:1995:JB

[299] Jack London. John Barleycorn, volume 318 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/jbarl10.zip.

Gregory:1995:WM

[300] Eliot Gregory. The Ways of Men, volume 319 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/waymn10.zip.

Anonymous:1995:VLA

[301] Anonymous. Vida de Lazarillo, Author Unknown, In Spanish, volume 320 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/lazae10.zip.

Norris:1995:MLL

[302] Frank Norris. Moran of the Lady Letty, volume 321 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/morll10.zip.

Stevenson:1995:SI

[303] Robert Louis Stevenson. St. Ives, volume 322 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/stive10.zip.

Kipling:1995:V

[304] Rudyard Kipling. Verses 1889–1896, volume 323 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/11kip10.zip.

Fox:1995:KC

[305] John Fox. A Knight of the Cumberland, volume 324 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/kcumb10.zip.

MacDonald:1995:PFR

[306] George MacDonald. Phantastes, A Faerie Romance, volume 325 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/phafr10.zip.

Muir:1995:ST

[307] John Muir. Steep Trails, volume 326 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext95/sttrl10.
zip.

Davis:1995:PA

[308] Richard Harding Davis. The Princess Aline, volume 327 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/palin10.zip.

Guest:1995:HOL

[309] Edgar A. (Edgar Albert) Guest. A Heap O’ Livin’, volume 328 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/olivn10.zip.

Stevenson:1995:INE

[310] Robert Louis Stevenson. Island Nights’ Entertainments, volume 329 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/islni10.zip.

Rinehart:1995:WTW

[311] Mary Roberts Rinehart. Where There’s A Will, volume 330 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/awill10.zip.

Burroughs:1995:M

[312] Edgar Rice Burroughs. The Mucker, volume 331 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/tmuck10.zip.

Page:1995:BG

[313] Thomas Nelson Page. The Burial of the Guns, volume 332 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/bguns10.zip.

Raleigh:1995:RLS

[314] Sir Walter Alexander Raleigh. Robert Louis Stevenson, volume 333 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/rlswr10.zip.

Davis:1995:VBL

[315] Richard Harding Davis. Van Bibber’s Life, volume 334 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/vanbb10.zip.

Davis:1995:FW

[316] Rebecca Harding Davis. Frances Waldeaux, volume 335 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/fwald10.zip.

Eastman:1995:IHG

[317] Charles A. Eastman. Indian Heroes and Great Chieftains, volume 336 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/indhe10.zip.

Eastman:1995:IB

[318] Charles A. Eastman. Indian Boyhood, volume 337 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/indbo10.zip.

Zitkala-Sa:1995:OIL

[319] Zitkala-Sa. Old Indian Legends, volume 338 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/indle10.zip.

Eastman:1995:OID

[320] Charles A. Eastman. Old Indian Days, volume 339 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/inday10.zip.

Eastman:1995:SI

[321] Charles A. Eastman. The Soul of the Indian, volume 340 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/indso10.zip.

McLaughlin:1995:MLS

[322] Marie L. McLaughlin, b. 1842. Myths and Legends of the Sioux, volume 341 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/sioux10.zip.

Barrie:1995:MO

[323] J. M. (James Matthew) Barrie. Margaret Ogilvy, volume 342 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/marog10.zip.

Stevenson:1995:F

[324] Robert Louis Stevenson. Fables, volume 343 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/fbrls10.zip.

Stevenson:1995:MM

[325] Robert Louis Stevenson. Merry Men, volume 344 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/mrmen10.zip.

Stoker:1995:D

[326] Bram (Abraham) Stoker. Dracula, volume 345 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/dracu10.zip.

Cather:1995:TGS

[327] Willa Cather. The Troll Garden, and selected stories, volume 346 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/troll10.zip.

Anonymous:1995:GSG

[328] Anonymous. Grettir the Strong Grettir the Strong, Icelandic Saga, Author Unknown, volume 347 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1995. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext95/grttr10.
zip.

Homer:1995:CHH

[329] Homer. Collection of Hesiod, Homer and Homerica, volume 348 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/homer10.zip.

Stratton-Porter:1995:H

[330] Gene Stratton-Porter. The Harvester, volume 349 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/tharv10.zip.

Ferber:1995:FH

[331] Edna Ferber. Fanny Herself, volume 350 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/fnher10.zip.

Maugham:1995:HB

[332] W. Somerset (William Somerset) Maugham. Of Human Bondage, volume 351 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/humbn10.zip.

Ferber:1995:BS

[333] Edna Ferber. Buttered Side Down, volume 352 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/bsdwn10.zip.

McCrae:1995:FF

[334] John McCrae. In Flanders Fields, volume 353 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/infla10.zip.

Shaw:1995:SP

[335] Anna Howard Shaw. The Story of a Pioneer, volume 354 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/stpio10.zip.

ConanDoyle:1995:P

[336] Sir Arthur Conan Doyle. The Parasite, volume 355 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/prsit10.zip.

ConanDoyle:1995:BC

[337] Sir Arthur Conan Doyle. Beyond the City, volume 356 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/bcity10.zip.

Morris:1995:DJB

[338] William Morris. A Dream of John Ball, etc., volume 357 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/jball10.zip.

Davis:1995:SC

[339] Richard Harding Davis. The Scarlet Car, volume 358 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/sccar10.zip.

Olcott:1995:GSH

[340] Frances Jenkins Olcott. Good Stories for Holidays, volume 359 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/sthol10.zip.

Proudhon:1995:WP

[341] P.-J. (Pierre-Joseph) Proudhon. What is Property?, volume 360 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/pprty10.zip.

Porter:1995:MBM

[342] Eleanor H. (Eleanor Hodgman) Porter. Miss Billie Married, volume 361 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/msbim10.zip.

Porter:1995:MBD

[343] Eleanor H. (Eleanor Hodgman) Porter. Miss Billie’s Decision, volume 362 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/msbid10.zip.

Burroughs:1995:OA

[344] Edgar Rice Burroughs. The Oakdale Affair, volume 363 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/oakda10.zip.

Burroughs:1995:MK

[345] Edgar Rice Burroughs. The Mad King, volume 364 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/mdkng10.zip.

Austin:1995:LLR

[346] Mary Hunter Austin. The Land of Little Rain, volume 365 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/lndlr10.zip.

Rinehart:1995:BS

[347] Mary Roberts Rinehart. Bab: A Sub-Deb, volume 366 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/babsu10.zip.

Jewett:1995:CPF

[348] Sarah Orne Jewett. The Country of the Pointed Firs, volume 367 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/pfirs11.zip.

Conwell:1995:AD

[349] Russell Herman Conwell. Acres of Diamonds, volume 368 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/acrdi10.zip.

Burroughs:1995:OT

[350] Edgar Rice Burroughs. The Outlaw of Torn, volume 369 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/otorn11.zip.

Defoe:1995:MF

[351] Daniel Defoe. Moll Flanders, volume 370 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/mollf10.zip.

Wheatley:1995:LB

[352] Henry Benjamin Wheatley. Literary Blunders, volume 371 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/litbl10.zip.

Stevenson:1995:PO

[353] Robert Louis Stevenson. Prince Otto, volume 372 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/prott10.zip.

Stevenson:1995:LM

[354] Robert Louis Stevenson. Lay Morals, volume 373 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/lamor10.zip.

Bierce:1995:FF

[355] Ambrose Bierce. Fantastic Fables, volume 374 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext95/fanfb10.zip.

Bierce:1995:OAO

[356] Ambrose Bierce. An Occurrence At Owl Creek Bridge, volume 375 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/owlcr10.zip.

Defoe:1995:JPY

[357] Daniel Defoe. A Journal of the Plague Year, volume 376 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/jplag10.zip.

Barker:1995:KWL

[358] Nettie Garmer Barker. Kansas Women in Literature, volume 377 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/kswom10.zip.

Rudder:1995:WKT

[359] Robert S. Rudder. The White Knight: Tirant Lo Blanc, volume 378 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/whitk10.zip.

Gregory:1995:WWB

[360] Eliot Gregory. Worldly Ways and Byways, volume 379 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/bywys10.zip.

Stevenson:1995:WH

[361] Robert Louis Stevenson. Weir of Hermiston, volume 380 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/weirh10.zip.

Stevenson:1995:MP

[362] Robert Louis Stevenson. Memories and Portraits, volume 381 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/mempo10.zip.

Stevenson:1995:EPN

[363] Robert Louis Stevenson. Edinburgh Picturesque Notes, volume 382 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext95/edinn10.zip.

Goldsmith:1995:SSC

[364] Oliver Goldsmith. She Stoops to Conquer, volume 383 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1995. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext95/sstcq10.zip.

Burnett:1996:LP

[365] Frances Hodgson Burnett. The Lost Prince, volume 384 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/lprin10.zip.

Grey:1996:RO

[366] Zane Grey. The Redheaded Outfield, volume 385 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rhout10.zip.

Stevenson:1996:VP

[367] Robert Louis Stevenson. Virginibus Puerisque, volume 386 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/virpr10.zip.

Stevenson:1996:VL

[368] Robert Louis Stevenson. Vailima Letters, volume 387 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/valma10.zip.

Churchill:1996:C

[369] Winston Churchill. The Crossing, volume 388 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tcros10.zip.

Machen:1996:GGP

[370] Arthur Machen. The Great God Pan, volume 389 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ggpan10.zip.

Cranmer-Byng:1996:LJC

[371] L. (Launcelot) Cranmer-Byng. A Lute of Jade (Chinese poetry), volume 390 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/ljade10.zip.

Moncrief:1996:SR

[372] Moncrief. The Song of Roland, volume 391 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/sorol10.zip.

Tasso:1996:JD

[373] Torquato Tasso. Jerusalem Delivered, volume 392 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/jrdel10.zip.

Stacpoole:1996:BLR

[374] H. De Vere (Henry De Vere) Stacpoole. The Blue lagoon: A Romance, volume 393 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/blago10.zip.

Gaskell:1996:C

[375] Elizabeth Cleghorn Gaskell. Cranford, volume 394 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/crnfd10.zip.

Bunyan:1996:HW

[376] John Bunyan. The Holy War, volume 395 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/hlywr10.zip.

Stockton:1996:LT

[377] Frank R. (Frank Richard) Stockton. The Lady, or the Tiger?, volume 396 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/ladyt10.zip.

Milton:1996:FP

[378] John Milton. Four Poems, volume 397 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/miltp10.zip.

Platt:1996:BAE

[379] Rutherford Hayes Platt. 1st Book of Adam and Eve, volume 398 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/1adam10.zip.

Alger:1996:CUB

[380] Horatio Alger. Cast Upon the Breakers, volume 399 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/cubrk10.zip.

Teasdale:1996:HT

[381] Sara Teasdale. Helen of Troy, volume 400 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/helen10.zip.

Norris:1996:B

[382] Frank Norris. Blix, volume 401 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/blixn10.
zip.

Tarkington:1996:P

[383] Booth Tarkington. Penrod, volume 402 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/pnrod10.zip.

Davis:1996:SF

[384] Richard Harding Davis. Soldiers of Fortune, volume 403 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/soldf10.zip.

Smiles:1996:IB

[385] Samuel Smiles. Industrial Biography, volume 404 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/inbio10.zip.

Davis:1996:ALR

[386] Richard Harding Davis. Adventures and Letters of Richard Harding Davis, volume 405 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/alrhd10.zip.

Anonymous:1996:ARH

[387] Anonymous. Appreciations of Richard Harding Davis, volume 406 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/aprhd10.zip.

Davis:1996:RWM

[388] Richard Harding Davis. The Reporter Who Made Himself King, volume 407 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/rking10.zip.

DuBois:1996:SBF

[389] W. E. B. (William Edward Burghardt) Du Bois. The Souls of Black Folk, volume 408 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/soulb10.zip.

Wheatley:1996:RMP

[390] Phillis Wheatley. Religious and Moral Poems, volume 409 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/whtly10.zip.

Fox:1996:HFS

[391] John Fox. Hell Fer Sartain and Other Stories, volume 410 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/hfers10.zip.

Davis:1996:KJ

[392] Richard Harding Davis. The King’s Jackal, volume 411 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/kjack10.zip.

Hubbard:1996:JJA

[393] Elbert Hubbard. John Jacob Astor, volume 412 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/jastr10.zip.

Stevenson:1996:B

[394] Robert Louis Stevenson. Ballads, volume 413 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rlsba10.zip.

Jefferies:1996:PS

[395] Richard Jefferies. Pageant of Summer, volume 414 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/psumr10.zip.

Borrow:1996:BS

[396] George Henry Borrow. The Bible in Spain, volume 415 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tbisp10.zip.

Anderson:1996:WO

[397] Sherwood Anderson. Winesburg, Ohio, volume 416 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/wnbrg11.zip.

Luther:1996:HSC

[398] Martin Luther. The Hymns and Small Cathechism of Martin Luther, volume 417 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/thoml11.zip.

Luther:1996:TGW

[399] Martin Luther. A Treatise on Good Works, volume 418 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/gwork10.zip.

Baum:1996:MO

[400] L. Frank (Lyman Frank) Baum. The Magic of Oz, volume 419 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/magoz10.zip.

Baum:1996:DWO

[401] L. Frank (Lyman Frank) Baum. Dorothy and the Wizard in Oz, volume 420 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/doroz10.zip.

Stevenson:1996:K

[402] Robert Louis Stevenson. Kidnapped, volume 421 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/kdnpd10.zip.

Borrow:1996:RR

[403] George Henry Borrow. The Romany Rye, volume 422 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/trrye10.zip.

ConanDoyle:1996:RRL

[404] Sir Arthur Conan Doyle. Round The Red Lamp, volume 423 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/rrlmp10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext96/rrlmp10h.zip   ((HTML)
).

Lindsay:1996:GBO

[405] Vachel Lindsay. General Booth, Other Poems, volume 424 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/gnbth10.zip.

Stevenson:1996:FSM

[406] Robert Louis Stevenson. Familiar Studies of Men and Books, volume 425 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/fsomb10.zip.

Stevenson:1996:TF

[407] Robert Louis Stevenson. Tales and Fantasies, volume 426 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tlfns10.zip.

Stockton:1996:GWS

[408] Frank R. (Frank Richard) Stockton. The Great War Syndicate, volume 427 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/wsynd10.zip.

Hope:1996:FC

[409] Anthony Hope. Frivolous Cupid, volume 428 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/frcup10.zip.

Stockton:1996:MEO

[410] Frank R. (Frank Richard) Stockton. The Magic Egg and Other Stories, volume 429 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/mgegg10.zip.

Phillips:1996:GD

[411] David Graham Phillips. The Grain of Dust, volume 430 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tdust10.zip.

Phillips:1996:FH

[412] David Graham Phillips. The Fortune Hunter, volume 431 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/frhnt10.zip.

James:1996:A

[413] Henry James. The Ambassadors, volume 432 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ambas10.zip.

Phillips:1996:Ca

[414] David Graham Phillips. The Conflict, volume 433 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tconf10.zip.

Rinehart:1996:CS

[415] Mary Roberts Rinehart. The Circular Staircase, volume 434 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/circs10.zip.

Houdini:1996:MME

[416] Harry Houdini. The Miracle Mongers, An Expose’, volume 435 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/mmong10.zip.

Baum:1996:MK

[417] L. Frank (Lyman Frank) Baum. The Master Key, volume 436 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/mstky10.zip.

Rudder:1996:LT

[418] Robert S. Rudder. Lazarillo of Tormes, volume 437 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lazro13.zip.

Stevenson:1996:U

[419] Robert Louis Stevenson. Underwoods, volume 438 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/undrw10.zip.

ConanDoyle:1996:VM

[420] Sir Arthur Conan Doyle. The Vital Message, volume 439 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/vtmsg10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext96/vtmsg10h.zip   ((HTML)
).

Porter:1996:JD

[421] Eleanor H. (Eleanor Hodgman) Porter. Just David, volume 440 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/jusdv10.zip.

Stevenson:1996:NP

[422] Robert Louis Stevenson. New Poems, volume 441 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rlsnp10.zip.

Teasdale:1996:LS

[423] Sara Teasdale. Love Songs, volume 442 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lsongs10.zip.

Field:1996:LAB

[424] Eugene Field. The Love Affairs Of A Bibliomaniac, volume 443 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/lbibl10.zip.

Proudhon:1996:PM

[425] P.-J. (Pierre-Joseph) Proudhon. The Philosophy of Misery, volume 444 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/pmisr10.zip.

LeBon:1996:CSP

[426] Gustave Le Bon. The Crowd: a Study of the Popular Mind, volume 445 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/tcrwd10.zip.

Irving:1996:BRC

[427] Henry Brodribb Irving. A Book of Remarkable Criminals, volume 446 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/rcrim10.zip.

Crane:1996:MGS

[428] Stephen Crane. Maggie, Girl of the Steets, volume 447 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/mgots11.zip.

LeBon:1996:PR

[429] Gustave Le Bon. The Psychology of Revolution, volume 448 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/psrev10.zip.

Healy:1996:PLE

[430] William Healy. Pathology of Lying, Etc., volume 449 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/pathl10.zip.

Phillips:1996:SLR

[431] David Graham Phillips. Susan Lenox, Rise, volume 450 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/lenox10.zip.

Conrad:1996:SL

[432] Joseph Conrad. The Shadow Line, volume 451 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/shlin10.zip.

Borrow:1996:L

[433] George Henry Borrow. Lavengro, volume 452 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lvgro10.zip.

Ferber:1996:EMC

[434] Edna Ferber. Emma McChesney and Co., volume 453 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/emamc10.zip.

Leland:1996:BB

[435] Charles Godfrey Leland. The Breitmann Ballads, volume 454 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/britm10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext96/britm10a.zip.

Parlette:1996:UHK

[436] Ralph Parlette. The University of Hard Knocks, volume 455 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/hdknk10.zip.

Wells:1996:DWA

[437] H. G. (Herbert George) Wells. The Door in the Wall, et al., volume 456 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/tditw10.zip.

Phillips:1996:PSP

[438] David Graham Phillips. The Price She Paid, volume 457 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tpspd10.zip.

Marquis:1996:DD

[439] Don Marquis. Dreams and Dust, volume 458 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ddust10.zip.

Burnett:1996:WP

[440] Frances Hodgson Burnett. The White People, volume 459 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/twpop10.zip.

Burnett:1996:DT

[441] Frances Hodgson Burnett. The Dawn of A To-morrow, volume 460 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/tdoat10.zip.

Gallienne:1996:QGG

[442] Richard le Gallienne. Quest of the Golden Girl, volume 461 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/ggirl10.zip.

Alger:1996:EB

[443] Horatio Alger. The Errand Boy, volume 462 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rndby10.zip.

Crane:1996:RBC

[444] Stephen Crane. The Red Badge of Courage, volume 463 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/badge10a.zip.

Stevenson:1996:SSa

[445] Robert Louis Stevenson. In the South Seas, volume 464 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/sseas10.zip.

White:1996:M

[446] Stewart Edward White. The Mountains, volume 465 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tmtns10.zip.

Steinmetz:1996:GTP

[447] Andrew Steinmetz. The Gaming Table [Pros, volume 466 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/tgamt10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext96/tgamt210.zip.

Lafayette:1996:PC

[448] Marie-Madeleine Pioche de La Vergne Lafayette, Comtesse de. The Princess of Cleves, volume 467 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/pclev10.
zip.

Prevost:1996:ML

[449] abbe Prevost. Manon Lescaut, volume 468 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lesco10.zip.

Balzac:1996:DL

[450] Honoré de Balzac. The Duchesse de Langeais, volume 469 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/dlang10.zip.

Chesterton:1996:H

[451] G. K. (Gilbert Keith) Chesterton. Heretics, volume 470 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/heret11.zip.

Scott:1996:BL

[452] Sir Walter Scott. Bride of Lammermoor, volume 471 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/brlam10.zip.

Chesnutt:1996:HBC

[453] Charles Waddell Chesnutt. The House Behind The Cedars, volume 472 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/thbtc10.zip.

Bryant:1996:STC

[454] Sara Cone Bryant. Stories To Tell To Children, volume 473 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/kids110.zip.

Bryant:1996:HTC

[455] Sara Cone Bryant. How To Tell Children Stories, volume 474 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/kids210.zip.

Booth:1996:DEW

[456] General Booth. In Darkest England and The Way Out, volume 475 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/detwo10.zip.

Nasmyth:1996:JNE

[457] James Nasmyth. James Nasmyth: Engineer, An Autobiography, volume 476 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/jnasm10.zip.

Ferri:1996:CS

[458] Enrico Ferri. Criminal Sociology, volume 477 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/crsoc10.zip.

Phillips:1996:Cb

[459] David Graham Phillips. The Cost, volume 478 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tcost10.zip.

Burnett:1996:LLF

[460] Frances Hodgson Burnett. Little Lord Fauntleroy, volume 479 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/fntlr10.zip.

Hutsko:1996:UN

[461] Joe Hutsko. Undo, A Novel, volume 480 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/undo10.zip.

Michelson:1996:BC

[462] Mirriam Michelson. In The Bishop’s Carriage, volume 481 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/bicar10.zip.

Hardy:1996:W

[463] Thomas Hardy. The Woodlanders, volume 482 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/woodl10.zip.

Tarkington:1996:CC

[464] Booth Tarkington. The Conquest of Canaan, volume 483 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/ccnan10.zip.

Optic:1996:PPK

[465] Oliver Optic. Poor and Proud [Katy Redburn], volume 484 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/porap10.zip.

Baum:1996:RO

[466] L. Frank (Lyman Frank) Baum. The Road to Oz, volume 485 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/rd2oz10.zip.

Baum:1996:OO

[467] L. Frank (Lyman Frank) Baum. Ozma of Oz, volume 486 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/ozmoz10.zip.

Stevenson:1996:ST

[468] Robert Louis Stevenson. Songs of Travel, volume 487 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/strvl10.zip.

MacClure:1996:SSA

[469] Victor MacClure. She Stands Accused, volume 488 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ssacc10.zip.

Ferber:1996:OB

[470] Edna Ferber. One Basket, volume 489 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/1bskt10.zip.

Burk:1996:L

[471] Marthy Cannary Burk. Life/Adventures, volume 490 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/cjane10.zip.

Atherton:1996:R

[472] Gertrude Franklin Horn Atherton. Rezanov, volume 491 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/reznv10.zip.

Stevenson:1996:AW

[473] Robert Louis Stevenson. The Art of Writing, volume 492 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/artow10.zip.

Conrad:1996:F

[474] Joseph Conrad. Falk, volume 493 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/falk10.
zip.

Conrad:1996:T

[475] Joseph Conrad. To-morrow, volume 494 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/2mrow10.zip.

Conrad:1996:AFa

[476] Joseph Conrad. Amy Foster, volume 495 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/afost10.zip.

Mulock:1996:LLP

[477] Miss Mulock. The Little Lame Prince, volume 496 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lamep10.zip.

Coke:1996:TRS

[478] Henry J. (Henry John) Coke. Tracks of a Rolling Stone, volume 497 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/trlst10.zip.

Wiggin:1996:RSF

[479] Kate Douglas Smith Wiggin. Rebecca Of Sunnybrook Farm, volume 498 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/snybk10.zip.

Appleton:1996:TS

[480] Victor Appleton. Tom Swift, volume 499 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tom2010.zip.

Collodi:1996:AP

[481] Carlo Collodi. Adventures of Pinocchio, volume 500 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/pnoco10.zip.

Lofting:1996:SDD

[482] Hugh Lofting. The Story of Doctor Dolittle, volume 501 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/dolit10.zip.

Grey:1996:DG

[483] Zane Grey. Desert Gold, volume 502 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/dgold10.
zip.

Lang:1996:BFB

[484] Andrew Lang. The Blue Fairy Book, volume 503 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/blfry10.zip.

Sousa:1996:FS

[485] John Philip Sousa. The Fifth String, volume 504 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/strng10.zip.

White:1996:HWS

[486] Andrew Dickson White. A History of the Warfare of Science with Theology in Christendom, volume 505 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/hwswt10.zip.

Burnett:1996:S

[487] Frances Hodgson Burnett. The Shuttle, volume 506 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tshtl10.zip.

Eliot:1996:AB

[488] George Eliot. Adam Bede, volume 507 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/adamb10.zip.

Hawthorne:1996:TTT

[489] Nathaniel Hawthorne. From Twice Told Tales, volume 508 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/2tale10.zip.

LeFanu:1996:PP

[490] Joseph Sheridan Le Fanu. The Purcell Papers, volume 509–511 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pclp110.zip   (Volume 1);
ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/pclp210.
zip (Volume   2); ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/pclp310.zip   (Volume 3).

Hawthorne:1996:MOM

[491] Nathaniel Hawthorne. From Mosses From An Old Manse, volume 512 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/manse10.zip.

Hawthorne:1996:SI

[492] Nathaniel Hawthorne. From The Snow Image, volume 513 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/snowi10.zip.

Alcott:1996:LW

[493] Louisa May Alcott. Little Women, volume 514 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lwmen10.zip.

Howth:1996:ST

[494] Margret Howth. A Story of To-day, volume 515 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/mhday10.zip.

Stevenson:1996:SSb

[495] Robert Louis Stevenson. The Silverado Squatters, volume 516 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/silvs10.zip.

Baum:1996:ECO

[496] L. Frank (Lyman Frank) Baum. The Emerald City of Oz, volume 517 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/emcty10.zip.

Baum:1996:EIY

[497] L. Frank (Lyman Frank) Baum. The Enchanted Island of Yew, volume 518 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/enyew10.zip.

Baum:1996:KSC

[498] L. Frank (Lyman Frank) Baum. A Kidnapped Santa Claus, volume 519 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/kdsta10.zip.

Baum:1996:L

[499] L. Frank (Lyman Frank) Baum. Life, volume 520 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lfsta10.zip.

Defoe:1996:RCa

[500] Daniel Defoe. Robinson Crusoe, volume 521 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rbcru10.zip.

Headland:1996:CBG

[501] Isaac Taylor Headland. The Chinese Boy and Girl, volume 522 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/chnbg10.zip.

Headland:1996:CLC

[502] Isaac Taylor Headland. Court Life in China, volume 523 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/clchi10.zip.

Wells:1996:AV

[503] H. G. (Herbert George) Wells. Ann Veronica, volume 524 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/anver10.zip.

Conrad:1996:Y

[504] Joseph Conrad. Youth, volume 525 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/youth10.
zip.

Conrad:1996:HD

[505] Joseph Conrad. Heart of Darkness, volume 526 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/hdark12a.zip.

Conrad:1996:ET

[506] Joseph Conrad. End of the Tether, volume 527 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tethr10.zip.

Alger:1996:JHB

[507] Horatio Alger. Joe The Hotel Boy, volume 528 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/jothb10.zip.

Steedman:1996:KA

[508] Amy Steedman. Knights of the Art, volume 529 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/knart10.zip.

Alger:1996:DH

[509] Horatio Alger. Driven From Home, volume 530 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/drvhm10.zip.

Steinmetz:1996:GT

[510] Andrew Steinmetz. The Gaming Table, volume 531 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tgamt210.zip   (Volume 2).

Stratton-Porter:1996:AFR

[511] Gene Stratton-Porter. At the Foot of the Rainbow, volume 532 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/frain10.zip.

Stratton-Porter:1996:SC

[512] Gene Stratton-Porter. The Song of the Cardinal, volume 533 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/scard10.zip.

Stevenson:1996:IV

[513] Robert Louis Stevenson. An Inland Voyage, volume 534 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/nvoyg10.zip.

Stevenson:1996:TDC

[514] Robert Louis Stevenson. Travels with a Donkey in the Cevennes, volume 535 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/ceven10.zip.

Stevenson:1996:FH

[515] Robert Louis Stevenson. A Footnote to History, volume 536 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/fnhst10.zip.

ConanDoyle:1996:TTM

[516] Sir Arthur Conan Doyle. Tales of Terror and Mystery, volume 537 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/totam10.zip.

Bower:1996:JL

[517] B. M. Bower. Jean of the Lazy A, volume 538 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lazya10.zip.

Tuckwell:1996:BSW

[518] Rev. W. Tuckwell. Biog Study of A. W. Kinglake, volume 539 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/awkbi10.zip.

Lang:1996:RFB

[519] Andrew Lang. The Red Fairy Book, volume 540 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rdfry10.zip.

Wharton:1996:AI

[520] Edith Wharton. The Age of Innocence, volume 541 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/agino10.zip.

Johnson:1996:LM

[521] Clarence Edgar Johnson. The Life of Me, volume 542 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/lfome10.zip.

Lewis:1996:MS

[522] Sinclair Lewis. Main Street, volume 543 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/mnstr10.zip.

Montgomery:1996:AHD

[523] L. M. (Lucy Maud) Montgomery. Anne’s House of Dreams, volume 544 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/annhd10.zip.

Burroughs:1996:AEC

[524] Edgar Rice Burroughs. At the Earth’s Core, volume 545 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/atcor10.zip.

Stout:1996:UA

[525] Rex Stout. Under the Andes, volume 546 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/andes10.zip.

Gaboriau:1996:BTV

[526] Emile Gaboriau. Baron Trigault’s Vengeance, volume 547 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/trvng10.zip.

Anonymous:1996:PTO

[527] Anonymous. Project Trinity, Official U.S. Government Report, volume 548 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/prjtr10.zip.

Azuela:1996:U

[528] Mariano Azuela. The Underdogs, volume 549 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ndrdg10.zip.

Eliot:1996:SMW

[529] George Eliot. Silas Marner, The Weaver of Raveloe, volume 550 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/smarn10.zip.

Burroughs:1996:LTF

[530] Edgar Rice Burroughs. The Land That Time Forgot, volume 551 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/tlttf10.zip.

Burroughs:1996:PT

[531] Edgar Rice Burroughs. People Out Of Time, volume 552 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/poftm10.zip.

Burroughs:1996:TA

[532] Edgar Rice Burroughs. Out Of Time’s Abyss, volume 553 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/ootma10.zip.

Tyler:1996:C

[533] Royall Tyler. The Contrast, volume 554 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tcntr10.zip.

Saki:1996:UB

[534] Saki. The Unbearable Bassington, volume 555 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/nbrbl10.zip.

Kipling:1996:RF

[535] Rudyard Kipling. Rewards and Fairies, volume 556 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rwfrs10.zip.

Kipling:1996:PPH

[536] Rudyard Kipling. Puck of Pook’s Hill, volume 557 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/pkpkh10.zip.

Buchan:1996:TS

[537] John Buchan. The Thirty-Nine Steps, volume 558 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/39stp10.zip.

Buchan:1996:G

[538] John Buchan. Greenmantle, volume 559 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/gmntl10.zip.

Buchan:1996:MS

[539] John Buchan. Mr. Standfast, volume 560 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/mstnf10.zip.

Defoe:1996:RCb

[540] Daniel Defoe. Robinson Crusoe, volume 561 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rbcr210.zip.

Kay:1996:GAB

[541] Ross Kay. Go Ahead Boys and Racing Motorboat, volume 562 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/gabrm10.zip.

Urides:1996:PMI

[542] “Eros Urides”. The Planet Mars and Its Inhabitants, volume 563 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/marsi10.zip.

Dickens:1996:MED

[543] Charles Dickens. The Mystery of Edwin Drood, volume 564 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/drood10.zip.

Borrow:1996:ZGS

[544] George Henry Borrow. Zincali, Gypsies of Spain, volume 565 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/zncli10.zip.

Schiller:1996:HTY

[545] Johann Christoph Friedrich von Schiller. The History of the Thirty Years’ War, volume 566 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/1jcfs10.zip.

Brann:1996:BI

[546] William Cowper Brann. Brann The Iconoclast, volume 567–569 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/bti0110.zip   (Vol.1); ftp:
//uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/bti1010.
zip   (Vol.10);   ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/bti1210.zip   (Vol.12).

Fries:1996:MG

[547] Adelaide L. (Adelaide Lisetta) Fries. The Moravians in Georgia, volume 570 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/mrvga10.zip.

Anonymous:1996:CWF

[548] Anonymous. The 1995 CIA World Factbook, volume 571 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/world95.zip.

Potter:1996:GBT

[549] Beatrix Potter. The Great Big Treasury of Beatrix Potter, volume 572 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/gbtbp10.zip.

Lamb:1996:TS

[550] Charles Lamb. Tales of Shakespeare, volume 573 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tshak10.zip.

Blake:1996:PWB

[551] William Blake. Poems of William Blake, volume 574 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/pblak10.zip.

Bacon:1996:EFB

[552] Francis Bacon. Essays of Francis Bacon, volume 575 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/ebacn10.zip.

Anonymous:1996:PGW

[553] Anonymous. The Project Gutenberg Web Pages, volume 576 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/pg.zip.

Johnson:1996:WSJ

[554] Samuel Johnson. The Works of Samuel Johnson, Volume 4 of 16, volume 577 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/sjv0410.zip.

Nakashima:1996:C

[555] Tadashi Nakashima. Down With The Cities, volume 578 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/dwtct10.zip.

Lanier:1996:PSL

[556] Sidney Lanier. The Poems of Sidney Lanier, volume 579 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/slanr10.zip.

Dickens:1996:PP

[557] Charles Dickens. The Pickwick Papers, volume 580 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/pwprs10.zip.

Jenkins:1996:GBS

[558] Edward Jenkins. Ginx’s Baby, A Satire, volume 581 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ginxb10.zip.

Potter:1996:CBP

[559] Beatrix Potter. A Collection of Beatrix Potter Stories, volume 582 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/bpsto10.zip.

Collins:1996:WW

[560] Wilkie Collins. The Woman in White, volume 583 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/wwhit10.zip.

Wilson:1996:ON

[561] Harriet E. Wilson. Our Nig, volume 584 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ourng10.zip.

Craft:1996:RTM

[562] William Craft. Running a Thousand Miles for Freedom, volume 585 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/runng10.zip.

Browne:1996:RMH

[563] Sir Thomas Browne. Religio Medici, Hydriotaphia, and the Letter to a Friend, volume 586 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/rmedi10.zip.

Marquis:1996:DOS

[564] Don Marquis. Danny’s Own Story, volume 587 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/dsown10.zip.

Dickens:1996:MHC

[565] Charles Dickens. Master Humphrey’s Clock, volume 588 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/mhmph10.zip.

Stevenson:1996:CK

[566] Robert Louis Stevenson. Catriona (Kidnapped2), volume 589 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/ctrna10.zip.

Japp:1996:RLS

[567] Alexander H. (Alexander Hay) Japp. Robert Louis Stevenson, A Memorial, volume 590 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/rlsjp10.zip.

Teasdale:1996:FS

[568] Sara Teasdale. Flame and Shadow, volume 591 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/fshad10.zip.

Lindsay:1996:CNO

[569] Vachel Lindsay. The Chinese Nightingale and Other Poems, volume 592 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/ngale10.zip.

Maupassant:1996:SWG

[570] Guy de (Henri Rene Albert) Maupassant. Selected Writings of Guy De Maupassant V. 1, volume 593 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/swgem10.
zip.

Anonymous:1996:TS

[571] Anonymous. Twilight Stories, volume 594 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/twils10.zip.

Aldrich:1996:ST

[572] Thomas Bailey Aldrich. The Sisters’ Tragedy, volume 595 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/sistr10.zip.

Teasdale:1996:RS

[573] Sara Teasdale. Rivers to the Sea, volume 596 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rivse10.zip.

Anonymous:1996:NSS

[574] Anonymous. Njal’s Saga The Story of Burnt Njal, Icelandic, volume 597 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/njals10.zip.

Anonymous:1996:HH

[575] Anonymous. Heimskringla Heimskringla, volume 598 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/hmskr10.zip.

Thackeray:1996:VF

[576] William Makepeace Thackeray. Vanity Fair, volume 599 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/vfair10.zip.

Dostoyevsky:1996:NU

[577] Fyodor Dostoyevsky. Notes From The Underground, volume 600 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/notun11.zip.

Lewis:1996:M

[578] Matthew Lewis. The Monk, volume 601 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tmonk10.zip.

Lucanus:1996:PCW

[579] Marcus Annaeus Lucanus (39-65 AD). Pharsalia [Civil War], volume 602 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/pcwar10.zip.

Borrow:1996:LGB

[580] George Henry Borrow. Letters of George Borrow, volume 603 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/ltbor10.zip.

Arnold:1996:GM

[581] Edwin Lester Linden Arnold. Gulliver of Mars, volume 604 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/gulvm10.zip.

Burroughs:1996:P

[582] Edgar Rice Burroughs. Pellucidar, volume 605 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/pellu10.zip.

Linderman:1996:IWS

[583] Frank Bird Linderman. Indian Why Stories, volume 606 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/inwhy10.zip.

Anonymous:1996:ACA

[584] Anonymous. The Augsburg Confession, The Augsburg Confession, volume 607 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/auglg10.zip.

Milton:1996:A

[585] John Milton. Aeropagitica, volume 608 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/areop10.zip.

Tingfang:1996:ASO

[586] Tingfang. America As Seen By Oriental Diplomat, volume 609 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/asbod10.zip.

Tennyson:1996:IK

[587] Alfred Tennyson. Idylls of the King, volume 610 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/idyll10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext96/idyll10a.zip.

Buchan:1996:PJ

[588] John Buchan. Prester John, volume 611 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/prsjn10.zip.

Anonymous:1996:CJ

[589] Anonymous. The Constitution of Japan, volume 612 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/jcnst10.zip.

Anonymous:1996:PCJ

[590] Anonymous. Previous Constitution of Japan, volume 613 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/cjold10.zip.

Stevenson:1996:AP

[591] Robert Louis Stevenson. Across The Plains, volume 614 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/axpln10.zip.

Ariosto:1996:OF

[592] Ludovico Ariosto. Orlando Furioso, volume 615 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/orfur10.zip.

Stevenson:1996:VPS

[593] Robert Louis Stevenson. Vailima Prayers and Sabbath Morn, volume 616 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/vpasm10.zip.

Seeger:1996:P

[594] Alan Seeger. Poems, volume 617 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/seegr10.
zip.

Anonymous:1996:CJB

[595] Anonymous. Codex Junius 11, Biblical Anglo-Saxon Translations, volume 618 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/codju10.zip.

Trollope:1996:W

[596] Anthony Trollope. The Warden, volume 619 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/twrdn10.zip.

Carroll:1996:SB

[597] Lewis Carroll. Sylvie and Bruno, volume 620 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/sbrun10.zip.

James:1996:VRE

[598] William James. Varieties of Religious Experience: . A Study in Human Nature, volume 621 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/varre10.zip.

Stevenson:1996:LRLa

[599] Robert Louis Stevenson. Letters of Robert Louis Stevenson Vol 1, volume 622 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/rlsl110.zip.

Swift:1996:BBO

[600] Jonathan Swift. The Battle of the Books And Other Short Pieces, volume 623 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/batbk10.zip. Preface – I. THE BATTLE

OF THE BOOKS – II. A MEDITATION UPON A BROOMSTICK. – III. PREDICTIONS FOR THE YEAR 1708. – IV. THE ACCOMPLISH-MENT OF THE FIRST OF MR. BICKERSTAFF’S PREDICTIONS. – V. BAUCIS AND PHILEMON. – VI. THE LOGICIANS REFUTED.

– VII. THE PUPPET SHOW. – VIII. CADENUS AND VANESSA. – IX. STELLA’S BIRTHDAYS – X. TO STELLA – XI. THE FIRST HE WROTE OCT. 17, 1727. – XII. THE SECOND PRAYER WAS WRIT-TEN NOV. 6, 1727. – XIII. THE BEASTS’ CONFESSION (1732). – XIV. ABOLISHING CHRISTIANITY – XV. HINTS TOWARDS AN ESSAY ON CONVERSATION. – XVI. THOUGHTS ON VARIOUS SUBJECTS.

Bellamy:1996:LB

[601] Edward Bellamy. Looking Backward, 2000 to 1887, volume 624 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/lkbak10.zip.

Aldrich:1996:PP

[602] Thomas Bailey Aldrich. Ponkapog Papers, volume 625 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/pnkpg10.zip.

Bury:1996:PRB

[603] Richard de Bury. The Philobiblon of Richard de Bury, volume 626 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/phlbb10.zip.

Stevenson:1996:ET

[604] Robert Louis Stevenson. Essays of Travel, volume 627 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/esstr10.zip.

Anonymous:1996:SROa

[605] Anonymous. The Square Root of 3 to one million digits, volume 628 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/3sqrt10.zip.

Anonymous:1996:SROb

[606] Anonymous. The Square Root of 5 to one million digits, volume 629 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/5sqrt10.zip.

Anonymous:1996:SROc

[607] Anonymous. The Square Root of 6 to one million digits, volume 630 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/6sqrt10.zip.

Anonymous:1996:SROd

[608] Anonymous. The Square Root of 7 to one million digits, volume 631 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/7sqrt10.zip.

Anonymous:1996:SROe

[609] Anonymous. The Square Root of 8 to one million digits, volume 632 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/8sqrt10.zip.

Anonymous:1996:GMR

[610] Anonymous. The Golden Mean or Ratio, volume 633 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/gmean10.zip.

Anonymous:1996:MMC

[611] Anonymous. Miscellaneous Mathematical Constants Etc, volume 634 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/miscc10.zip.

Anonymous:1996:SROf

[612] Anonymous. The Square Root of 10 to one million digits, volume 635 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/10srt10.zip.

MacKay:1996:MPDa

[613] Charles MacKay. Memoirs of Popular Delusions V1, volume 636 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/ppdel10.zip.

Stevenson:1996:LRLb

[614] Robert Louis Stevenson. Letters of Robert Louis Stevenson Vol 2, volume 637 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/rlsl210.zip.

Conrad:1996:OI

[615] Joseph Conrad. An Outcast of the Islands, volume 638 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/outci10.zip.

Gamble:1996:GA

[616] Eliza Burt Gamble. God-Idea of the Ancients, volume 639 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/gidan10.zip.

Lang:1996:YFB

[617] Andrew Lang. The Yellow Fairy Book, volume 640 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ylfry10.zip.

Lang:1996:VFB

[618] Andrew Lang. The Violet Fairy Book, volume 641 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/vifry10.zip.

James:1996:AD

[619] Henry James. The Altar of the Dead, volume 642 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/altdd10.zip.

James:1996:DL

[620] Henry James. The Death of the Lion, volume 643 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/dlion10.zip.

Dickens:1996:HMG

[621] Charles Dickens. The Haunted Man and the Ghost’s Bargain, volume 644 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/hntmn10.zip.

James:1996:FC

[622] Henry James. The Figure in the Carpet, volume 645 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/fgcpt10.zip.

Ballantyne:1996:CI

[623] R. M. (Robert Michael) Ballantyne. The Coral Island, volume 646 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/coril10.zip.

Stevenson:1996:D

[624] Robert Louis Stevenson. The Dynamiter, volume 647 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/dynmt10.zip.

Borrow:1996:WW

[625] George Henry Borrow. Wild Wales, volume 648 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/wwals10.zip.

Bell:1996:APB

[626] Robert Bell. Ancient Poems, Ballads and Songs of the Peasantry of England, volume 649 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/oleng10.zip.

Dickens:1996:PI

[627] Charles Dickens. Pictures From Italy, volume 650 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/picit10.zip.

Carroll:1996:POP

[628] Lewis Carroll. Phantasmagoria and Other Poems, volume 651 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/fntsm10.zip.

Johnson:1996:RPA

[629] Samuel Johnson. Rasselas, Prince of Abyssinia, volume 652 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/rslas10.zip.

Dickens:1996:C

[630] Charles Dickens. The Chimes, volume 653 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tchms11.zip.

Bunyan:1996:GAC

[631] John Bunyan. Grace Abounding to Chief of Sinners, volume 654 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/gacos10.zip.

Orr:1996:LLR

[632] Sutherland Orr, Mrs. Life and Letters of Robert Browning, volume 655 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/orrbr10.zip.

Sharp:1996:LRB

[633] William Sharp. Life of Robert Browning, volume 656 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/shabr10.zip.

Ingram:1996:AC

[634] J. H. (James Henry) Ingram. The Anglo-Saxon Chronicle, volume 657 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/angsx10.zip.

Smyrnaeus:1996:FT

[635] Quintus Smyrnaeus. The Fall of Troy, volume 658 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ftroy10.zip.

Alger:1996:PP

[636] Horatio Alger. Paul The Peddler, volume 659 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ptped10.zip.

Anonymous:1996:GWUa

[637] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [AB], volume 660 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgwab04.zip.

Anonymous:1996:GWUb

[638] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [C], volume 661 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgwc04.zip.

Anonymous:1996:GWUc

[639] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [DE], volume 662 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgwde04.zip.

Anonymous:1996:GWUd

[640] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [FH], volume 663 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgwfh04.zip.

Anonymous:1996:GWUe

[641] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [IL], volume 664 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgwil04.zip.

Anonymous:1996:GWUf

[642] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [MO], volume 665 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/pgwmo04.zip.

Anonymous:1996:GWUg

[643] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [PQ], volume 666 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgwpq04.zip.

Anonymous:1996:GWUh

[644] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [R], volume 667 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgwr04.zip.

Anonymous:1996:GWUi

[645] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [S], volume 668 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgws04.zip.

Anonymous:1996:GWUj

[646] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [TW], volume 669 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/pgwtw04.zip.

Anonymous:1996:GWUk

[647] Anonymous. The Gutenberg Webster’s Unabridged Dictionary [XZ], volume 670 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext96/pgwxz04.zip.

Alger:1996:PF

[648] Horatio Alger. Phil, the Fiddler, volume 671 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/phidl10.zip.

Walter:1996:SGC

[649] Russ Walter. The Secret Guide to Computers, volume 672 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/sgcpv10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext96/sgcwp10.zip.

Anonymous:1996:GWUl

[650] Anonymous. Gutenberg Webster’s Unabridged Dictionary, volume 673 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/pgwht04.zip.

Clough:1996:PL

[651] Arthur Hugh Clough. Plutarch’s Lives, volume 674 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/plivs10.zip.

Dickens:1996:AN

[652] Charles Dickens. American Notes, volume 675 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/amnts10.zip.

Dickens:1996:BL

[653] Charles Dickens. The Battle of Life, volume 676 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/batlf10.zip.

Kingsley:1996:H

[654] Charles Kingsley. Heroes, volume 677 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/ghros10.
zip.

Dickens:1996:CH

[655] Charles Dickens. The Cricket on the Hearth, volume 678 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tcoth11.zip.

Harper:1996:P

[656] Frances Ellen Watkins Harper. Poems, volume 679 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/pfewh09.zip.

Naidu:1996:GT

[657] Sarojini Naidu. The Golden Threshold, volume 680 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/gldth10.zip.

Gorky:1996:COW

[658] Maxim Gorky. Creatures That Once Were Men, volume 681 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/crmen10.zip.

Anonymous:1996:CCR

[659] Anonymous. Catalan’s Constant [Ramanujan’s Formula], volume 682 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/ctcst10.zip.

Walton:1996:CA

[660] Izaak Walton. The Compleat Angler, volume 683 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tcang10.zip.

US:1996:WEN

[661] U.S. Arms Control and Disarmament Agency. Worldwide Effects of Nuclear War-Some Perspectives, volume 684 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/nukwr10.zip.

Anonymous:1996:ABH

[662] Anonymous. The Atomic Bombings of Hiroshima and Nagasaki, volume 685 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/abomb10.zip.

Anonymous:1996:TEU

[663] Anonymous. The Treaty of the European Union [Maastricht], volume 686 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/maast10.zip.

Conrad:1996:PR

[664] Joseph Conrad. A Personal Record, volume 687 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/aprjc10.zip.

Dunbar:1996:GSR

[665] Alice Ruth Moore Dunbar. The Goodness of St. Rocque and other stories, volume 688 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/stroq10.zip.

Tolstoy:1996:KSO

[666] Count Leo Nikoleyevich Tolstoy. The Kreutzer Sonata and other stories, volume 689 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/krson10.

zip. Kreutzer Sonata – Ivan the Fool – A Lost Opportunity – Polikushka – The Candle..

Russell:1996:PRF

[667] Bertrand Arthur William 3rd Russell, Earl. Proposed Roads to Freedom, volume 690 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/rfree10.zip.

Riley:1996:CWJa

[668] James Whitcomb Riley. Complete Works of James Whitcomb Riley, Vol I, volume 691 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/01jwr10.zip.

Riley:1996:CWJb

[669] James Whitcomb Riley. Complete Works of James Whitcomb Riley, Vol X, volume 692 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/10jwr10.zip.

Mitchell:1996:AQ

[670] S. Weir (Silar Weir) Mitchell. The Autobiography of a Quack, volume 693 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/auqak10.zip.

Anonymous:1996:SEM

[671] Anonymous. Stories from Everybody’s Magazine, 1910, volume 694 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/10evm10.zip.

Kingsley:1996:G

[672] Charles Kingsley. Glaucus, volume 695 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/glcus10.zip.

Walpole:1996:CO

[673] Horace Walpole. The Castle of Otranto, volume 696 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/cotrt10.zip.

MacDonald:1996:LP

[674] George MacDonald. The Light Princess, volume 697 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/ltprn10.zip.

Stevenson:1996:MFJ

[675] Robert Louis Stevenson. Memoir of Fleeming Jenkin, volume 698 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/fleem10.zip.

Dickens:1996:CHE

[676] Charles Dickens. A Child’s History of England, volume 699 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/achoe10.zip.

Dickens:1996:OCS

[677] Charles Dickens. The Old Curiosity Shop, volume 700 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/curio10.zip.

Ruskin:1996:KGR

[678] John Ruskin. The King of the Golden River, volume 701 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tkogr10.zip.

Young:1996:SLG

[679] Martha Young. Somebody’s Little Girl, volume 702 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/slgrl10.zip.

Lovelace:1996:LP

[680] Richard Lovelace. The Lucasta Poems, volume 703 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lcsta10.zip.

Dyke:1996:M

[681] Henry van Dyke. The Mansion, volume 704 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/tmans10.zip.

Barber:1996:R

[682] Margaret Fairless Barber. The Roadmender, volume 705 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/rmend10.zip.

Hornung:1996:AC

[683] E. W. (Ernest William) Hornung. The Amateur Cracksman, volume 706 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/amatc10.zip.

Hornung:1996:RFA

[684] E. W. (Ernest William) Hornung. Raffles, Further Adventures, volume 707 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/rafls10.zip.

MacDonald:1996:PG

[685] George MacDonald. The Princess and the Goblin, volume 708 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/prgob10.zip.

MacDonald:1996:PC

[686] George MacDonald. The Princess and Curdie, volume 709 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/prcur10.zip.

London:1996:LLO

[687] Jack London. Love of Life and other stories, volume 710 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/llife10.zip.

Haggard:1996:AQ

[688] H. Rider Haggard. Allan Quatermain, volume 711 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/allnq10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext96/allnq10a.zip.

Ellis:1996:TJ

[689] Edward Sylvester Ellis. Thomas Jefferson, volume 712 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tjeff10.zip.

MacKay:1996:MPDb

[690] Charles MacKay. Memoirs of Popular Delusions V2, volume 713 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/2ppdl10.zip.

Hope:1996:BTC

[691] Laura Lee Hope. Bobbsey Twins in the Country, volume 714 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tbtic10.zip.

Buchan:1996:MET

[692] John Buchan. Moon Endureth [Tales, volume 715 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ndrth10.zip.

Marquis:1996:CJB

[693] Don Marquis. The Cruise of the Jasper B., volume 716 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/jsprb10.zip.

Hearn:1996:CML

[694] Lafcadio Hearn. Chita: A Memory of Last Island, volume 717 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/chita10.zip.

Wells:1996:TB

[695] H. G. (Herbert George) Wells. Tono Bungay, volume 718 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tonob10.zip.

Henley:1996:PWH

[696] William Ernest Henley. Plays of William E. Henley and Robert Louis Stevenson, volume 719 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/tpohs10.
zip.

Conrad:1996:AFb

[697] Joseph Conrad. Almayer’s Folly, volume 720 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/lmyer10.zip.

Wiggin:1996:BCC

[698] Kate Douglas Smith Wiggin. The Birds’ Christmas Carol, volume 721 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/tbscc10.zip.

Ridpath:1996:JOP

[699] John Clark Ridpath. James Otis The Pre-Revolutionist, volume 722 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/jotis10.zip.

Howells:1996:HJJ

[700] William Dean Howells. Henry James, Jr., volume 723 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/jimjr10.zip.

Howells:1996:MLM

[701] William Dean Howells. The Man of Letters as a Man of Business, volume 724 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/tmlmb10.zip.

Smiles:1996:MII

[702] Samuel Smiles. Men of Invention and Industry, volume 725 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/moiai10.zip.

Howells:1996:PCR

[703] William Dean Howells. A Psychological Counter-Current in recent fiction, volume 726 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/pccmf10.zip.

Carpenter:1996:SB

[704] John Carpenter. The Star-Spangled Banner, volume 727 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996.

Howells:1996:EZ

[705] William Dean Howells. Emile Zola, volume 728 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/ezola10.zip.

Levy:1996:HCR

[706] Steven Levy. Hackers, Computer Revolution Heroes, volume 729 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/hckrs10.zip.

Dickens:1996:OT

[707] Charles Dickens. Oliver Twist, volume 730 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/olivr10.zip.

Gibbon:1996:HDF

[708] Edward Gibbon. The History of the Decline and Fall Of The Roman Empire, volume 731–736 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/1dfre10.zip   (V1)
; ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/
2dfre10.zip   (V2); ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/3dfre10.zip   (V3); ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/etext96/4dfre10.zip   (V4); ftp://uiarchive.
cso.uiuc.edu/pub/etext/gutenberg/etext96/5dfre10.zip   (V5)
; ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext96/
6dfre10.zip   (V6).

Hope:1996:BTA

[709] Laura Lee Hope. The Bobbsey Twins at School, volume 737 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/tbtas10.zip.

Lang:1996:PDL

[710] Andrew Lang. The Puzzle of Dickens’s Last Plot, volume 738 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/pldlp10.zip.

Anonymous:1996:HCR

[711] Anonymous. Henry Clay’s Remarks in House and Senate, volume 739 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/hcrhs10.zip.

Anonymous:1996:JCC

[712] Anonymous. John C. Calhoun’s Remarks in the Senate, volume 740 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/jccrs10.zip.

Anonymous:1996:THB

[713] Anonymous. Thomas Hart Benton’s Remarks to the Senate, volume 741 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/thbrs10.zip.

Brisbane:1996:EHN

[714] Arthur Brisbane. Editorials from the Hearst Newspapers, volume 742 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/ehnab10.zip.

Godwin:1996:TMH

[715] William Godwin. Thoughts on Man, His Nature, etc., volume 743 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/tmnwg10.zip.

Anonymous:1996:GMM

[716] Anonymous. The Golden Mean, To A Million Digits, volume 744 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/gmean11.zip.

Anonymous:1996:ODP

[717] Anonymous. One Divided by Pi, To A Million Digits, volume 745 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/onepi10.zip.

London:1996:BD

[718] Jack London. Burning Daylight, volume 746 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/bdlit10.zip.

Gould:1996:ACM

[719] George M. (George Milbry) Gould. Anomalies and Curiosities of Medicine, volume 747 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/aacom10.zip.

Yates:1996:BD

[720] Dornford Yates. The Brother of Daphne, volume 748 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/bdaph10.zip.

John:1996:BI

[721] Saint John of Damascus. Barlaam and Ioasaph, volume 749 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/

gutenberg/etext96/bioas10.zip. The text of this edition is based on that published as ST. JOHN DAMASCENE: BARLAAM AND IOAS-APH (Trans: G. R. Woodward and H. Mattingly; Harvard University Press, Cambridge MA, 1914)..

Anonymous:1996:HHH

[722] Anonymous. The High History of the Holy Graal, Author Unknown, volume 750 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext96/hhohg10.zip.

Holmes:1996:ABT

[723] Oliver Wendell Holmes. Autocrat of the Breakfast Table, volume 751 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/aofbt10.zip.

Freud:1996:YGD

[724] Sigmund Freud, 1856-1939. A Young Girl’s Diary, and Letter of Sigmund Freud, volume 752 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext96/ygdsf10.zip.

White:1996:AN

[725] Stewart Edward White. Arizona Nights, volume 753 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/aznit10.zip.

Loon:1996:SM

[726] Hendrik van Loon. The Story of Mankind, volume 754 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/hmank10.zip.

Bird:1996:LLR

[727] Isabella L. (Isabella Lucy) Bird. Lady’s Life in Rocky Mountains, volume 755 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/llirm10.zip.

Munk:1996:AS

[728] J. A. (Joseph Amasa) Munk. Arizona Sketches, volume 756 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/azsks10.zip.

Gilbert:1996:BB

[729] W. S. (William Schwenck) Gilbert. 50 Bab Ballads, volume 757 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/50bab10.zip.

Anonymous:1996:LPW

[730] Anonymous. LandSat Picture of Washington, DC, USA, volume 758 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext96/wdcls10.zip.

Beerbohm:1996:JP

[731] Sir Max Beerbohm. James Pethel, volume 759 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/pethl10.zip.

Beerbohm:1996:ES

[732] Sir Max Beerbohm. Enoch Soames, volume 760 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/enoch11.zip.

Beerbohm:1996:VL

[733] Sir Max Beerbohm. A. V. Laider, volume 761 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/laidr10.zip.

Whale:1996:BAP

[734] George Whale. British Airships, Past/Present, volume 762 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/bappf10.zip.

Murray:1996:R

[735] John Murray. The Round-Up, volume 763 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/rndup10.zip.

Dodge:1996:HBS

[736] Mary Mapes Dodge. Hans Brinker/Silver Skates, volume 764 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext96/hboss10.zip.

Merritt:1996:MP

[737] Abraham Merritt. The Moon Pool, volume 765 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/mpool10.zip.

Dickens:1996:DC

[738] Charles Dickens. David Copperfield, volume 766 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/cprfd10.zip.

Bronte:1996:AG

[739] Anne Brontë. Agnes Grey, volume 767 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/agnsg10.zip.

Bronte:1996:WH

[740] Emily Brontë. Wuthering Heights, volume 768 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext96/wuthr10.zip.

Okakura:1997:BT

[741] Kakuzo Okakura. The Book of Tea, volume 769 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/tboft10.zip.

Nesbit:1997:STS

[742] E. (Edith) Nesbit. The Story of the Treasure Seekers, volume 770 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/tsots10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext97/tsots10a.zip.

Bronte:1997:TSP

[743] Charlotte Brontë. Two Short Pieces by Charlotte Brontë: Biographical Notice of Ellis and Acton Bell; Preface to ’Wuthering Heights’, volume 771 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/brnte10.zip.

Stevenson:1997:ME

[744] Robert Louis Stevenson. Moral Emblems, volume 772 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/morem10.zip.

Wilde:1997:LAS

[745] Oscar Wilde. Lord Arthur Savile’s Crime and other stories, volume 773 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/ldasc10.zip.

Wilde:1997:EL

[746] Oscar Wilde. Essays and Lectures, volume 774 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/sandl10.zip.

Wells:1997:WSW

[747] H. G. (Herbert George) Wells. When the Sleeper Wakes, volume 775 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/wtslw10.zip.

Marquis:1997:HGT

[748] Don Marquis. Hermione’s Group of Thinkers, volume 776 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/hlgst11.zip.

Claxton:1997:MA

[749] William J. Claxton. The Mastery of the Air, volume 777 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tmota10.zip.

Nesbit:1997:FCI

[750] E. (Edith) Nesbit. Five Children and It, volume 778 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/fivit10.zip.

Marlowe:1997:DFa

[751] Christopher Marlowe. Dr. Faustus, volume 779 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/drfst10.zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/drfsta10.zip;   ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/etext97/drfst10a.zip   (without footnotes)

.

Wells:1997:WA

[752] H. G. (Herbert George) Wells. The War in the Air, volume 780 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/wrair10.zip.

Marshall:1997:STG

[753] Logan Marshall. Sinking of the Titanic and Great Sea Disasters, volume 781 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/ttnic10.zip. A Detailed and Accurate Ac-

count of the Most Awful Marine Disaster in History, Constructed from the Real Facts as Obtained from Those on Board Who Survived.

Anonymous:1997:TSJ

[754] Anonymous. The Travels of Sir John Mandeville, Author Unknown, volume 782 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/tosjm10.zip.

Badger:1997:LC

[755] Joseph E. Badger. The Lost City, volume 783 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/locty10.zip.

Boyesen:1997:BN

[756] Hjalmar Hjorth Boyesen. Boyhood in Norway, volume 784 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/bnrwy10.zip.

Lucretius:1997:NT

[757] Lucretius. Nature of Things, volume 785 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/natng10.zip.

Dickens:1997:HT

[758] Charles Dickens. Hard Times, volume 786 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/hardt10.zip.

Barr:1997:MB

[759] Amelia E. Barr, Edith Huddleston. The Man Between, volume 787 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/mnbtw10.zip.

London:1997:RO

[760] Jack London. The Red One, volume 788 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/tred110.zip.

Fairless:1997:GBH

[761] Michael Fairless. Gathering of Brother Hilarius, volume 789 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tgobh10.zip.

Wilde:1997:LWF

[762] Oscar Wilde. Lady Windermere’s Fan, volume 790 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/lwfan10.zip.

Tennyson:1997:P

[763] Alfred Tennyson. The Princess, volume 791 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/prncs09.zip.

Brown:1997:WT

[764] Charles Brockden Brown. Wieland, The Transformation, volume 792 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/welnd10.zip.

Talbot:1997:ADW

[765] Frederick Arthur Ambrose Talbot. Aeroplanes and Dirigibles of War, volume 793 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/aadow10.zip.

Nesbit:1997:W

[766] E. (Edith) Nesbit. The Wouldbegoods, volume 794 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/twbgd10.zip.

Lang:1997:BLP

[767] Andrew Lang. Ballads Lyrics and Poems of Old France, volume 795 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/blpof10.zip.

Stendhal:1997:CP

[768] (Henri Beyle) Stendhal. La Chartreuse de Parme, volume 796 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/8parm07.zip.

Stendhal:1997:LCE

[769] (Henri Beyle) Stendhal. L’Abbesse de Castro etc., volume 797 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/8cast07.zip.

Stendhal:1997:RN

[770] (Henri Beyle) Stendhal. Le Rouge et Le Noir, volume 798 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/8roug07.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext97/8roug08.zip.

Verne:1997:TL

[771] Jules Verne. De La Terre à La Lune, volume 799 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/8lune07.zip.

Verne:1997:TDM

[772] Jules Verne. Tour Du Mond 80 Jours, volume 800 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/880jr07.zip.

Stendhal:1997:C

[773] (Henri Beyle) Stendhal. Les Cenci, volume 801 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/8cnci07.zip.

Stendhal:1997:VA

[774] (Henri Beyle) Stendhal. Vittoria Accoramboni, volume 802 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/8vtra07.zip.

Stendhal:1997:DP

[775] (Henri Beyle) Stendhal. La Duchesse de Palliano, volume 803 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/8plno07.zip.

Sterne:1997:SJ

[776] Laurence Sterne. A Sentimental Journey, volume 804 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/senjr10.zip.

Fitzgerald:1997:SP

[777] F. Scott (Francis Scott) Fitzgerald. This Side of Paradise, volume 805 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/tsopr09.zip.

Sophocles:1997:P

[778] Sophocles. Philoktetes, volume 806 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/phlok10.
zip.

Dickens:1997:H

[779] Charles Dickens. Hunted Down, volume 807 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/hntdn10.zip.

Gilbert:1997:CPG

[780] W. S. (William Schwenck) Gilbert. The Complete Plays of Gilbert and Sullivan, volume 808 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext97/cpogs10.zip.

Dickens:1997:HR

[781] Charles Dickens. Holiday Romance, volume 809 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/hldrm10.zip.

Dickens:1997:GSE

[782] Charles Dickens. George Silverman’s Explanation, volume 810 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/gsilx10.zip.

Marlowe:1997:DFb

[783] Christopher Marlowe. Dr. Faustus, volume 811 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/drfsta10.zip.

Anonymous:1997:CCP

[784] Anonymous. Catalan’s Constant to 1.5M Places, volume 812 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tspar10.zip.

Tolstoi:1997:RT

[785] Ilia Lvovich Tolstoi, Graf. Reminiscences of Tolstoy, volume 813 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/rtlst10.zip.

Trollope:1997:HS

[786] Anthony Trollope. Hunting Sketches, volume 814 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/hntsk10.zip.

Toqueville:1997:DA

[787] Alexis de Toqueville. Democracy In America, volume 815–816 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/1dina10.zip   (Vol 1); ftp:
//uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/2dina10.
zip   (Vol 2).

Anonymous:1997:JF

[788] Anonymous. The Jargon File, 4.0.0, volume 817 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/jarg400.zip.

Barber:1997:AS

[789] H. (Horatio) Barber. The Aeroplane Speaks, volume 818 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/arspk10.zip.

Casson:1997:HT

[790] Herbert Newton Casson. The History of the Telephone, volume 819 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/thott10.zip.

Dyer:1997:EHL

[791] Frank Lewis Dyer. Edison, His Life and Inventions, volume 820 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ehlai10.zip.

Dickens:1997:DS

[792] Charles Dickens. Dombey and Son, volume 821 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/domby10.zip.

Tymon:1997:TE

[793] Frank Tymon. The Tarn of Eternity, volume 822 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/etarn10.zip.

Anonymous:1997:LSD

[794] Anonymous. The Life of St. Declan of Ardmore, volume 823 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/stdec10.zip.

Dickens:1997:SLS

[795] Charles Dickens. Speeches: Literary and Social, volume 824 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dslas10.zip.

Anonymous:1997:LVB

[796] Anonymous. Latin Vulgate Bible: Book Titles and Names, volume 825 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/lvbtn10.zip.

Anonymous:1997:LVEa

[797] Anonymous. Latin Vulgate, Esther: Liber Esther, volume 826 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/lvest10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext97/lvest10a.zip.

Anonymous:1997:LVEb

[798] Anonymous. Latin Vulgate, Esther, volume 827 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/lvest10a.zip.

Anonymous:1997:LVD

[799] Anonymous. Latin Vulgate, Daniel: Prophetia Danielis, volume 828 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/lvdan10.zip.

Swift:1997:GT

[800] Jonathan Swift. Gulliver’s Travels, volume 829 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/gltrv10.zip.

Rhodius:1997:A

[801] Apollonius Rhodius. The Argonautica, volume 830 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1argn10.zip.

DeTroyes:1997:FAR

[802] Chretien DeTroyes. Four Arthurian Romances, volume 831 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dtroy10.zip.

McSpadden:1997:RH

[803] J. Walker (Joseph Walker) McSpadden. Robin Hood, volume 832 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1rbnh10.zip.

Veblen:1997:TLC

[804] Thorstein Veblen. Theory of the Leisure Class, volume 833 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/totlc10.zip.

ConanDoyle:1997:MSH

[805] Sir Arthur Conan Doyle. Memoirs of Sherlock Holmes, volume 834 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/memho11.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext97/memho10h.zip   ((HTML)
).

Fairless:1997:GB

[806] Michael Fairless. The Gray Brethren, volume 835 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/grybr10.zip.

Nesbit:1997:PC

[807] E. (Edith) Nesbit. The Phoenix and the Carpet, volume 836 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/phcar10.zip.

Nesbit:1997:SA

[808] E. (Edith) Nesbit. The Story of the Amulet, volume 837 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tsamu10.zip.

Smiles:1997:JBP

[809] Samuel Smiles. Jasmin: Barber, Poet, Philanthropist, volume 838 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/jasmn10.zip.

Stevenson:1997:NAN

[810] Robert Louis Stevenson. New Arabian Nights, volume 839 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/narab10.zip.

Blackmore:1997:LDR

[811] R. D. (Richard Doddridge) Blackmore. Lorna Doone, A Romance of Exmoor, volume 840 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext97/lorna10.zip.

Lowell:1997:MWG

[812] Amy Lowell. Men, Women and Ghosts, volume 841 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/almwg10.zip.

Brown:1997:CB

[813] Charles Brockden Brown. Carwin the Biloquist, volume 842 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/moctb10.zip.

Simms:1997:LFM

[814] W. Gilmore (William Gilmore) Simms. The Life of Francis Marion, volume 843 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/1sfox10.zip.

Wilde:1997:IBE

[815] Oscar Wilde. The Importance of Being Earnest, volume 844 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tiobe10.zip.

Timrod:1997:PHT

[816] Henry Timrod. The Poems of Henry Timrod, volume 845 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/htimr10.zip.

Weems:1997:LGF

[817] M. L. (Mason Locke) Weems. The Life of Gen. Francis Marion, volume 846 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/wfmar10.zip.

Macaulay:1997:LAR

[818] Thomas Babington Macaulay Macaulay, Baron. Lays of Ancient Rome, volume 847 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/lrome10.zip.

Stevenson:1997:BA

[819] Robert Louis Stevenson. The Black Arrow, volume 848 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/blcka10.zip.

Jerome:1997:ITI

[820] Jerome K. Jerome, (Jerome Klapka). Idle Thoughts of an Idle Fellow, volume 849 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/jjidl10.zip.

Smith:1997:TG

[821] F. Hopkinson (Francis Hopkinson) Smith. Tom Grogan, volume 850 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/grogn10.zip.

Rowlandson:1997:NCR

[822] Mary White Rowlandson. Narrative of the Captivity and Restoration of Mrs. Mary Rowlandson, volume 851 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/crmmr10.zip.

Dewey:1997:DE

[823] John Dewey. Democracy and Education, volume 852 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dmedu10.zip.

Reu:1997:CP

[824] Johann Michael Reu. The Confutatio Pontificia, volume 853 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/cfpnt10.zip.

Wilde:1997:WNI

[825] Oscar Wilde. A Woman of No Importance, volume 854 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/awoni10.zip.

Jerome:1997:C

[826] Jerome K. Jerome, (Jerome Klapka). Clocks, volume 855 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jjclk10.zip.

Jerome:1997:D

[827] Jerome K. Jerome, (Jerome Klapka). Dreams, volume 856 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jjdrm10.zip.

Jerome:1997:E

[828] Jerome K. Jerome, (Jerome Klapka). Evergreens, volume 857 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jjevg10.zip.

Jerome:1997:S

[829] Jerome K. Jerome, (Jerome Klapka). Stage-Land, volume 858 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jjstg10.zip.

Mayo:1997:PC

[830] Margaret Mayo. Polly of the Circus, volume 859 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/pcrcs10.zip.

Mayo:1997:BM

[831] Margaret Mayo. Baby Mine, volume 860 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/bmine10.zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/bminex10.zip.

Bacon:1997:DA

[832] John Mackenzie Bacon. The Dominion of the Air, volume 861 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dmair10.zip.

Wister:1997:P

[833] Owen Wister. Philosophy 4, volume 862 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/phil410.zip.

Christie:1997:MAA

[834] Agatha Christie. The Mysterious Affair at Styles, volume 863 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/masac10.zip.

Stevenson:1997:MB

[835] Robert Louis Stevenson. Master of Ballantrae, volume 864 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/blntr10.zip.

Jerome:1997:PTF

[836] Jerome K. Jerome, (Jerome Klapka). Passing of the Third Floor Back, volume 865 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/jjp3b10.zip.

Jerome:1997:CK

[837] Jerome K. Jerome, (Jerome Klapka). The Cost of Kindness, volume 866 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/jjknd10.zip.

Jerome:1997:MKS

[838] Jerome K. Jerome, (Jerome Klapka). Mrs. Korner Sins Her Mercies, volume 867 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/jjkor10.zip.

Jerome:1997:PJ

[839] Jerome K. Jerome, (Jerome Klapka). The Philosopher’s Joke, volume 868 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/jjphj10.zip.

Jerome:1997:SNS

[840] Jerome K. Jerome, (Jerome Klapka). The Soul of Nicholas Snyders, volume 869 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/jjsny10.zip.

Jerome:1997:LUN

[841] Jerome K. Jerome, (Jerome Klapka). The Love of Ulrich Nebendahl, volume 870 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/jjulr10.zip.

Epictetus:1997:GSE

[842] Epictetus. The Golden Sayings of Epictetus, volume 871 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/epict10.zip.

Dickens:1997:RP

[843] Charles Dickens. Reprinted Pieces, volume 872 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/cdrpr10.zip.

Wilde:1997:HP

[844] Oscar Wilde. A House of Pomegranates, volume 873 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/hpomg10.zip.

Vivian:1997:HA

[845] Evelyn Charles Vivian. A History of Aeronautics, volume 874 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/haero10.zip.

Wilde:1997:DP

[846] Oscar Wilde. The Duchess of Padua, volume 875 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/dpdua10.zip.

Davis:1997:LI

[847] Rebecca Harding Davis. Life in the Iron-Mills, volume 876 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ironm10.zip.

Irving:1997:LB

[848] Washington Irving. Little Britain, volume 877 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/lbrit10.zip.

Whittier:1997:YG

[849] John Greenleaf Whittier. Yankee Gypsies, volume 878 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ynkgp10.zip.

Whittier:1997:BC

[850] John Greenleaf Whittier. The Boy Captives, volume 879 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/bcptv10.zip.

Lowell:1997:MGA

[851] James Russell Lowell. My Garden Acquaintance, volume 880 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/mgacq10.zip.

Stoddard:1997:LVH

[852] Elizabeth Drew Stoddard. Lemorne Versus Huell, volume 881 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/lvssh10.zip.

Dickens:1997:SB

[853] Charles Dickens. Sketches by Boz, volume 882 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/sbboz10.zip.

Dickens:1997:OMF

[854] Charles Dickens. Our Mutual Friend, volume 883 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/mfrnd10.zip.

MacKay:1997:MPD

[855] Charles MacKay. Memoirs of Popular Delusions V3, volume 884 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/3ppdl10.zip.

Wilde:1997:IH

[856] Oscar Wilde. An Ideal Husband, volume 885 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/ihsbn10.zip.

Gordon:1997:LC

[857] Lady Duff Gordon. Letters from the Cape, volume 886 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/lddfg10.zip.

Wilde:1997:I

[858] Oscar Wilde. Intentions, volume 887 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/ntntn10.
zip.

Dickens:1997:LTT

[859] Charles Dickens. Lazy Tour of Two Idle Apprentices, volume 888 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/lttia10.zip.

Ling:1997:TYF

[860] Princess Der Ling. Two Years in the Forbidden City, volume 889 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/tyifc10.zip.

Gibbon:1997:HDFa

[861] Edward Gibbon. The History of the Decline and Fall Of The Roman Empire, volume 890–895 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext97/dfre110.zip   (V1)
; ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/
dfre210.zip   (V2); ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dfre310.zip   (V3); ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/etext97/dfre410.zip   (V4); ftp://uiarchive.
cso.uiuc.edu/pub/etext/gutenberg/etext97/dfre510.zip   (V5)
; ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/
dfre610.zip   (V6).

Adams:1997:O

[862] John Quincy Adams. Orations, volume 896 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/objqa10.zip.

Thackeray:1997:RR

[863] William Makepeace Thackeray. The Rose and the Ring, volume 897 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/rsrng10.zip.

James:1997:LM

[864] Henry James. The Lesson of the Master, volume 898 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tlotm10.zip.

Marion:1997:WBA

[865] F. (Fulgence) Marion. Wonderful Balloon Ascents, volume 899 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/wonba10.zip.

Gibbon:1997:HDFb

[866] Edward Gibbon. The History of the Decline and Fall Of The Roman Empire, volume 900 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext97/dfr310.zip.

Marlowe:1997:JM

[867] Christopher Marlowe. The Jew of Malta, volume 901 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jmlta10.zip.

Wilde:1997:HPO

[868] Oscar Wilde. The Happy Prince and Other Tales, volume 902 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/hpaot10.zip.

ConanDoyle:1997:WC

[869] Sir Arthur Conan Doyle. The White Company, volume 903 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/whtco10.zip.

Stratton-Porter:1997:HFD

[870] Gene Stratton-Porter. Her Father’s Daughter, volume 904 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/hfdtr10.zip.

Dana:1997:WL

[871] Marvin Dana. Within the Law, volume 905 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/wnlaw10.zip.

Lowell:1997:AL

[872] James Russell Lowell. Abraham Lincoln, volume 906 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1lncn10.zip.

Jackman:1997:FM

[873] William James Jackman. Flying Machine, volume 907 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/flymc10.zip.

Shaw:1997:TPC

[874] George Bernard Shaw. Treatise on Parents and Children, volume 908 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/topac10.zip.

Filson:1997:ACD

[875] John Filson. Adventures of Col. Daniel Boone, volume 909 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1boon10.zip.

London:1997:WF

[876] Jack London. White Fang, volume 910 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/wtfng10.zip.

London:1997:TFP

[877] Jack London. Tales of the Fish Patrol, volume 911 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/totfp10.zip.

Dickens:1997:MOS

[878] Charles Dickens. Mudfog and Other Sketches, volume 912 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/mdfog10.zip.

Lermontov:1997:HOT

[879] Mikhail Yurevich Lermontov. A Hero of Our Time, volume 913 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ahero10.zip.

Dickens:1997:UT

[880] Charles Dickens. The Uncommercial Traveller, volume 914 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/unctr10.zip.

Hazeltine:1997:LWC

[881] Alice Isabel Hazeltine. Library Work with Children, volume 915 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/lwwch10.zip.

Dickens:1997:SYC

[882] Charles Dickens. Sketches of Young Couples, volume 916 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/yngcp10.zip.

Dickens:1997:BRR

[883] Charles Dickens. Barnaby Rudge, 80’s Riots, volume 917 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/rudge10.zip.

Dickens:1997:SYG

[884] Charles Dickens. Sketches of Young Gentlemen, volume 918 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/skygm10.zip.

Spinoza:1997:Ea

[885] Benedict de Spinoza. Ethics, volume 919 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1spne10.zip.

Spinoza:1997:Eb

[886] Benedict de Spinoza. Ethics, volume 920 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/2spne10.zip.

Wilde:1997:P

[887] Oscar Wilde. De Profundis, volume 921 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/dprof10.zip.

Dickens:1997:SUT

[888] Charles Dickens. Sunday Under Three Heads, volume 922 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/suths10.zip.

James:1997:LFM

[889] William Dobein James. Life of Francis Marion, volume 923 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jjmar11.zip.

Dickens:1997:BRA

[890] Charles Dickens. To Be Read At Dusk, volume 924 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/rddsk10.zip.

Anonymous:1997:USP

[891] Anonymous. United States Presidents’ Inaugural Speeches, volume 925 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/uspis10.zip.

Miller:1997:DI

[892] Gustavus Hindman Miller. 10,000 Dreams Interpreted, volume 926 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/drmnt10.zip.

Dickens:1997:L

[893] Charles Dickens. The Lamplighter, volume 927 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/lmplt10.zip.

Carroll:1997:AWH

[894] Lewis Carroll. Alice In Wonderland, HTML Version of 30th Edition, volume 928 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/alice30h.htm;   ftp://
uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/alice30h.
zip.

StJude:1997:CF

[895] St. Jude. The Cyberpunk Fakebook, volume 929 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/fakeb10.zip.

Water:1997:CD

[896] Mrs. W. G. Water. The Cook’s Decameron, volume 930 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ckdec10.zip.

Gilbert:1997:BB

[897] W. S. (William Schwenck) Gilbert. The Bab Ballads, volume 931 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/2babb10.zip.

Poe:1997:FHU

[898] Edgar Allan Poe. Fall of the House of Usher, volume 932 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/usher10.zip.

Gilbert:1997:MBB

[899] W. S. (William Schwenck) Gilbert. More Bab Ballads, volume 933 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/3babb10.zip.

Gilbert:1997:SS

[900] W. S. (William Schwenck) Gilbert. Songs of a Savoyard, volume 934 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/svyrd10.zip.

Smiles:1997:SHC

[901] Samuel Smiles. Self Help; Conduct and Perseverance, volume 935 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/selfh10.zip.

Wiggin:1997:VW

[902] Kate Douglas Smith Wiggin. The Village Watch-Tower, volume 936 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/vilwt10.zip.

Ryan:1997:PPR

[903] Kenneth Ryan, Father. Poems: Patriotic, Religious, etc., volume 937 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/fryan10.zip.

Bower:1997:GI

[904] B. M. Bower. Good Indian, volume 938 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/gndin10.zip.

Smiles:1997:LTT

[905] Samuel Smiles. Life of Thomas Telford, volume 939 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/tlfrd10.zip.

Cooper:1997:LM

[906] James Fenimore Cooper. Last of the Mohicans, volume 940 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/mohic10.zip.

Guest:1997:JF

[907] Edgar A. (Edgar Albert) Guest. Just Folks, volume 941 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jfolk10.zip.

Hudson:1997:GM

[908] W. H. (William Henry) Hudson. Green Mansions, volume 942 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/gmans10.zip.

Shaw:1997:M

[909] George Bernard Shaw. Misalliance, volume 943 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/msali10.zip.

Darwin:1997:VB

[910] Charles Darwin. The Voyage of the Beagle, volume 944 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/vbgle10.zip.

Haldeman-Julius:1997:D

[911] E. (Emanuel) Haldeman-Julius. Dust, volume 945 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/dsthj10.zip.

Austen:1997:LS

[912] Jane Austen. Lady Susan, volume 946 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/lsusn10.zip.

Southey:1997:LHL

[913] Robert Southey. The Life of Horatio Lord Nelson, volume 947 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/hnlsn10.zip.

Spinoza:1997:Ec

[914] Benedict de Spinoza. Ethics, volume 948 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/3spne10.zip.

Appleton:1997:TSHa

[915] Victor Appleton. Tom Swift and his Submarine Boat, volume 949 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/04tom10.zip.

Appleton:1997:TSHb

[916] Victor Appleton. Tom Swift and his Electric Runabout, volume 950 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/05tom10.zip.

Appleton:1997:TSHc

[917] Victor Appleton. Tom Swift and his Sky Racer, volume 951 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/09tom10.zip.

Appleton:1997:TSHd

[918] Victor Appleton. Tom Swift and his Air Glider, volume 952 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/12tom10.zip.

Appleton:1997:TSHe

[919] Victor Appleton. Tom Swift and his Big Tunnel, volume 953 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/19tom10.zip.

Appleton:1997:TSHf

[920] Victor Appleton. Tom Swift and his War Tank, volume 954 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/21tom10.zip.

Baum:1997:PGO

[921] L. Frank (Lyman Frank) Baum. The Patchwork Girl of Oz, volume 955 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/07woz10.zip.

Baum:1997:TO

[922] L. Frank (Lyman Frank) Baum. Tik-Tok of Oz, volume 956 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/08woz10.zip.

Baum:1997:SO

[923] L. Frank (Lyman Frank) Baum. The Scarecrow of Oz, volume 957 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/09woz10.zip.

Baum:1997:RO

[924] L. Frank (Lyman Frank) Baum. Rinkitink In Oz, volume 958 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/10woz10.zip.

Baum:1997:LPO

[925] L. Frank (Lyman Frank) Baum. The Lost Princess of Oz, volume 959 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/11woz11.zip.

Baum:1997:TWO

[926] L. Frank (Lyman Frank) Baum. The Tin Woodman of Oz, volume 960 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/12woz10.zip.

Baum:1997:GO

[927] L. Frank (Lyman Frank) Baum. Glinda of Oz, volume 961 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/14woz10.zip.

Kendall:1997:PHK

[928] Henry Kendall. The Poems of Henry Kendall, volume 962 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/phknd10.zip.

Dickens:1997:LD

[929] Charles Dickens. Little Dorrit, volume 963 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/ldort10.zip.

Pyle:1997:ARH

[930] Howard Pyle. The Adventures of Robin Hood, volume 964 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/2rbnh10.zip.

Dumas:1997:BT

[931] Alexandre Dumas. The Black Tulip, volume 965 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/tbtlp10.zip.

Peacock:1997:MM

[932] Thomas Love Peacock. Maid Marian, volume 966 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/maidm10.zip.

Dickens:1997:NN

[933] Charles Dickens. Nicholas Nickleby, volume 967 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/ncklb10.zip.

Dickens:1997:MC

[934] Charles Dickens. Martin Chuzzlewit, volume 968 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/chuzz10.zip.

Bronte:1997:TWH

[935] Anne Brontë. The Tenant of Wildfell Hall, volume 969 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/wldfl10.zip.

Stewart:1997:UJP

[936] Cal Stewart. Uncle Josh’s Punkin Centre Stories, volume 970 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ncjsh10.zip.

Spinoza:1997:Ed

[937] Benedict de Spinoza. Ethics, volume 971 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/4spne10.zip.

Bierce:1997:DD

[938] Ambrose Bierce. The Devil’s Dictionary, volume 972 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dvldc10.zip.

Pyle:1997:HPB

[939] Howard Pyle. Howard Pyle’s Book of Pirates, volume 973 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/hpprt10.zip.

Conrad:1997:SA

[940] Joseph Conrad. The Secret Agent, volume 974 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/agent10.zip.

Spinoza:1997:Ee

[941] Benedict de Spinoza. Ethics, volume 975 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/5spne10.zip.

Hawthorne:1997:TT

[942] Nathaniel Hawthorne. Tanglewood Tales, volume 976 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tnglw10.zip.

Kipling:1997:AN

[943] Rudyard Kipling. American Notes, volume 977 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/amrnt10.zip.

Freeman:1997:YP

[944] Mary Eleanor Wilkins Freeman. The Yates Pride, volume 978 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ytspr10.zip.

Munro:1997:HT

[945] J. Munro. Heroes of the Telegraph, volume 979 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/htgrf10.zip.

Tarkington:1997:AA

[946] Booth Tarkington. Alice Adams, volume 980 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/aladm10.zip.

Gummere:1997:B

[947] Frances Barton Gummere. Beowulf, volume 981 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/bwulf10.zip.

Lear:1997:BN

[948] Edward Lear. The Book of Nonsense, volume 982 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/nnsns10.zip.

Defoe:1997:ECE

[949] Daniel Defoe. Eastern Counties of England, volume 983 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ttece10.zip.

Gordon:1997:WWW

[950] Irwin Leslie Gordon. Who Was Who: 5000 BC – 1914, volume 984 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/wwasw10.zip.

Tolstoy:1997:FS

[951] Count Leo Nikoleyevich Tolstoy. Father Sergius, volume 985 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/fsrgs10.zip.

Tolstoy:1997:MM

[952] Count Leo Nikoleyevich Tolstoy. Master and Man, volume 986 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/mramn10.zip.

Anonymous:1997:PSM

[953] Anonymous. Popular Science Monthly, Oct, Nov, Dec, 1915 V.86, volume 987 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/86psm10.zip.

Key:1997:EC

[954] Ellen Karolina Sofia Key. The Education of the Child, volume 988 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/edkid10.zip.

Spinoza:1997:TTP

[955] Benedict de Spinoza. Theologico-Political Treatise P1, volume 989–992 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/etext97/1spnt10.zip   (Part 1);   ftp:/
/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/2spnt10.
zip (Part   2); ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/3spnt10.zip   (Part 3);   ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/etext97/4spnt10.zip (Part   4).

Higginson:1997:MOR

[956] Thomas Wentworth Higginson. Malbone: An Oldport Romance, volume 993 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/malbn10.zip.

Synge:1997:RS

[957] John Millington Synge. Riders to the Sea, volume 994 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/rdrse10.zip.

Service:1997:BB

[958] Robert William Service. Ballads of a Bohemian, volume 995 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/blbhm10.zip.

CervantesSaavedra:1997:DQ

[959] Miguel de Cervantes Saavedra. Don Quixote, volume 996 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1donq10.zip.

DanteAlighieri:1997:DCDa

[960] Dante Alighieri. Divina Commedia di Dante: Inferno, 7-bit Italian, volume 997 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/1ddcd09.zip.

DanteAlighieri:1997:DCDb

[961] Dante Alighieri. Divina Commedia di Dante: Purgatorio 7-bit Italian, volume 998 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/2ddcd09.zip.

DanteAlighieri:1997:DCDc

[962] Dante Alighieri. Divina Commedia di Dante: Paradiso, 7-bit Italian, volume 999 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/3ddcd09.zip.

DanteAlighieri:1997:DCDd

[963] Dante Alighieri. La Divina Commedia di Dante in Italian, 7-bit text, volume 1000 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/0ddcd09.zip.

DanteAlighieri:1997:LTDa

[964] Dante Alighieri. Longfellow’s Translation of Dante, Inferno, volume 1001 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ddcl10.zip.

DanteAlighieri:1997:LTDb

[965] Dante Alighieri. Longfellow’s Translation of Dante, Purgatory, volume 1002 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/2ddcl10.zip.

DanteAlighieri:1997:LTDc

[966] Dante Alighieri. Longfellow’s Translation of Dante, Paradise, volume 1003 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/3ddcl10.zip.

DanteAlighieri:1997:LTDd

[967] Dante Alighieri. Longfellow’s Translation of Dante, Entire Comedy, volume 1004 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/0ddcl10.zip.

DanteAlighieri:1997:HFCa

[968] Dante Alighieri. H. F. Cary’s Translation of Dante, Hell, volume 1005 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ddcc10.zip.

DanteAlighieri:1997:HFCb

[969] Dante Alighieri. H. F. Cary’s Translation of Dante, Purgatory, volume 1006 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/2ddcc10.zip.

DanteAlighieri:1997:HFCc

[970] Dante Alighieri. H. F. Cary’s Translation of Dante, Paradise, volume 1007 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/3ddcc10.zip.

DanteAlighieri:1997:HFCd

[971] Dante Alighieri. H. F. Cary’s Translation of Dante, Entire Comedy, volume 1008 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/0ddcc10.zip.

DanteAlighieri:1997:DCDe

[972] Dante Alighieri. Divina Commedia di Dante: Inferno, volume 1009 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ddc809a.zip.

DanteAlighieri:1997:DCDf

[973] Dante Alighieri. Divina Commedia di Dante: Purgatorio, volume 1010 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/2ddc809a.zip.

DanteAlighieri:1997:DCDg

[974] Dante Alighieri. Divina Commedia di Dante: Paradiso, volume 1011 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/3ddc809a.zip.

DanteAlighieri:1997:DCDh

[975] Dante Alighieri. La Divina Commedia di Dante in Italian, 8-bit text, volume 1012 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/0ddc809a.zip.

Wells:1997:FMM

[976] H. G. (Herbert George) Wells. The First Men In The Moon, volume 1013 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/fmitm10.zip.

Bower:1997:LDT

[977] B. M. Bower. The Lure of the Dim Trails, volume 1014 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ldmtr10.zip.

Parkman:1997:OT

[978] Francis Parkman. The Oregon Trail, volume 1015 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/ortrl10.zip.

Spinoza:1997:IU

[979] Benedict de Spinoza. Improvement of Understanding, volume 1016 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/spint10.zip.

Wilde:1997:SM

[980] Oscar Wilde. The Soul of Man, volume 1017 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/slman10.zip.

Kingsley:1997:W

[981] Charles Kingsley. The Water-Babies, volume 1018 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/wtrbs10.zip.

Bronte:1997:PBS

[982] Anne Brontë. Poems by the Brontë Sisters, volume 1019 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/brntp10.zip.

Lowell:1997:SBP

[983] Amy Lowell. Sword Blades and Poppy Seed, volume 1020 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/sbaps10.zip.

Lindsay:1997:COP

[984] Vachel Lindsay. The Congo and Other Poems, volume 1021 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/cngop10.zip.

Thoreau:1997:W

[985] Henry David Thoreau. Walking, volume 1022 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/wlkng10.zip.

Dickens:1997:BH

[986] Charles Dickens. Bleak House, volume 1023 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/blkhs10.zip.

Stevenson:1997:W

[987] Robert Louis Stevenson. The Wrecker, volume 1024 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/wrckr10.zip.

Rumford:1997:E

[988] Benjamin Rumford, Graf von. Essays, volume 1025 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/essbr10.zip.

Grossmith:1997:DN

[989] George Grossmith. Diary of a Nobody, volume 1026 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/dnbdy10.zip.

Grey:1997:LSR

[990] Zane Grey. The Lone Star Ranger, volume 1027 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/lrngr10.zip.

Bronte:1997:P

[991] Charlotte Brontë. The Professor, volume 1028 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/tprof10.zip.

London:1997:N

[992] Jack London. The Night-Born, volume 1029 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/ntbrn10.zip.

MacKay:1997:CSB

[993] Charles MacKay. The Cavalier Songs and Ballads of England from 1642 to 1684, volume 1030 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext97/csboe10.zip.

Wilde:1997:COP

[994] Oscar Wilde. Charmides and Other Poems, volume 1031 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/crmds10.zip.

James:1997:P

[995] Henry James. The Pupil, volume 1032 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/pupil10.zip.

Wiggin:1997:ROR

[996] Kate Douglas Smith Wiggin. Rose O’ the River, volume 1033 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/rorvr10.zip.

Owen:1997:P

[997] Wilfred Owen. Poems, volume 1034 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/wowen11.
zip.

Robinson:1997:MAS

[998] Edwin Arlington Robinson. The Man against the Sky, volume 1035 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/tmats11.zip.

Lawson:1997:JWH

[999] Henry Lawson. Joe Wilson and His Mates, volume 1036 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jwahm11.zip.

Venables:1997:LJB

[1000] Edmund Venables. The Life of John Bunyan, volume 1037 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/lfbyn10.zip.

Raleigh:1997:S

[1001] Sir Walter Alexander Raleigh. Style, volume 1038 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/style10.zip.

Livingstone:1997:MTS

[1002] David Livingstone. Missionary Travels in South Africa, volume 1039 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/mtrav10.zip.

Robinson:1997:TT

[1003] Edwin Arlington Robinson. The Three Taverns, volume 1040 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/3tavs10.zip.

Shakespeare:1997:SSa

[1004] William Shakespeare. Shakespeare’s Sonnets, volume 1041 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/wssnt10.zip.

Meredith:1997:RLO

[1005] George Meredith. A Reading of Life, Other Poems, volume 1042 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/rdlif10.zip.

McCabe:1997:SE

[1006] Joseph McCabe. The Story of Evolution, volume 1043 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tsoev10.zip.

Twain:1997:CSV

[1007] Mark Twain. Captain Stormfield’s Visit to Heaven, volume 1044 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/cptsf10.zip.

Shakespeare:1997:VA

[1008] William Shakespeare. Venus and Adonis, volume 1045 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/wsvns10.zip.

Wells:1997:GIK

[1009] H. G. (Herbert George) Wells. God The Invisible King, volume 1046 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/godik10.zip.

Wells:1997:NM

[1010] H. G. (Herbert George) Wells. The New Machiavelli, volume 1047 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/nmchv10.zip.

Dyke:1997:RP

[1011] Henry van Dyke. The Ruling Passion, volume 1048 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/rlpsn10.zip.

Summerhayes:1997:VA

[1012] Martha Summerhayes. Vanished Arizona, volume 1049 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/variz10.zip.

Shaw:1997:DLS

[1013] George Bernard Shaw. Dark Lady of the Sonnets, volume 1050 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dlots10.zip.

Carlyle:1997:SR

[1014] Thomas Carlyle. Sartor Resartus, volume 1051 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/srtrs10.zip.

Anonymous:1997:SST

[1015] Anonymous. Step by Step; or Tidy’s Way to Freedom, volume 1052 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/tidys11.zip.

Conrad:1997:WT

[1016] Joseph Conrad. Within The Tides, volume 1053 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/wthnt10.zip.

Lang:1997:CB

[1017] Andrew Lang. A Collection of Ballads, volume 1054 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/cblad10.zip.

Conrad:1997:LS

[1018] Joseph Conrad. ’Twixt Land and Sea, volume 1055 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/twxls10.zip.

London:1997:ME

[1019] Jack London. Martin Eden, volume 1056 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/meden10.zip.

Wilde:1997:BRG

[1020] Oscar Wilde. Ballad of Reading Gaol, and other poems, volume 1057 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/pmwld10.zip.

Conrad:1997:MS

[1021] Joseph Conrad. The Mirror of the Sea, volume 1058 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/tmots10.zip.

Wells:1997:WSF

[1022] H. G. (Herbert George) Wells. The World Set Free, volume 1059 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/twsfr10.zip.

Lang:1997:GP

[1023] Andrew Lang. Grass of Parnassus, volume 1060 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/grprn10.zip.

Fiske:1997:MM

[1024] John Fiske. Myths and Myth-Makers, volume 1061 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/mythm10.zip.

Poe:1997:PCE

[1025] Edgar Allan Poe. 1st PG Collection of Edgar Allan Poe, volume 1062– 1065 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/1epoe10.zip.

Anonymous:1997:PMU

[1026] Anonymous. Personal Memoirs of U. S. Grant, volume 1067–1068 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/1musg10.zip   (Volume 1);
ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/2musg10.
zip (Volume   2).

Zola:1997:PCE

[1027] Emile Zola. 1st PG Collection of Emile Zola, volume 1069–1073 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1zola10.zip.

London:1997:SW

[1028] Jack London. The Sea Wolf, volume 1074 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/cwolf10.zip.

London:1997:SSO

[1029] Jack London. The Strength of the Strong, and other tales, volume 1075 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/sstrg10.zip.

Bramah:1997:WKL

[1030] Ernest Bramah. The Wallet of Kai Lung, volume 1076 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/wklng10.zip.

Bramah:1997:MK

[1031] Ernest Bramah. The Mirror of Kong Ho, volume 1077 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/kongh10.zip.

Altsheler:1997:SV

[1032] Joseph A. (Joseph Alexander) Altsheler. The Scouts of the Valley, volume 1078 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext97/sctvl10.zip.

Sterne:1997:LTS

[1033] Laurence Sterne. Life of Tristram Shandy, volume 1079 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/shndy10.zip.

Swift:1997:MP

[1034] Jonathan Swift. A Modest Proposal, volume 1080 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/mdprp10.zip.

Gogol:1997:DSa

[1035] Nikolai Vasilievich Gogol. Dead Souls, volume 1081 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/dsols10.zip.

Bishop:1997:VPC

[1036] Nathaniel H. (Nathaniel Holmes) Bishop. Voyage of The Paper Canoe, volume 1082 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext97/pprcn11.zip.

Gogol:1997:DSb

[1037] Nikolai Vasilievich Gogol. Dead Souls, volume 1083 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/dsols10.zip.

Presbyterian:1997:RTT

[1038] Presbyterian Ladies’ Aid. Recipes Tried and True, volume 1084 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/tandt10.zip.

Carlyle:1997:LJS

[1039] Thomas Carlyle. Life of John Sterling, volume 1085 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/strlg10.zip.

Twain:1997:HT

[1040] Mark Twain. A Horse’s Tale, volume 1086 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/hrstl10.zip.

Roth:1997:B

[1041] Lewis Roth. Baartock, volume 1087 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1997. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext97/brtck10.
zip.

Seton:1997:RW

[1042] Ernest Thompson Seton. Rolf in the Woods, volume 1088 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/rolfw11.zip.

London:1997:MOS

[1043] Jack London. Moon-Face and Other Stories, volume 1089 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/mface10.zip.

Swift:1997:BP

[1044] Jonathan Swift. Bickerstaff-Partridge Papers, volume 1090 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/bstaf10.zip.

Carlyle:1997:HHW

[1045] Thomas Carlyle. Heroes and Hero Worship, volume 1091 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/heros10.zip.

Cambrensis:1997:DW

[1046] Giraldus Cambrensis. The Description of Wales, volume 1092 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dscwl10.zip.

James:1997:BJ

[1047] Henry James. The Beast in the Jungle, volume 1093 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/bstjg10.zip.

Marlowe:1997:TGP

[1048] Christopher Marlowe. Tamburlaine the Great PT 1, volume 1094 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/tmbn110.zip.

Grey:1997:LWS

[1049] Zane Grey. Light of the Western Stars, volume 1095 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/lwstr10.zip.

London:1997:FM

[1050] Jack London. The Faith of Men, volume 1096 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/fthmn10.zip.

Shaw:1997:MWP

[1051] George Bernard Shaw. Mrs. Warren’s Profession, volume 1097 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/wrpro10.zip.

Tarkington:1997:TN

[1052] Booth Tarkington. The Turmoil, A novel, volume 1098 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/turmo10.zip.

White:1997:R

[1053] Stewart Edward White. The Riverman, volume 1099 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/rvrmn11.zip.

Shakespeare:1997:KHVa

[1054] William Shakespeare. King Henry VI, Part 1, volume 1100 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws0110.zip.

Shakespeare:1997:KHVb

[1055] William Shakespeare. King Henry VI, Part 2, volume 1101 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws0210.zip.

Shakespeare:1997:KHVc

[1056] William Shakespeare. King Henry VI, Part 3, volume 1102 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws0310.zip.

Shakespeare:1997:KRIa

[1057] William Shakespeare. King Richard III, volume 1103 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws0411.zip.

Shakespeare:1997:CE

[1058] William Shakespeare. The Comedy of Errors, volume 1104 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws0610.zip.

Shakespeare:1997:SSb

[1059] William Shakespeare. The Shakespearian Sonnets, volume 1105 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws0710.zip.

Shakespeare:1997:TA

[1060] William Shakespeare. Titus Andronicus, volume 1106 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws0910.zip.

Shakespeare:1997:TS

[1061] William Shakespeare. The Taming of the Shrew, volume 1107 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws1010.zip.

Shakespeare:1997:TGV

[1062] William Shakespeare. Two Gentlemen of Verona, volume 1108 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws1110.zip.

Shakespeare:1997:LLL

[1063] William Shakespeare. Love’s Labour’s Lost, volume 1109 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws1210.zip.

Shakespeare:1997:KJ

[1064] William Shakespeare. King John, volume 1110 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws1410.zip.

Shakespeare:1997:KRIb

[1065] William Shakespeare. King Richard II, volume 1111 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws1510.zip.

Shakespeare:1997:RJ

[1066] William Shakespeare. Romeo and Juliet, volume 1112 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws1610.zip.

Shakespeare:1997:MND

[1067] William Shakespeare. A Midsummer Night’s Dream, volume 1113 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ws1710.zip.

Shakespeare:1997:MV

[1068] William Shakespeare. The Merchant of Venice, volume 1114 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws1810.zip.

Shakespeare:1997:KHIa

[1069] William Shakespeare. King Henry IV Part 1, volume 1115 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws1910.zip.

Shakespeare:1997:MWW

[1070] William Shakespeare. The Merry Wives of Windsor, volume 1116 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ws2010.zip.

Shakespeare:1997:KHIb

[1071] William Shakespeare. King Henry IV, Part 2, volume 1117 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws2110.zip.

Shakespeare:1997:MAA

[1072] William Shakespeare. Much Ado about Nothing, volume 1118 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws2210.zip.

Shakespeare:1997:KHVd

[1073] William Shakespeare. King Henry V, volume 1119 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws2310.zip.

Shakespeare:1997:JC

[1074] William Shakespeare. Julius Caesar, volume 1120 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws2410.zip.

Shakespeare:1997:YLI

[1075] William Shakespeare. As You Like It, volume 1121 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws2510.zip.

Shakespeare:1997:HPD

[1076] William Shakespeare. Hamlet, Prince of Denmark, volume 1122 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ws2610.zip.

Shakespeare:1997:TNW

[1077] William Shakespeare. Twelfth Night; or What You Will, volume 1123 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ws2810.zip.

Shakespeare:1997:HTC

[1078] William Shakespeare. History of Troilus and Cressida, volume 1124 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ws2910.zip.

Shakespeare:1997:AWE

[1079] William Shakespeare. All’s Well that End’s Well, volume 1125 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws3010.zip.

Shakespeare:1997:MM

[1080] William Shakespeare. Measure for Measure, volume 1126 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws3110.zip.

Shakespeare:1997:OMV

[1081] William Shakespeare. Othello, The Moor of Venice, volume 1127 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ws3210.zip.

Shakespeare:1997:KL

[1082] William Shakespeare. King Lear, volume 1128 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws3310.zip.

Shakespeare:1997:M

[1083] William Shakespeare. Macbeth, volume 1129 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws3410.zip.

Shakespeare:1997:AC

[1084] William Shakespeare. Antony and Cleopatra, volume 1130 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws3510.zip.

Shakespeare:1997:TC

[1085] William Shakespeare. The Tragedy of Coriolanus, volume 1131 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws3610.zip.

Shakespeare:1997:LTA

[1086] William Shakespeare. The Life of Timon of Athens, volume 1132 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/1ws3710.zip.

Shakespeare:1997:C

[1087] William Shakespeare. Cymbeline, volume 1133 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws3910.zip.

Shakespeare:1997:WT

[1088] William Shakespeare. The Winter’s Tale, volume 1134 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws4010.zip.

Shakespeare:1997:T

[1089] William Shakespeare. Tempest, volume 1135 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/1ws4110.zip.

Shakespeare:1997:KHVe

[1090] William Shakespeare. King Henry VIII, volume 1136 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws4210.zip.

Shakespeare:1997:LC

[1091] William Shakespeare. A Lover’s Complaint, volume 1137 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/1ws4410.zip.

Wells:1997:RM

[1092] H. G. (Herbert George) Wells. The Research Magnificent, volume 1138 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/rschm10.zip.

Dyke:1997:FL

[1093] Henry van Dyke. Fisherman’s Luck, volume 1139 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/fshlk10.zip.

Carlyle:1997:LP

[1094] Thomas Carlyle. Latter-Day Pamphlets, volume 1140 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/latda10.zip.

Wilde:1997:SPO

[1095] Oscar Wilde. Selected Poems of Oscar Wilde, volume 1141 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/spoow10.zip.

Conrad:1997:T

[1096] Joseph Conrad. Typhoon, volume 1142 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/typhn10.zip.

Conrad:1997:NLL

[1097] Joseph Conrad. Notes on Life and Letters, volume 1143 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/ntlfl10.zip.

James:1997:C

[1098] Henry James. In the Cage, volume 1144 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/incag10.zip.

Hope:1997:RH

[1099] Anthony Hope. Rupert of Hentzau, volume 1145 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/rprhn10.zip.

Fielding:1997:JVL

[1100] Henry Fielding. Journal of A Voyage to Lisbon, volume 1146 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jlsbn10.zip.

Fielding:1997:WN

[1101] Henry Fielding. From This World to the Next, volume 1147 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/jtwtn10.zip.

Cambrensis:1997:IBW

[1102] Giraldus Cambrensis. Itinerary of Baldwin in Wales, volume 1148 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext97/itwls10.zip.

Defoe:1997:LLE

[1103] Daniel Defoe. From London to Land’s End, volume 1149 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/lndle10.zip.

Grammaticus:1997:DH

[1104] Saxo Grammaticus. The Danish History, volume 1150 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/dnhst10.zip.

Anonymous:1997:N

[1105] Anonymous. The Nibelungenlied, volume 1151 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext97/nblng10.zip.

Anonymous:1997:SV

[1106] Anonymous. The Story of the Volsungs, volume 1152 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1997. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext97/vlsng10.zip.

Burroughs:1998:CM

[1107] Edgar Rice Burroughs. The Chessman of Mars, volume 1153 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/cmars10.zip.

Lofting:1998:VDD

[1108] Hugh Lofting. Voyages of Dr. Doolittle, volume 1154 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/vdrdl10.zip.

Christie:1998:SA

[1109] Agatha Christie. Secret Adversary, volume 1155 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/secad10.zip.

Lewis:1998:B

[1110] Sinclair Lewis. Babbit, volume 1156 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/babit10.
zip.

Sinclair:1998:DG

[1111] Upton Sinclair. Damaged Goods, volume 1157 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/

etext98/dmgds10.zip. A novelization of the play “Les Avaries”, by Eugène Brieux.

Tarkington:1998:PS

[1112] Booth Tarkington. Penrod and Sam, volume 1158 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/pnrds10.zip.

Rohmer:1998:F

[1113] Sax Rohmer. Fire-Tongue, volume 1159 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/firtg10.zip.

London:1998:G

[1114] Jack London. The Game, volume 1160 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/tgame10.zip.

London:1998:JI

[1115] Jack London. Jerry of the Islands, volume 1161 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/jrisl10.zip.

London:1998:JS

[1116] Jack London. The Jacket (Star-Rover), volume 1162 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/jaket10.zip.

London:1998:A

[1117] Jack London. Adventure, volume 1163 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/advnt10.
zip.

London:1998:IH

[1118] Jack London. The Iron Heel, volume 1164 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/irnhl10.zip.

Rittenhouse:1998:LBM

[1119] Jessie Belle Rittenhouse. The Little Book of Modern Verse, volume 1165 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/lbkmv10.zip.

Rittenhouse:1998:SBM

[1120] Jessie Belle Rittenhouse. The Second Book of Modern Verse, volume 1166 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/sbkmv10.zip.

Green:1998:SD

[1121] Anna Katharine Green. A Strange Disappearance, volume 1167 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/sdspr10.zip.

Duncan:1998:PD

[1122] Sara Jeannette Duncan. The Pool in the Desert, volume 1168 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/pldst10.zip.

Xenophon:1998:Aa

[1123] 431-355 BC Xenophon. Agesilaus, volume 1169 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/agsls10.zip.

Xenophon:1998:Ab

[1124] 431-355 BC Xenophon. Anabasis, volume 1170 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/anbss10.zip.

Xenophon:1998:Ac

[1125] 431-355 BC Xenophon. The Apology, volume 1171 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/aplgy10.zip.

Xenophon:1998:CG

[1126] 431-355 BC Xenophon. The Cavalry General, volume 1172 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/cvlry10.zip.

Xenophon:1998:E

[1127] 431-355 BC Xenophon. The Economist, volume 1173 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/econm10.zip.

Xenophon:1998:Ha

[1128] 431-355 BC Xenophon. Hellenica, volume 1174 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/hllnc10.zip.

Xenophon:1998:Hb

[1129] 431-355 BC Xenophon. Hiero, volume 1175 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/hiero10.zip.

Xenophon:1998:Hc

[1130] 431-355 BC Xenophon. On Horsemanship, volume 1176 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/hrsmn10.zip.

Xenophon:1998:M

[1131] 431-355 BC Xenophon. The Memorabilia, volume 1177 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mmrbi10.zip.

Xenophon:1998:PAL

[1132] 431-355 BC Xenophon. Polity Athenians and Lacedaemonians, volume 1178 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/pltis10.zip.

Xenophon:1998:R

[1133] 431-355 BC Xenophon. On Revenues, volume 1179 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/rvnue10.zip.

Xenophon:1998:Sa

[1134] 431-355 BC Xenophon. The Sportsman, volume 1180 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/sport10.zip.

Xenophon:1998:Sb

[1135] 431-355 BC Xenophon. The Symposium, volume 1181 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/sympm10.zip.

Rohmer:1998:D

[1136] Sax Rohmer. Dope, volume 1182 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/dope10.
zip.

Rohmer:1998:RDF

[1137] Sax Rohmer. The Return of Dr. Fu-Manchu, volume 1183 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/rfumn10.zip.

Dumas:1998:CMC

[1138] Alexandre Dumas. The Count of Monte Cristo, volume 1184 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/crsto10.zip.

Lamb:1998:TSa

[1139] Charles Lamb. Tales From Shakespeare, volume 1185 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/shlmb10.zip.

Meynell:1998:P

[1140] Alice Meynell. Poems, volume 1186 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/pomam10.
zip.

London:1998:WC

[1141] Jack London. War of the Classes, volume 1187 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/wrcls10.zip.

Stoker:1998:LWW

[1142] Bram (Abraham) Stoker. Lair of the White Worm, volume 1188 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/lrwhw10.zip.

Balzac:1998:M

[1143] Honoré de Balzac. The Message, volume 1189 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/msage10.zip.

James:1998:JC

[1144] Henry James. The Jolly Corner, volume 1190 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/jllyc10.zip.

Congreve:1998:D

[1145] William Congreve. The Double-Dealer, volume 1191 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/dbdlr10.zip.

Congreve:1998:OB

[1146] William Congreve. The Old Bachelor, volume 1192 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/oldba10.zip.

James:1998:CF

[1147] Henry James. The Coxon Fund, volume 1193 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/coxon10.zip.

Rougemont:1998:ALR

[1148] Louis de Rougemont. Adventures of Louis de Rougemont, volume 1194 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/advlr10.zip.

James:1998:G

[1149] Henry James. Glasses, volume 1195 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/glses10.
zip.

Balzac:1998:Pa

[1150] Honoré de Balzac. The Purse, volume 1196 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/purse10.zip.

Gogol:1998:TBO

[1151] Nikolai Vasilievich Gogol. Taras Bulba and Other Tales, volume 1197 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/taras10.zip.

Boldrewood:1998:RUA

[1152] Rolf Boldrewood. Robbery Under Arms, volume 1198 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/robry10.zip.

Stevens:1998:AAV

[1153] Bertram Stevens. An Anthology of Australian Verse, volume 1199 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/ozvrs10.zip.

Rabelais:1998:GP

[1154] Francis Rabelais. Gargantua and Pantagruel, volume 1200 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ggpnt10.zip.

Spooner:1998:ETJ

[1155] Lysander Spooner. Essay on the Trial By Jury, volume 1201 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/tbjry10.zip.

Conrad:1998:TU

[1156] Joseph Conrad. Tales of Unrest, volume 1202 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/tnrst10.zip.

Hope:1998:DD

[1157] Anthony Hope. Dolly Dialogues, volume 1203 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/dlydl10.zip.

Bower:1998:CF

[1158] B. M. Bower. Cabin Fever, volume 1204 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/cabfv10.zip.

Meynell:1998:CL

[1159] Alice Meynell. The Colour of Life, volume 1205 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/clrlf10.zip.

Bower:1998:FUR

[1160] B. M. Bower. The Flying U Ranch, volume 1206 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/flurn10.zip.

Haggard:1998:NL

[1161] H. Rider Haggard. Nada the Lily, volume 1207 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/ndlly10.zip.

London:1998:SST

[1162] Jack London. South Sea Tales, volume 1208 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/soset10.zip.

Prescott:1998:HCPa

[1163] William Hickling Prescott. History Of The Conquest Of Peru, volume 1209 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/hcpru10.zip.

Hearn:1998:K

[1164] Lafcadio Hearn. Kwaidan, volume 1210 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/kwidn10.zip.

Herrick:1998:LPR

[1165] Robert Herrick. From The Lyrical Poems Of Robert Herrick, volume 1211 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/lporh10.zip.

Austen:1998:LFA

[1166] Jane Austen. Love and Friendship, et al., volume 1212 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/lvfnd10.zip.

Twain:1998:MCH

[1167] Mark Twain. Man That Corrupted Hadleyburg, volume 1213 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/hdlyb10.zip.

Rinehart:1998:SSS

[1168] Mary Roberts Rinehart. The Street of Seven Stars, volume 1214 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/sstrs10.zip.

Balzac:1998:EL

[1169] Honoré de Balzac. The Elixir of Life, volume 1215 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/lxrlf10.zip.

Babbage:1998:DSE

[1170] Charles Babbage. Decline of Science in England, volume 1216 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/dosie10.zip.

Wiggin:1998:PES

[1171] Kate Douglas Smith Wiggin. Penelope’s Experiences in Scotland, volume 1217 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/pesct10.zip.

Packard:1998:AJD

[1172] Frank L. (Frank Lucius) Packard. The Adventures of Jimmie Dale, volume 1218 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/advjd10.zip.

Meredith:1998:ECC

[1173] George Meredith. Essay on Comedy, Comic Spirit, volume 1219 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/esycm10.zip.

Balzac:1998:AM

[1174] Honoré de Balzac. The Atheist’s Mass, volume 1220 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/athms10.zip.

Coulter:1998:NAS

[1175] John Merle Coulter. North American Species of Cactus, volume 1221 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/nasoc10.zip.

Cobb:1998:CA

[1176] Irvin S. (Irvin Shrewsbury) Cobb. Cobb’s Anatomy, volume 1222 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/canat10.zip.

Balzac:1998:U

[1177] Honoré de Balzac. Ursula, volume 1223 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/rsula10.zip.

Mims:1998:BSL

[1178] Edwin Mims. A Biography of Sidney Lanier, volume 1224 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/lanrb10.zip.

Tyndall:1998:FD

[1179] John Tyndall. Faraday As A Discoverer, volume 1225 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/fdayd10.zip.

Anonymous:1998:JAP

[1180] Anonymous. The Journal of Abnormal Psychology: Volume 10, volume 1226 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/jap1010.zip.

Darwin:1998:EEM

[1181] Charles Darwin. The Expression of Emotion in Man and Animals, volume 1227 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/eemaa10.zip.

Darwin:1998:OS

[1182] Charles Darwin. On the Origin of Species, volume 1228 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/otoof10.zip.

Lanier:1998:SPS

[1183] Sidney Lanier. Select Poems of Sidney Lanier, volume 1229 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/sposl10.zip.

Balzac:1998:PG

[1184] Honoré de Balzac. Pierre Grassou, volume 1230 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/prgrs10.zip.

Lawson:1998:T

[1185] Henry Lawson. On the Track, volume 1231 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/ontrk10.zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ontrk10h.zip.

Machiavelli:1998:PE

[1186] Nicolo Machiavelli. The Prince, etc., volume 1232 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/tprnc11.zip.

Crile:1998:ONE

[1187] George W. (Washington) Crile. Origin and Nature of Emotions, volume 1233 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/oanoe10.zip.

Conant:1998:OS

[1188] James Bryant Conant. Organic Syntheses, volume 1234 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/rgsyn10.zip.

Gautier:1998:CF

[1189] Theophile Gautier. Captain Fracasse, volume 1235 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/cptnf10.zip.

Brooks:1998:FAC

[1190] Noah Brooks. First Across the Continent, volume 1236 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/landc10.zip.

Balzac:1998:FG

[1191] Honoré de Balzac. Father Goriot, volume 1237 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/frgrt10.zip.

Thomson:1998:CDN

[1192] James Thomson. The City of Dreadful Night, volume 1238 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ctdnt10.zip.

Grey:1998:SB

[1193] Zane Grey. The Spirit of the Border, volume 1239 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/sprtb10.zip.

Synge:1998:PWW

[1194] John Millington Synge. The Playboy of the Western World, volume 1240 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/potww10.zip.

Synge:1998:WS

[1195] John Millington Synge. The Well of the Saints, volume 1241 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/welst10.zip.

Balzac:1998:UC

[1196] Honoré de Balzac. Unconscious Comedians, volume 1242 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/nccmd10.zip.

Meynell:1998:HC

[1197] Alice Meynell. Hearts of Controversy, volume 1243 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/hrtcn10.zip.

Congreve:1998:LL

[1198] William Congreve. Love for Love, volume 1244 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/lv4lv10.zip.

Woolf:1998:ND

[1199] Virginia Woolf. Night and Day, volume 1245 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/niday10.zip.

Aiken:1998:HD

[1200] Conrad Potter Aiken. The House of Dust, volume 1246 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/hdust10.zip.

Millay:1998:SA

[1201] Edna St. Vincent Millay. Second April, volume 1247 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/april10.zip.

Wetmore:1998:LGS

[1202] Helen Cody Wetmore. Last of the Great Scouts, volume 1248 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/bbill10.zip.

Rand:1998:ACA

[1203] Ayn Rand. Anthem: Comparison of anthm10 and 10a, volume 1249 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/anthm10z.zip.

Rand:1998:A

[1204] Ayn Rand. Anthem, volume 1250 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/anthm10.
zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/
anthm10a.zip.

Malory:1998:MD

[1205] Thomas Malory. Le Mort d’Arthur, volume 1251–1252 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/1mart10.zip   (Volume 1);   ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext98/2mart10.zip (Volume
 2).

Flaubert:1998:SS

[1206] Gustave Flaubert. A Simple Soul, volume 1253 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/sseng10.zip.

Rostand:1998:CBa

[1207] Edmond Rostand. Cyrano de Bergerac, volume 1254 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/cdben10.zip.

Rostand:1998:CBb

[1208] Edmond Rostand. Cyrano de Bergerac, volume 1255 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/cdbfr10h.zip.

Rostand:1998:CBc

[1209] Edmond Rostand. Cyrano de Bergerac, volume 1256 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/cdbfr10.zip.

Dumas:1998:TM

[1210] Alexandre Dumas. The Three Musketeers, volume 1257 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/1musk10.zip.

Dumas:1998:TYL

[1211] Alexandre Dumas. Ten Years Later, volume 1258 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2musk10.zip.

Dumas:1998:TYA

[1212] Alexandre Dumas. Twenty Years After, volume 1259 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/3musk10.zip.

Bronte:1998:JE

[1213] Charlotte Brontë. Jane Eyre, volume 1260 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/janey10.zip.

Grey:1998:BZ

[1214] Zane Grey. Betty Zane, volume 1261 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/bzane10.
zip.

Grey:1998:HD

[1215] Zane Grey. Heritage of the Desert, volume 1262 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/hdsrt10.zip.

Wharton:1998:GM

[1216] Edith Wharton. The Glimpses of the Moon, volume 1263 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/tgotm10.zip.

Wells:1998:WCB

[1217] H. G. (Herbert George) Wells. Wheels of Chance and Bicycling Idyll, volume 1264 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/wchnc10.zip.

Strachey:1998:QV

[1218] Lytton Strachey. Queen Victoria, volume 1265 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/qvctr10.zip.

Reed:1998:LOL

[1219] Myrtle Reed. Lavender and Old Lace, volume 1266 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/lvolc10.zip.

Bramah:1998:KLG

[1220] Ernest Bramah. Kai Lung’s Golden Hours, volume 1267 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/klsgh10.zip.

Verne:1998:MI

[1221] Jules Verne. The Mysterious Island, volume 1268 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/milnd10.zip.

Wells:1998:SB

[1222] H. G. (Herbert George) Wells. Soul of a Bishop, volume 1269 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/sbshp10.zip.

Mencken:1998:DW

[1223] H. L. (Henry Louis) Mencken. In Defense of Women, volume 1270 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/ndwmn11.zip.

Redgrove:1998:BB

[1224] H. Stanley (Herbert Stanley) Redgrove. Bygone Beliefs, volume 1271 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/byblf10.zip.

LeFebure:1998:CSP

[1225] Victor LeFebure. Chemical Strategy in Peace and War, volume 1272 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/rrhin10.zip.

LeFebure:1998:RR

[1226] Victor LeFebure. The Riddle of the Rhine, volume 1272 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/rrhin10.zip.

Lyall:1998:AS

[1227] Edna Lyall. The Autobiography of a Slander, volume 1273 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/autos10.zip.

Masefield:1998:MHD

[1228] John Masefield. Martin Hyde the Duke’s Messenger, volume 1274 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/mhyde10.zip.

Kingsley:1998:AHS

[1229] Charles Kingsley. Alexandria and her Schools, volume 1275 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/alxsc10.zip.

Meynell:1998:RL

[1230] Alice Meynell. The Rhythm of Life, volume 1276 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/rhyml10.zip.

Balzac:1998:MR

[1231] Honoré de Balzac. Melmoth Reconciled, volume 1277 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mlmth10.zip.

Wiggin:1998:PEE

[1232] Kate Douglas Smith Wiggin. Penelope’s English Experiences, volume 1278 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/penex10.zip.

Burns:1998:PSR

[1233] Robert Burns. Poems and Songs of Robert Burns, volume 1279 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/psorb10.zip.

Masters:1998:SRA

[1234] Edgar Lee Masters. Spoon River Anthology, volume 1280 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/sprvr10.zip.

Appleton:1998:TSHa

[1235] Victor Appleton. Tom Swift and His Aerial Warship, volume 1281 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/18tom10.zip.

Appleton:1998:TSAa

[1236] Victor Appleton. Tom Swift Among the Diamond Makers, volume 1282 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/07tom10.zip.

Appleton:1998:TSHb

[1237] Victor Appleton. Tom Swift and His Wizard Camera, volume 1283 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/14tom10.zip.

Appleton:1998:TSHc

[1238] Victor Appleton. Tom Swift and His Air Scout, volume 1284 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/22tom10.zip.

Butler:1998:WGA

[1239] Ellis Parker Butler. The Water Goats et al., volume 1285 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/twgts10.zip.

Lamb:1998:TSb

[1240] Charles Lamb. Tales of Shakespeare, volume 1286 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/tshak10a.zip.

Goethe:1998:PG

[1241] Johann Wolfgang von Goethe. The Poems of Goethe, volume 1287 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/tpgth10.zip.

Grahame:1998:DD

[1242] Kenneth Grahame. Dream Days, volume 1288 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/drday10a.zip.

Dickens:1998:TGS

[1243] Charles Dickens. Three Ghost Stories, volume 1289 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/3ghst10.zip.

Flaubert:1998:S

[1244] Gustave Flaubert. Salammbo, volume 1290 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/slmmb10.zip.

Flaubert:1998:H

[1245] Gustave Flaubert. Herodias, volume 1291 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/hrods10.zip.

Congreve:1998:WW

[1246] William Congreve. The Way of the World, volume 1292 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/wwrld10.zip.

Ruskin:1998:SL

[1247] John Ruskin. Sesame and Lilies, volume 1293 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/sesli10.zip.

Balzac:1998:FN

[1248] Honoré de Balzac. The Firm of Nucingen, volume 1294 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ncngn10.zip.

Meynell:1998:CR

[1249] Alice Meynell. Cérès’ Runaway, volume 1295 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/crnwy10.zip.

Galt:1998:P

[1250] John Galt. The Provost, volume 1296 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/prvst10.
zip.

Davis:1998:IP

[1251] James J. (James John) Davis. The Iron Puddler, volume 1297 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/tirnp10.zip.

Wister:1998:VHP

[1252] Owen Wister. The Virginian, Horseman Of The Plains, volume 1298 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/vrgnn10.zip.

Bower:1998:HS

[1253] B. M. Bower. The Heritage of the Sioux, volume 1299 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/hrtsu10.zip.

Grey:1998:RPS

[1254] Zane Grey. Riders of the Purple Sage, volume 1300 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/prpsg10.zip.

Carlyle:1998:FR

[1255] Thomas Carlyle. The French Revolution, volume 1301 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/frrev10.zip.

Blades:1998:EB

[1256] William Blades. Enemies of Books, volume 1302 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/nmybk10.zip.

Caine:1998:S

[1257] Sir Hall Caine. The Scapegoat, volume 1303 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/scpgt10.zip.

Anonymous:1998:PGB

[1258] Anonymous. Project Gutenberg’s Book of English Verse, volume 1304 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/pgbev10.zip.

Balzac:1998:BAS

[1259] Honoré de Balzac. The Ball at Sceaux, volume 1305 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/blsco10.zip.

Beerbohm:1998:SM

[1260] Sir Max Beerbohm. Seven Men, volume 1306 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/svnmn10.zip.

Balzac:1998:MS

[1261] Honoré de Balzac. The Magic Skin, volume 1307 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/mgcsk10.zip.

Wilde:1998:OWM

[1262] Oscar Wilde. Oscar Wilde Miscellaneous, volume 1308 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/wldms10.zip.

Meynell:1998:SPA

[1263] Alice Meynell. The Spirit of Place, et al., volume 1309 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/sptpl10.zip.

Galt:1998:AP

[1264] John Galt. The Annals of the Parish, volume 1310 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/anapr10.zip.

Dunsany:1998:I

[1265] Lord Dunsany. If, volume 1311 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/ifdun10.
zip.

Harte:1998:SS

[1266] Bret Harte. Selected Stories, volume 1312 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/harte10.zip.

Lawson:1998:S

[1267] Henry Lawson. Over The Sliprails, volume 1313 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/oslip10.zip.

Oppenheim:1998:M

[1268] E. Phillips (Edward Phillips) Oppenheim. The Malefactor, volume 1314 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/mlfct11.zip.

Huxley:1998:ASE

[1269] Thomas Henry Huxley. Autobiography and Selected Essays, volume 1315 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/aseth10.zip.

Conrad:1998:SR

[1270] Joseph Conrad. Some Reminiscences, volume 1316 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/rmnis10.zip.

Paterson:1998:SBJ

[1271] A. B. (Andrew Barton) Paterson. Saltbush Bill J.P., volume 1317 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/biljp10.zip.

Reynolds:1998:TH

[1272] John N. Reynolds. The Twin Hells, volume 1318 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/twnhl10.zip.

Scott:1998:IEB

[1273] Walter Dill Scott. Increasing Efficiency In Business, volume 1319 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/ihdib10.zip.

Gross:1998:CP

[1274] Hans Gustav Adolf Gross. Criminal Psychology, volume 1320 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/crmsy10.zip.

Eliot:1998:WL

[1275] T. S. (Thomas Stearns) Eliot. The Waste Land, volume 1321 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/wslnd11.zip.

Whitman:1998:LG

[1276] Walt Whitman. Leaves of Grass, volume 1322 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/lvgrs10.zip.

Prescott:1998:HCPb

[1277] William Hickling Prescott. History Of The Conquest Of Peru, volume 1323 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/hcpru10a.zip.

Ransome:1998:R

[1278] Arthur Ransome. Russia in 1919, volume 1324 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/19rus10.zip.

Addams:1998:TYA

[1279] Jane Addams. Twenty Years At Hull House, volume 1325 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/20yhh10.zip.

Ransome:1998:CR

[1280] Arthur Ransome. The Crisis in Russia, volume 1326 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/crrus10.zip.

Elizabeth:1998:EHG

[1281] “Elizabeth”. Elizabeth and her German Garden, volume 1327 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/lzgdn10.zip.

Synge:1998:TW

[1282] John Millington Synge. The Tinker’s Wedding, volume 1328 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/tnkwd10.zip.

Lindsay:1998:VA

[1283] David Lindsay. A Voyage to Arcturus, volume 1329 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/vrctr10.zip.

Bannerman:1998:TS

[1284] Helen Bannerman. Two Stories, volume 1330 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/sambo10.zip.

Oliver:1998:AS

[1285] Charles A. (Charles Alexander) Oliver. ABC’s of Science, volume 1331 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/abcos10.zip.

Barrie:1998:PPK

[1286] J. M. (James Matthew) Barrie. Peter Pan in Kensington Gardens, volume 1332 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/ppikg10.zip.

Lang:1998:RFM

[1287] Andrew Lang. R. F. Murray: His Poems with a Memoir, volume 1333 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/rfmur10.zip.

Jerome:1998:PK

[1288] Jerome K. Jerome, (Jerome Klapka). Paul Kelver, volume 1334 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/pklvr10.zip.

Kingsley:1998:AR

[1289] Charles Kingsley. The Ancien Regime, volume 1335 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/anrgm10.zip.

Thompson:1998:S

[1290] Francis Thompson. Shelley, volume 1336 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/tshly10.zip.

Waterlow:1998:S

[1291] Sydney Philip Perigal Waterlow. Shelley, volume 1337 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/wshly10.zip.

Wilde:1998:SPO

[1292] Oscar Wilde. Selected Prose of Oscar Wilde, volume 1338 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/slpwl10.zip.

Wilde:1998:Sa

[1293] Oscar Wilde. Salome, volume 1339 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/salme10.
zip.

Wilde:1998:Sb

[1294] Oscar Wilde. Salome, volume 1340 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. Projected HTML version, not yet available.

Cardozo:1998:AP

[1295] Benjamin N. (Benjamin Nathan) Cardozo. The Altruist in Politics, volume 1341 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/ltplt10.zip.

Austen:1998:PP

[1296] Jane Austen. Pride and Prejudice, volume 1342 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/pandp10.zip.

Balzac:1998:B

[1297] Honoré de Balzac. Bureaucracy, volume 1343 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/brcrc10.zip.

Balzac:1998:SPC

[1298] Honoré de Balzac. Secrets of the Princesse de Cadignan, volume 1344 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/sdpdc10.zip.

Balzac:1998:VT

[1299] Honoré de Balzac. The Vicar of Tours, volume 1345 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/vcrtr10.zip.

Marx:1998:EBL

[1300] Karl Marx. Eighteenth Brumaire of Louis Bonaparte, volume 1346 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/mar1810.zip.

Roy:1998:NPH

[1301] Édouard le Roy. A New Philosophy: Henri Bergson, volume 1347 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/anphb10.zip.

McCarter:1998:MD

[1302] Margaret Hill McCarter. A Master’s Degree, volume 1348 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/amsdg10.zip.

Wallace:1998:R

[1303] Sir Donald Mackenzie Wallace. Russia, volume 1349 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/rsdmw10.zip.

Balzac:1998:CD

[1304] Honoré de Balzac. The Country Doctor, volume 1350 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ctrdr10.zip.

Trueman:1998:CIF

[1305] Howard Trueman. Chignecto Isthmus; First Settlers, volume 1351 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/chgnt1-.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext98/chgnt10.zip.

Balzac:1998:OM

[1306] Honoré de Balzac. An Old Maid, volume 1352 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/omaid10.zip.

Verne:1998:C

[1307] Jules Verne. Off on a Comet, volume 1353 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/comet10.zip.

Montgomery:1998:CA

[1308] L. M. (Lucy Maud) Montgomery. Chronicles of Avonlea, volume 1354 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/avnle08.zip.

Verne:1998:UC

[1309] Jules Verne. The Underground City, volume 1355 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/ucity10.zip.

Bullen:1998:CC

[1310] Frank T. Bullen. The Cruise of the Cachalot, volume 1356 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/cchlt10.zip.

Balzac:1998:MF

[1311] Honoré de Balzac. Madame Firmiani, volume 1357 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/frmni10.zip.

Tennyson:1998:EAC

[1312] Alfred Tennyson. Enoch Arden, and c., volume 1358 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/enchr11.zip.

Nield:1998:BHN

[1313] Jonathan Nield. Best Historical Novels and Tales, volume 1359 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/bhnat10.zip.

Kingsley:1998:HLE

[1314] Charles Kingsley. Historical Lectures and Essays, volume 1360 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/hstle10.zip.

Appleton:1998:TSHd

[1315] Victor Appleton. Tom Swift And His Giant Cannon, volume 1361 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/16tom10.zip.

Appleton:1998:TSHe

[1316] Victor Appleton. Tom Swift And His Undersea Search, volume 1362 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/23tom10.zip.

Appleton:1998:TSAb

[1317] Victor Appleton. Tom Swift Among The Fire Fighters, volume 1363 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/24tom10.zip.

Appleton:1998:TSHf

[1318] Victor Appleton. Tom Swift And His Electric Locomotive, volume 1364 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/25tom10.zip.

Longfellow:1998:CPH

[1319] Henry Wadsworth Longfellow. The Complete Poems of Henry Wadsworth Longfellow, volume 1365 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/cphwl10.
zip.

Reade:1998:CH

[1320] Charles Reade. The Cloister and the Hearth, volume 1366 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/chrth10.zip.

Ramee:1998:F

[1321] Louise de la Ramee. Findelkind, volume 1367 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/fndkd10.zip.

Haggard:1998:WWS

[1322] H. Rider Haggard. When the World Shook, volume 1368 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/wtwsh10.zip.

Balzac:1998:Pb

[1323] Honoré de Balzac. Paz, volume 1369 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/pzhdb10.
zip.

White:1998:AAD

[1324] Andrew Dickson White. Autobiography of Andrew Dickson White, volume 1370 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/1aadw10.zip   (Volume I)
;   ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/
2aadw10.zip (Volume   II).

Irving:1998:A

[1325] Washington Irving. Astoria, volume 1371 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/stria10.zip.

Irving:1998:ACB

[1326] Washington Irving. The Adventures of Captain Bonneville, volume 1372 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/taocb10.zip.

Balzac:1998:SW

[1327] Honoré de Balzac. Study of a Woman, volume 1373 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/sowmn10.zip.

Balzac:1998:Va

[1328] Honoré de Balzac. Vendetta, volume 1374 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/vndta10.zip.

Wiggin:1998:NCR

[1329] Kate Douglas Smith Wiggin. New Chronicles of Rebecca, volume 1375 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/tncor10.zip.

Barrie:1998:LWB

[1330] J. M. (James Matthew) Barrie. The Little White Bird, volume 1376 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/tlwbd10.zip.

Scott:1998:T

[1331] Sir Walter Scott. The Talisman, volume 1377 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/tlsmn10.zip.

White:1998:LF

[1332] Stewart Edward White. The Land of Footprints, volume 1378 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/fprnt10.zip.

Wister:1998:SD

[1333] Owen Wister. A Straight Deal, volume 1379 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/strdl10.zip.

Balzac:1998:TB

[1334] Honoré de Balzac. The Two Brothers, volume 1380 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/brthr10.zip.

Meredith:1998:P

[1335] George Meredith. Poems, volume 1381–1383 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/pmgm110.zip   (Volume 1);   ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/etext98/pmgm210.zip (Volume   2); ftp:
//uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/pmgm310.
zip   (Volume 3).

Galt:1998:AL

[1336] John Galt. The Ayrshire Legatees, volume 1384 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/ayrlg.zip;   ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ayrlg10.zip.

Wister:1998:LM

[1337] Owen Wister. Lin McLean, volume 1385 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/lmcln10.zip.

Wister:1998:LB

[1338] Owen Wister. Lady Baltimore, volume 1386 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/lbalt10.zip.

Wister:1998:M

[1339] Owen Wister. Mother, volume 1387 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/mothr10.
zip.

Wister:1998:PI

[1340] Owen Wister. Padre Ignacio, volume 1388 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/ignco10.zip.

Balzac:1998:Ga

[1341] Honoré de Balzac. Gobseck, volume 1389 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/gbsek10.zip.

Wister:1998:JBO

[1342] Owen Wister. The Jimmyjohn Boss and Other Stories, volume 1390 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/jmyjn10.zip.

Wiggin:1998:PIE

[1343] Kate Douglas Smith Wiggin. Penelope’s Irish Experiences, volume 1391 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/pnlie10.zip.

Dickens:1998:SPT

[1344] Charles Dickens. The Seven Poor Travellers, volume 1392 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/svprt10.zip.

Clough:1998:AV

[1345] Arthur Hugh Clough. Amours de Voyage, volume 1393 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mrvyg10.zip.

Dickens:1998:H

[1346] Charles Dickens. The Holly-Tree, volume 1394 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/hlytr10.zip.

Lang:1998:LL

[1347] Andrew Lang. Letters on Literature, volume 1395 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/ltlit10.zip.

Lytton:1998:RLR

[1348] Edward Bulwer Lytton Lytton, Baron. Rienzi, the last of the Roman tribunes, volume 1396 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext98/rienz10.zip.

Volney:1998:R

[1349] C. F. (Constantin François de) Volney. The Ruins, volume 1397 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ruins10.zip.

Troward:1998:DLM

[1350] T. (Thomas) Troward. Dore Lectures on Mental Science, volume 1398 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/dorel10.zip.

Tolstoy:1998:AK

[1351] Count Leo Nikoleyevich Tolstoy. Anna Karenina, volume 1399 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/nkrnn10.zip.

Dickens:1998:GE

[1352] Charles Dickens. Great Expectations, volume 1400 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/grexp10.zip.

Burroughs:1998:TU

[1353] Edgar Rice Burroughs. Tarzan the Untamed, volume 1401 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/tarz710.zip.

Morley:1998:WBB

[1354] Christopher Morley. Where the Blue Begins, volume 1402 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/wtbbg10.zip.

Balzac:1998:SL

[1355] Honoré de Balzac. A Start in Life, volume 1403 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/stlif10.zip.

Hamilton:1998:FP

[1356] Alexander Hamilton. The Federalist Papers, volume 1404 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/feder10a.zip.

Balzac:1998:CA

[1357] Honoré de Balzac. The Collection of Antiquities, volume 1405 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/clntq10.zip.

Dickens:1998:PCE

[1358] Charles Dickens. Perils of Certain English Prisoners, volume 1406 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/pocep10.zip.

Dickens:1998:MS

[1359] Charles Dickens. A Message From the Sea, volume 1407 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/amfts10.zip.

White:1998:NHS

[1360] Gilbert White. The Natural History of Selborne, volume 1408 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/tnhos10.zip.

Lowell:1998:SFE

[1361] Percival Lowell. The Soul of the Far East, volume 1409 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/sofre10.zip.

Balzac:1998:CL

[1362] Honoré de Balzac. The Commission in Lunacy, volume 1410 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/lunac10.zip.

Balzac:1998:DP

[1363] Honoré de Balzac. Domestic Peace, volume 1411 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/dmspc10.zip.

Marryat:1998:MR

[1364] Frederick Marryat. Masterman Ready, volume 1412 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/mmrdy10.zip.

Dickens:1998:TTG

[1365] Charles Dickens. Tom Tiddler’s Ground, volume 1413 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ttgnd10.zip.

Dickens:1998:SL

[1366] Charles Dickens. Somebody’s Luggage, volume 1414 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/smlgg10.zip.

Dickens:1998:DM

[1367] Charles Dickens. Doctor Marigold, volume 1415 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/drmrg10.zip.

Dickens:1998:MLLa

[1368] Charles Dickens. Mrs. Lirriper’s Lodgings, volume 1416 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mlldg10.zip.

Balzac:1998:SS

[1369] Honoré de Balzac. Sons of the Soil, volume 1417 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/ssoil10.zip.

Graves:1998:CS

[1370] Robert Ranke Graves. Country Sentiment, volume 1418 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/csent10.zip.

Dickens:1998:MJ

[1371] Charles Dickens. Mugby Junction, volume 1419 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/mgjnc10.zip.

Holmes:1998:LU

[1372] Thomas Holmes. London’s Underworld, volume 1420 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/lndwd10.zip.

Dickens:1998:MLLb

[1373] Charles Dickens. Mrs. Lirriper’s Legacy, volume 1421 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mlrlg10.zip.

Dickens:1998:GS

[1374] Charles Dickens. Going into Society, volume 1422 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/gisoc10.zip.

Dickens:1998:NT

[1375] Charles Dickens. No Thoroughfare, volume 1423 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/notho10.zip.

Edgeworth:1998:CR

[1376] Maria Edgeworth. Castle Rackrent, volume 1424 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/rkrnt10.zip.

Balzac:1998:Vb

[1377] Honoré de Balzac. El Verdugo, volume 1425 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/vrdug10.zip.

Balzac:1998:R

[1378] Honoré de Balzac. The Recruit, volume 1426 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/recrt10.zip.

Balzac:1998:DS

[1379] Honoré de Balzac. A Drama on the Seashore, volume 1427 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/seshr10.zip.

Balzac:1998:Gb

[1380] Honoré de Balzac. La Grenadiere, volume 1428 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/grndr10.zip.

Mansfield:1998:GPa

[1381] Katherine Mansfield. The Garden Party, volume 1429 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/gprty10.zip.

Nesbit:1998:BSS

[1382] E. (Edith) Nesbit. Beautiful Stories from Shakespeare, volume 1430 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/bsshk10.zip.

Schreiner:1998:TPH

[1383] Olive Schreiner. Trooper Peter Halket, volume 1431 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/trptr10.zip.

Balzac:1998:S

[1384] Honoré de Balzac. Seraphita, volume 1432 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/sraph10.zip.

Balzac:1998:RI

[1385] Honoré de Balzac. The Red Inn, volume 1433 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/rdinn10.zip.

Meynell:1998:E

[1386] Alice Meynell. Essays, volume 1434 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/esyam10.
zip.

Dickens:1998:MP

[1387] Charles Dickens. Miscellaneous Papers, volume 1435 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mspcd10.zip.

Lobo:1998:VAC

[1388] Jeronymo Lobo. A Voyage to Abyssinia /—cby Father Jerome Lobo, volume 1436 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/vygab10.zip.

Balzac:1998:J

[1389] Honoré de Balzac. Juana, volume 1437 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/juana10.zip.

Collins:1998:NN

[1390] Wilkie Collins. No Name, volume 1438 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/nname10.zip.

Schreiner:1998:D

[1391] Olive Schreiner. Dreams, volume 1439 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/drmos10.
zip.

Schreiner:1998:WL

[1392] Olive Schreiner. Woman and Labour, volume 1440 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/wmlbr10.zip.

Schreiner:1998:SAF

[1393] Olive Schreiner. The Story of an African Farm, volume 1441 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/aafrm10.zip.

Oppenheim:1998:KB

[1394] E. Phillips (Edward Phillips) Oppenheim. The Kingdom of the Blind, volume 1442 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/kblnd10.zip.

Balzac:1998:TP

[1395] Honoré de Balzac. Two Poets, volume 1443 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2poet10.zip.

Henry:1998:VC

[1396] O. Henry. The Voice of the City, volume 1444 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/vccty10.zip.

Zerbe:1998:A

[1397] James Slough Zerbe. Aeroplanes, volume 1445 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/aeroz10.zip.

Stewart:1998:PB

[1398] Donald Ogden Stewart. Perfect Behavior, volume 1446 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/pbhvr10.zip.

Oppenheim:1998:IP

[1399] E. Phillips (Edward Phillips) Oppenheim. The Illustrious Prince, volume 1447 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/iprnc10.zip.

Spyri:1998:H

[1400] Johanna Spyri. Heidi, volume 1448 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/heidi10.
zip.

London:1998:VM

[1401] Jack London. The Valley of the Moon, volume 1449 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/vlymn10.zip.

Porter:1998:P

[1402] Eleanor H. (Eleanor Hodgman) Porter. Pollyanna, volume 1450 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/plyna10.zip.

Tilden:1998:ALT

[1403] William (Bill) Tatem Tilden. The Art of Lawn Tennis, volume 1451 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/tenis10.zip.

Plaatje:1998:NLS

[1404] Sol Plaatje. Native Life in South Africa, volume 1452 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/nlisa10.zip.

Balzac:1998:Aa

[1405] Honoré de Balzac. The Alkahest, volume 1453 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/lkhst10.zip.

Balzac:1998:MCa

[1406] Honoré de Balzac. Maitre Cornelius, volume 1454 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/crnls10.zip.

Balzac:1998:HS

[1407] Honoré de Balzac. The Hated Son, volume 1455 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/htdsn10.zip.

Balzac:1998:EUT

[1408] Honoré de Balzac. An Episode Under the Terror, volume 1456 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/aeutt10.zip.

Sabatini:1998:MW

[1409] Rafael Sabatini. Mistress Wilding, volume 1457 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/wldng10.zip.

Schreiner:1998:DLR

[1410] Olive Schreiner. Dream Life and Real Life, volume 1458 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/dlarl10.zip.

Eliot:1998:POO

[1411] T. S. (Thomas Stearns) Eliot. Prufrock and Other Observations, volume 1459 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/prfrk10.zip.

Scott:1998:BD

[1412] Sir Walter Scott. The Black Dwarf, volume 1460 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/bdwrf10.zip.

Scott:1998:LM

[1413] Sir Walter Scott. A Legend of Montrose, volume 1461 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mntrs10.zip.

Thackeray:1998:SRP

[1414] William Makepeace Thackeray. Some Roundabout Papers, volume 1462 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/rndbt10.zip.

Gissing:1998:PPH

[1415] George Gissing. The Private Papers of Henry Ryecroft, volume 1463 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/ppohr10.zip.

Dickens:1998:CAY

[1416] Charles Dickens. Contributions to: All The Year Round, volume 1464 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/allyr10.zip.

Dickens:1998:WGM

[1417] Charles Dickens. Wreck of the Golden Mary, volume 1465 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/wrkgm10.zip.

Gorky:1998:COW

[1418] Maxim Gorky. Creatures That Once Were Men, and other stories, volume 1466 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/crmen10a.zip.

Dickens:1998:SCS

[1419] Charles Dickens. Some Christmas Stories, volume 1467 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/cdscs10.zip.

Macaulay:1998:HEA

[1420] Thomas Babington Macaulay Macaulay, Baron. The History of England from the Accession of James II, Vol. 1, volume 1468 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/1hoej10.zip.

Thompson:1998:P

[1421] Francis Thompson. Poems, volume 1469 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/pmbft10.zip.

Keene:1998:FME

[1422] H. G. Keene. The Fall of the Moghul Empire of Hindustan, volume 1470 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/tfmeh10.zip.

Thompson:1998:NP

[1423] Francis Thompson. New Poems, volume 1471 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/npbft10.zip.

Mansfield:1998:GPb

[1424] Katherine Mansfield. In a German Pension, volume 1472 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/inagp10.zip.

Edgeworth:1998:A

[1425] Maria Edgeworth. The Absentee, volume 1473 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/bsnte10.zip.

Balzac:1998:IG

[1426] Honoré de Balzac. The Illustrious Gaudissart, volume 1474 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/1gdsr10.zip.

Balzac:1998:GI

[1427] Honoré de Balzac. Gaudissart II, volume 1475 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2gdsr10.zip.

Conrad:1998:C

[1428] Joseph Conrad. Chance, volume 1476 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/chanc10.
zip.

Saki:1998:TP

[1429] Saki. The Toys of Peace, volume 1477 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/

etext98/toypc10.zip. The Toys of Peace – Louise – Tea – The Disappearance of Crispina Umberleigh – The Wolves of Cernogratz – Louis

– The Guests – The Penance – The Phantom Luncheon – A Bread and Butter Miss – Bertie’s Christmas Eve – Forewarned – The Interlopers – Quail Seed – Canossa – The Threat – Excepting Mrs. Pentherby – Mark

– The Hedgehog – The Mappined Life – Fate – The Bull – Morlvera – Shock Tactics – The Seven Cream Jugs – The Occasional Garden – The Sheep – The Oversight – Hyacinth – The Image of the Lost Soul – The Purple of the Balkan Kings – The Cupboard of the Yesterdays – For the Duration of the War.

Stewart:1998:POH

[1430] Donald Ogden Stewart. A Parody Outline of History, volume 1478 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/apooh10.zip.

CunninghameGraham:1998:VAB

[1431] R. B. (Robert Bontine) Cunninghame Graham. A Vanished Arcadia, being some account of the Jesuits in Paraguay, 1607 to 1767, volume 1479 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/vajip10.zip;   ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext98/vajip10h.zip.

Hughes:1998:TBS

[1432] Thomas Hughes. Tom Brown’s School Days, volume 1480 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/tbssd10.zip.

Balzac:1998:DE

[1433] Honoré de Balzac. A Daughter of Eve, volume 1481 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/doeve10.zip.

Balzac:1998:MM

[1434] Honoré de Balzac. Modeste Mignon, volume 1482 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/mdmgn10.zip.

Carvalho:1998:FCI

[1435] David N. Carvalho. Forty Centuries of Ink, volume 1483 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/40cnk10.zip.

BlascoIbanez:1998:FHA

[1436] Vicente Blasco Ibanez. The Four Horsemen of the Apocalypse, volume 1484 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/4hrsm10.zip.

Nation:1998:UNL

[1437] Carrie A. Nation. The Use and Need of the Life of Carrie A. Nation, volume 1485 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/crntn10.zip.

Fiske:1998:UWO

[1438] John Fiske. The Unseen World and Other Essays, volume 1486 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/nswoe10.zip.

Shaw:1998:PWC

[1439] George Bernard Shaw. The Perfect Wagnerite, Commentary on the Ring, volume 1487 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/sring10.zip.

Brooks:1998:TSC

[1440] Elbridge S. (Streeter) Brooks. The True Story of Christopher Columbus, Called the Great Admiral, volume 1488 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/ttscc10.zip.

Anonymous:1998:NMFa

[1441] Anonymous. The New McGuffey First Reader, volume 1489 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/1nmcg10.zip.

Anonymous:1998:NMFb

[1442] Anonymous. The New McGuffey Fourth Reader, volume 1490 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/4nmcg10.zip.

Lang:1998:LDA

[1443] Andrew Lang. Letters to Dead Authors, volume 1491 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/letda10.zip.

Hale:1998:LCC

[1444] Edward Everett Hale. The Life of Christopher Columbus from his own Letters and Journals, volume 1492 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/tlocc10.zip.

Ginzberg:1998:LJ

[1445] Louis Ginzberg. The Legends of the Jews, volume 1493–1494 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/1lotj10.zip   (v1); ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/etext98/2lotj10.zip   (v2).

Steele:1998:GCM

[1446] Chester K. Steele. The Golf Course Mystery, volume 1495 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/glfms10.zip.

Marlowe:1998:MAP

[1447] Christopher Marlowe. Massacre at Paris, volume 1496 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/msprs10.zip.

Plato:1998:R

[1448] Plato. The Republic, volume 1497 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/repub11.
zip.

Lord:1998:BLH

[1449] John Lord. Beacon Lights of History, volume 1498–1499 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/31blh10.zip   (v3.1); ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext98/32blh10.zip   (v3.2).

Shakespeare:1998:KHVa

[1450] William Shakespeare. King Henry VI, Part 1, volume 1500 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws0110.zip.

Shakespeare:1998:KHVb

[1451] William Shakespeare. King Henry VI, Part 2, volume 1501 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws0210.zip.

Shakespeare:1998:KHVc

[1452] William Shakespeare. King Henry VI, Part 3, volume 1502 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws0310.zip.

Shakespeare:1998:KRIa

[1453] William Shakespeare. King Richard III, volume 1503 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws0410.zip.

Shakespeare:1998:CE

[1454] William Shakespeare. The Comedy of Errors, volume 1504 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws0610.zip.

Shakespeare:1998:RLa

[1455] William Shakespeare. The Rape of Lucrece, volume 1505 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws0810.zip.

Shakespeare:1998:RLb

[1456] William Shakespeare. The Rape of Lucrece, volume 1506 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/3ws0810.zip.

Shakespeare:1998:TTA

[1457] William Shakespeare. The Tragedy of Titus Andronicus, volume 1507 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/2ws0910.zip.

Shakespeare:1998:TS

[1458] William Shakespeare. The Taming of the Shrew, volume 1508 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws1010.zip.

Shakespeare:1998:TGV

[1459] William Shakespeare. Two Gentlemen of Verona, volume 1509 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws1110.zip.

Shakespeare:1998:LLL

[1460] William Shakespeare. Love’s Labour’s Lost, volume 1510 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws1210.zip.

Shakespeare:1998:KJ

[1461] William Shakespeare. King John, volume 1511 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws1410.zip.

Shakespeare:1998:KRIb

[1462] William Shakespeare. King Richard II, volume 1512 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws1510.zip.

Shakespeare:1998:RJ

[1463] William Shakespeare. Romeo and Juliet, volume 1513 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws1610.zip.

Shakespeare:1998:MND

[1464] William Shakespeare. A Midsummer Night’s Dream, volume 1514 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/2ws1710.zip.

Shakespeare:1998:MV

[1465] William Shakespeare. The Merchant of Venice, volume 1515 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws1810.zip.

Shakespeare:1998:KHIa

[1466] William Shakespeare. King Henry IV, Part 1, volume 1516 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws1910.zip.

Shakespeare:1998:MWW

[1467] William Shakespeare. The Merry Wives of Windsor, volume 1517 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/2ws2010.zip.

Shakespeare:1998:KHIb

[1468] William Shakespeare. King Henry IV, Part 2, volume 1518 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws2110.zip.

Shakespeare:1998:MAAa

[1469] William Shakespeare. Much Ado About Nothing, volume 1519 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws2210.zip.

Shakespeare:1998:MAAb

[1470] William Shakespeare. Much Ado About Nothing, volume 1520 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/3ws2210.zip.

Shakespeare:1998:KHVd

[1471] William Shakespeare. King Henry V, volume 1521 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws2310.zip.

Shakespeare:1998:JC

[1472] William Shakespeare. Julius Caesar, volume 1522 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws2410.zip.

Shakespeare:1998:YLI

[1473] William Shakespeare. As You Like It, volume 1523 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws2510.zip.

Shakespeare:1998:H

[1474] William Shakespeare. Hamlet, volume 1524 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws2610.zip.

Shakespeare:1998:PT

[1475] William Shakespeare. The Phoenix and the Turtle, volume 1525 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws2710.zip.

Shakespeare:1998:TNa

[1476] William Shakespeare. Twelfth Night, volume 1526 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws2810.zip.

Shakespeare:1998:TNb

[1477] William Shakespeare. Twelfth Night, volume 1527 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/3ws2810.zip.

Shakespeare:1998:TC

[1478] William Shakespeare. Troilus and Cressida, volume 1528 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws2910.zip.

Shakespeare:1998:AWE

[1479] William Shakespeare. All’s Well That Ends Well, volume 1529 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws3010.zip.

Shakespeare:1998:MM

[1480] William Shakespeare. Measure for Measure, volume 1530 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws3110.zip.

Shakespeare:1998:O

[1481] William Shakespeare. Othello, volume 1531 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws3210.zip.

Shakespeare:1998:KL

[1482] William Shakespeare. King Lear, volume 1532 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws3310.zip.

Shakespeare:1998:M

[1483] William Shakespeare. Macbeth, volume 1533 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws3410.zip.

Shakespeare:1998:AC

[1484] William Shakespeare. Antony and Cleopatra, volume 1534 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws3510.zip.

Shakespeare:1998:Ca

[1485] William Shakespeare. Coriolanus, volume 1535 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws3610.zip.

Shakespeare:1998:TA

[1486] William Shakespeare. Timon of Athens, volume 1536 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws3710.zip.

Shakespeare:1998:P

[1487] William Shakespeare. Pericles, volume 1537 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws3810.zip.

Shakespeare:1998:Cb

[1488] William Shakespeare. Cymbeline, volume 1538 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws3910.zip.

Shakespeare:1998:WT

[1489] William Shakespeare. The Winter’s Tale, volume 1539 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws4010.zip.

Shakespeare:1998:T

[1490] William Shakespeare. The Tempest, volume 1540 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/2ws4110.zip.

Shakespeare:1998:KHVe

[1491] William Shakespeare. King Henry VIII, volume 1541 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws4310.zip.

Shakespeare:1998:TNK

[1492] William Shakespeare. The Two Noble Kinsmen, Shakespeare Apocrypha, volume 1542 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/2ws4310.zip.

Shakespeare:1998:LC

[1493] William Shakespeare. A Lover’s Complaint, volume 1543 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws4410.zip.

Shakespeare:1998:PPa

[1494] William Shakespeare. The Passionate Pilgrim, volume 1544 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/2ws4510.zip.

Shakespeare:1998:PPb

[1495] William Shakespeare. The Passionate Pilgrim, volume 1545 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/3ws4510.zip.

Shakespeare:1998:S

[1496] William Shakespeare. Sonnets, volume 1546 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/1ws4610.zip.

Shakespeare:1998:STM

[1497] William Shakespeare. Sir Thomas More, Shakespeare Apocrypha, volume 1547 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/1ws4710.zip.

Shakespeare:1998:L

[1498] William Shakespeare. Locrine, volume 1548 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/1ws4810.zip.

Luther:1998:CG

[1499] Martin Luther. Commentary on Galatians, volume 1549 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mlglt10.zip.

Burnett:1998:LQ

[1500] Frances Hodgson Burnett. A Lady of Quality, volume 1550 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ladyq10.zip.

Wiggin:1998:CC

[1501] Kate Douglas Smith Wiggin. A Cathedral Courtship, volume 1551 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/cthrc10.zip.

Hawthorne:1998:MIS

[1502] Julian Hawthorne. The Most Interesting Stories of All Nations, volume 1552 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/misan10.zip.

Balzac:1998:HM

[1503] Honoré de Balzac. The Hidden Masterpiece, volume 1553 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/hmstp10.zip.

Balzac:1998:Ab

[1504] Honoré de Balzac. Adieu, volume 1554 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/adieu10.zip.

Balzac:1998:PD

[1505] Honoré de Balzac. A Passion in the Desert, volume 1555 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/apitd10.zip.

Balzac:1998:MCb

[1506] Honoré de Balzac. The Marriage Contract, volume 1556 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/mrgct10.zip.

Pyle:1998:MI

[1507] Ernie Howard Pyle. Men of Iron, volume 1557 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/femen10.zip.

Sinclair:1998:PRE

[1508] Upton Sinclair. The Profits of Religion: An Essay in Economic Interpretation, volume 1558 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext98/prfrl10.zip.

Morris:1998:SFC

[1509] Charles Morris. The San Francisco Calamity by Earthquake and Fire, volume 1560 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/sfclm10.zip.

Carpenter:1998:PCC

[1510] Edward Carpenter. Pagan and Christian Creeds: Their Origin and Meaning, volume 1561 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/pchrc10.
zip.

Dyke:1998:LRB

[1511] Henry van Dyke. Little Rivers: a book of essays in profitable idleness, volume 1562 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/ltrvs10.zip.

LeBlanc:1998:CS

[1512] Maurice LeBlanc. The Crystal Stopper, volume 1563 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/cstpr10.zip.

Boswell:1998:BLJ

[1513] James Boswell. Boswell’s Life of Johnson, volume 1564 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ljnsn10.zip.

Lytton:1998:LDP

[1514] Edward Bulwer Lytton Lytton, Baron. The Last Days of Pompeii, volume 1565 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/tldop10.zip.

Osler:1998:EMM

[1515] William Osler. The Evolution of Modern Medicine, volume 1566 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/teomm10.zip.

Eliot:1998:P

[1516] T. S. (Thomas Stearns) Eliot. Poems, volume 1567 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/tsepm10.zip.

Henley:1998:P

[1517] William Ernest Henley. Poems, volume 1568 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/pmweh10.zip.

Balzac:1998:LV

[1518] Honoré de Balzac. The Lily of the Valley, volume 1569 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/tlotv10.zip.

Dumont:1998:PC

[1519] Theron Q. Dumont. The Power of Concentration, volume 1570 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/prcon10.zip.

Plato:1998:Ca

[1520] Plato. Critias, volume 1571 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/criti10.
zip.

Plato:1998:T

[1521] Plato. Timaeus, volume 1572 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/tmeus10.
zip.

Alger:1998:FCF

[1522] Horatio Alger. Frank’s Campaign or the Farm and the Camp, volume 1573 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext98/frcmp10.zip.

Brooks:1998:HG

[1523] Elbridge S. (Streeter) Brooks. Historic Girls, volume 1574 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/hgrls10.zip.

Myerson:1998:FP

[1524] Abraham Myerson. The Foundations of Personality, volume 1575 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext98/prsty10.zip.

Benton:1998:LPB

[1525] Joel Benton. The Life of Phineas T. Barnum, volume 1576 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext98/ptbnm10.zip.

Phillpotts:1998:GR

[1526] Eden Phillpotts. The Grey Room, volume 1577 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/gryrm10.zip.

Lang:1998:AN

[1527] Andrew Lang. Aucassin and Nicolete, volume 1578 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1998. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext98/aucnc10.zip.

Plato:1998:La

[1528] Plato. Lysis, volume 1579 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/lysis10.
zip.

Plato:1998:Cb

[1529] Plato. Charmides, volume 1580 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/chmds10.
zip.

Anonymous:1998:HBDa

[1530] Anonymous. Holy Bible, Douay Rheims Version, Both Testaments, volume 1581 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/0drvb10.zip.

Anonymous:1998:HBDb

[1531] Anonymous. The Holy Bible, Douay Rheims Version, New Testament, volume 1582 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext98/3drvb10.zip.

Henry:1998:O

[1532] O. Henry. Options, volume 1583 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/optns10.
zip.

Plato:1998:Lb

[1533] Plato. Laches, volume 1584 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1998. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext98/lachs10.
zip.

Stevenson:1999:WB

[1534] Robert Louis Stevenson. The Wrong Box, volume 1585 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/wrngb10.zip.

Collins:1999:MW

[1535] Wilkie Collins. Man and Wife, volume 1586 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/mandw10.zip.

Collins:1999:BR

[1536] Wilkie Collins. The Black Robe, volume 1587 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/blkrb10.zip.

Collins:1999:RL

[1537] Wilkie Collins. A Rogue’s Life, volume 1588 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/rgslf10.zip.

Marlowe:1999:TGP

[1538] Christopher Marlowe. Tamburlaine the Great, part 2, volume 1589 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/tmbn210.zip.

Rinehart:1999:AI

[1539] Mary Roberts Rinehart. The Amazing Interlude, volume 1590 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/mzgnt10.zip.

Plato:1999:Pa

[1540] Plato. Protagoras, volume 1591 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/prtgs10.
zip.

McAfee:1999:SKJ

[1541] Cleland Boyd McAfee. Study of the King James Bible, volume 1592 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/sokjv10.zip.

Wood:1999:HTB

[1542] Robert Williams Wood. How to Tell the Birds from the Flowers, volume 1593 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/httbf10h.zip;   ftp://
uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/httbf10h/
httbf10h.htm.

Lang:1999:EL

[1543] Andrew Lang. Essays in Little, volume 1594 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/esltt10.zip.

Henry:1999:W

[1544] O. Henry. Whirligigs, volume 1595 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/whrlg10.
zip.

London:1999:SB

[1545] Jack London. Smoke Bellew, volume 1596 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/smkbl10.zip.

Andersen:1999:AFT

[1546] Hans Christian Andersen. Andersen’s Fairy Tales, volume 1597 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/hcaft10.zip.

Plato:1999:Ea

[1547] Plato. Euthydemus, volume 1598 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/uthyd10.
zip.

Anonymous:1999:CLG

[1548] Anonymous. Cinderella; or, The Little Glass Slipper, and other stories, volume 1599 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/cndrl10.zip.

Plato:1999:Sa

[1549] Plato. Symposium, volume 1600 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/sympo10.
zip.

Rinehart:1999:BP

[1550] Mary Roberts Rinehart. The Breaking Point, volume 1601 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/brkpt10.zip.

Ferber:1999:DOG

[1551] Edna Ferber. Dawn O’Hara, The Girl Who Laughed, volume 1602 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/dwnhr10.zip.

Dyke:1999:BF

[1552] Henry van Dyke. The Blue Flower, volume 1603 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/blflr10.zip.

Stevenson:1999:E

[1553] Robert Louis Stevenson. The Ebb-Tide, volume 1604 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ebtid10.zip.

Stephens:1999:CG

[1554] James Stephens. The Crock of Gold, volume 1605 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/crkgd10.zip.

Scott:1999:K

[1555] Sir Walter Scott. Kenilworth, volume 1606 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/knlwt10.zip.

Astor:1999:JOW

[1556] John Jacob Astor. A Journey in Other Worlds, volume 1607 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ajiow10.zip.

Dumas:1999:CDA

[1557] Alexandre Dumas. Camille (La Dame aux Camilias), volume 1608 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/cmlle10.zip.

Anonymous:1999:HBD

[1558] Anonymous. The Holy Bible, Douay-Rheims Version, O.T., volume 1609–1610 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/1drvb10.zip   (Part 1)
;   ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/
2drvb10.zip (Part   2).

Tarkington:1999:S

[1559] Booth Tarkington. Seventeen, volume 1611 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/svntn10.zip.

Conkling:1999:PLG

[1560] Hilda Conkling. Poems By a Little Girl, volume 1612 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/pbalg10.zip.

Clausen:1999:CB

[1561] J. Storer Clausen. Count Bunker, volume 1613 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/cbnkr10.zip.

Hawthorne:1999:GF

[1562] Julian Hawthorne. The Golden Fleece, volume 1614 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/gldfl10.zip.

Savage:1999:OEL

[1563] Ernest A. Savage. Old English Libraries: The Making, Collection, and Use of Books During the Middle Ages, volume 1615 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/nglbs10.zip.

Plato:1999:Ca

[1564] Plato. Cratylus, volume 1616 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/crtls10.
zip.

Wilkins:1999:WRO

[1565] Mary Wilkins. The Wind in the Rose-Bush, and Other Stories Of The Supernatural, volume 1617 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/sotsn10.
zip.

Synge:1999:SG

[1566] John Millington Synge. In Shadow of the Glen, volume 1618 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/sglen10.zip.

Rojas:1999:C

[1567] Fernando de Rojas. La Celestina, volume 1619 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/clstn10.zip.

Davis:1999:LU

[1568] Richard Harding Davis. Lion and the Unicorn, volume 1620 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/liuni10.zip.

Collins:1999:MM

[1569] Wilkie Collins. Miss or Mrs?, volume 1621 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/mioms10.zip.

Collins:1999:LL

[1570] Wilkie Collins. The Law and the Lady, volume 1622 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/lwldy10.zip.

Collins:1999:NM

[1571] Wilkie Collins. The New Magdalen, volume 1623 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/nmgdl10.zip.

Collins:1999:TD

[1572] Wilkie Collins. The Two Destinies, volume 1624 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/2dstn10.zip.

Collins:1999:FD

[1573] Wilkie Collins. The Frozen Deep, volume 1625 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/frzdp10.zip.

Collins:1999:AD

[1574] Wilkie Collins. After Dark, volume 1626 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/ftrdk10.zip.

Collins:1999:EG

[1575] Wilkie Collins. The Evil Genius, volume 1627 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/vlgns10.zip.

Collins:1999:MLM

[1576] Wilkie Collins. My Lady’s Money, volume 1628 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/mlmny10.zip.

Collins:1999:SN

[1577] Wilkie Collins. I Say No, volume 1629 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/isyan10.zip.

Collins:1999:LN

[1578] Wilkie Collins. Little Novels, volume 1630 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/lnvls10.zip.

Lang:1999:MF

[1579] Andrew Lang. A Monk of Fife, volume 1631 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/mnkff10.zip.

Whibley:1999:BS

[1580] Charles Whibley. A Book of Scoundrels, volume 1632 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/abkos10.zip. Captain Hind.–Moll Cutpurse and

Jonathan Wild.–Ralph Briscoe.– Gilderoy and Sixteen-String Jack.– Thomas Pureney.–Sheppard and Cartouche.–Vaux.–George Barrington.– The Switcher and Gentleman Harry.–Deacon Brodie and Charles Peace.– The Man in the grey suit.–Monsieur l’abbe.

Hale:1999:BMA

[1581] Edward Everett Hale. The Brick Moon, et al., volume 1633 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/brkmn10.zip.

Dixon:1999:FV

[1582] Thomas Dixon. The Foolish Virgin, volume 1634 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/fvrgn10.zip.

Plato:1999:I

[1583] Plato. Ion, volume 1635 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/ionbp10.
zip.

Plato:1999:Pb

[1584] Plato. Phaedrus, volume 1636 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/phdrs10.
zip.

Kingsley:1999:SSL

[1585] Charles Kingsley. Sanitary and Social Lectures and Essays, volume 1637 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/sasle10.zip. Woman’s Work in a Country

Parish – The Science of Health – The Two Breaths – Thrift – Nausicaa in London; or, the Lower Education of Women – The Air-Mothers – The Tree of Knowledge – Great Cities and their Influence for Good and Evil – Heroism – The Massacre of the Innocents – “A mad world, my masters.”.

ConanDoyle:1999:NR

[1586] Sir Arthur Conan Doyle. The New Revelation, volume 1638 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/nrvln10.zip.

Balzac:1999:ED

[1587] Honoré de Balzac. Eve and David, volume 1639 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/evdvd10.zip.

MacDonald:1999:L

[1588] George MacDonald. Lilith, volume 1640 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/lilth10.zip.

Balzac:1999:LB

[1589] Honoré de Balzac. The Lesser Bourgeoisie, volume 1641 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/lsbrg10.zip.

Plato:1999:Eb

[1590] Plato. Euthyphro, volume 1642 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/uthph10.
zip.

Plato:1999:Ma

[1591] Plato. Meno, volume 1643 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/1meno10.
zip.

ConanDoyle:1999:AG

[1592] Sir Arthur Conan Doyle. The Adventures of Gerard, volume 1644 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/agrrd10.zip.

Lang:1999:RM

[1593] Andrew Lang. Rhymes a la Mode, volume 1645 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/rmalm10.zip.

Henry:1999:RD

[1594] O. Henry. Roads of Destiny, volume 1646 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/rdstn10.zip.

Chesterton:1999:MWKa

[1595] G. K. (Gilbert Keith) Chesterton. The Man Who Knew Too Much, volume 1647 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/mwktm10.zip.

Chesterton:1999:TPa

[1596] G. K. (Gilbert Keith) Chesterton. The Trees of Pride, volume 1648 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/trprd10.zip.

Balzac:1999:F

[1597] Honoré de Balzac. Ferragus, volume 1649 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/frrgs10.zip.

Burgoyne:1999:LES

[1598] Thomas H. Burgoyne. The Light of Egypt, or the Science of the Soul and the Stars, Volume II, volume 1650 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/

etext99/2tloe10.zip. “Dictated by the author from the subjective plane of life” through Belle M. Wagner.

Gaboriau:1999:MO

[1599] Emile Gaboriau. The Mystery of Orcival, volume 1651 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/orcvl10.zip.

Verne:1999:SCa

[1600] Jules Verne. The Survivors of the Chancellor, volume 1652 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/tsotc10.zip.

Kempis:1999:IC

[1601] Thomas a Kempis. The Imitation of Christ, volume 1653 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/mcrst10.zip.

Shaw:1999:US

[1602] George Bernard Shaw. An Unsocial Socialist, volume 1654 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/unsoc10.zip.

London:1999:TK

[1603] Jack London. Tales of the Klondyke, volume 1655 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/klndk10.zip.

Plato:1999:Aa

[1604] Plato. Apology, volume 1656 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/pplgy10.
zip.

Plato:1999:Cb

[1605] Plato. Crito, volume 1657 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/crito10.
zip.

Plato:1999:Pc

[1606] Plato. Phaedo, volume 1658 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/phado10.
zip.

Balzac:1999:GGE

[1607] Honoré de Balzac. The Girl with the Golden Eyes, volume 1659 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/gwtgi10.zip.

Balzac:1999:SCL

[1608] Honoré de Balzac. Scenes from a Courtesan’s Life, volume 1660 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/sfacl10.zip.

ConanDoyle:1999:ASH

[1609] Sir Arthur Conan Doyle. The Adventures of Sherlock Holmes, volume 1661 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext99/advsh11.zip.

Anonymous:1999:CWF

[1610] Anonymous. The 1997 CIA World Factbook, volume 1662 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/world97.zip.

Foster:1999:WMS

[1611] H. D. Foster. Webster’s March 7th Speech, volume 1663 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/wsm7s10.zip.

Farrar:1999:SP

[1612] John Farrar. Songs for Parents, volume 1664 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/sfpar10.zip.

Lyall:1999:DV

[1613] Edna Lyall. Derrick Vaughan–Novelist, volume 1665 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/dvnvl10.zip.

Africanus:1999:GA

[1614] Lucius Apuleius “Africanus”. The Golden Asse, volume 1666 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/gldns10.zip.

Scott:1999:MAM

[1615] Sir Walter Scott. My Aunt Margaret’s Mirror, volume 1667 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/mamsm10.zip.

Scott:1999:DLJ

[1616] Sir Walter Scott. Death of the Laird’s Jock, volume 1668 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/tpsch10.zip.

Scott:1999:TC

[1617] Sir Walter Scott. The Tapestried Chamber, volume 1668 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/tpsch10.zip.

London:1999:HDO

[1618] Jack London. The Human Drift, and other stories, volume 1669 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/hmndr10.zip.

Luther:1999:MLS

[1619] Martin Luther. Martin Luther’s Small Catechism, Trns. by R. Smith, volume 1670 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/smlct10.zip.

Rinehart:1999:WMM

[1620] Mary Roberts Rinehart. When a Man Marries, volume 1671 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/whamm10.zip.

Plato:1999:G

[1621] Plato. Gorgias, volume 1672 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/grgis10.
zip.

Plato:1999:LH

[1622] Plato. Lesser Hippias, volume 1673 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/lhpps10.
zip.

Anonymous:1999:NST

[1623] Anonymous. The Narrative of Sojourner Truth, volume 1674 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/sjrnr10.zip.

Brown:1999:NFO

[1624] Arthur Judson Brown. New Forces in Old China, volume 1675 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ldchn10.zip.

Plato:1999:Ab

[1625] Plato. Alcibiades I, volume 1676 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/1lcbd10.
zip.

PlatonicImitator:1999:AI

[1626] Platonic Imitator. Alcibiades II, volume 1677 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/2lcbd10.zip.

Balzac:1999:HM

[1627] Honoré de Balzac. An Historical Mystery, volume 1678 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/hmyst10.zip.

Todd:1999:HYF

[1628] Burbank L. Todd. Hiram The Young Farmer, volume 1679 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/hrmyf10.zip.

Balzac:1999:ASC

[1629] Honoré de Balzac. At the Sign of the Cat and Racket, volume 1680 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/ctrkt10.zip.

PlatonicImitator:1999:E

[1630] Platonic Imitator. Eryxias, volume 1681 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/ryxis10.zip.

Plato:1999:Mb

[1631] Plato. Menexenus, volume 1682 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/mnxns10.
zip.

Balzac:1999:H

[1632] Honoré de Balzac. Honorine, volume 1683 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/hnrne10.zip.

Meredith:1999:E

[1633] George Meredith. The Egoist, volume 1684 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/egost10.zip.

Leroux:1999:MYR

[1634] Gaston Leroux. Mystery of the Yellow Room, volume 1685 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ylorm10.zip.

Leroux:1999:SN

[1635] Gaston Leroux. The Secret of the Night, volume 1686 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/tsotn10.zip.

Plato:1999:Pd

[1636] Plato. Parmenides, volume 1687 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/prmds10.
zip.

London:1999:PA

[1637] Jack London. The People of the Abyss, volume 1688 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/tpota10.zip.

Sanger:1999:PC

[1638] Margaret Sanger. The Pivot of Civilization, volume 1689 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/pvcvl10.zip.

Haggard:1999:M

[1639] H. Rider Haggard. Marie, volume 1690 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/marie10.zip.

King:1999:DML

[1640] Martin Luther King, Jr. I Have A Dream, Martin Luther King, Jr., volume 1691 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/dream10.zip.

Johnston:1999:

[1641] Mary Johnston. 1492, volume 1692 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/c149210.
zip.

Rinehart:1999:DD

[1642] Mary Roberts Rinehart. Dangerous Days, volume 1693 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ddays10.zip.

Reilly:1999:OLH

[1643] S. A. Reilly. Our Legal Heritage, volume 1694 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/rlglh10.zip.

Chesterton:1999:MWW

[1644] G. K. (Gilbert Keith) Chesterton. The Man Who Was Thursday, volume 1695 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext99/tmwht10.zip.

Chesterton:1999:CQT

[1645] G. K. (Gilbert Keith) Chesterton. The Club of Queer Trades, volume 1696 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext99/tcoqt10.zip.

Kingsley:1999:MHL

[1646] Charles Kingsley. Madam How and Lady Why, volume 1697 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/hwwhy10.zip.

Verne:1999:SCb

[1647] Jules Verne. The Survivors of the Chancellor, volume 1698 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/tsotc10.zip.

Oppenheim:1999:VM

[1648] E. Phillips (Edward Phillips) Oppenheim. The Vanished Messenger, volume 1699 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/vmsgr10.zip.

Gaskell:1999:LCBa

[1649] Elizabeth Cleghorn Gaskell. Life of Charlotte Brontë, V2, volume 1700 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/2locb10.zip.

Wiggin:1999:SWB

[1650] Kate Douglas Smith Wiggin. Story Of Waitstill Baxter, volume 1701 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/tsowb10.zip.

Iles:1999:CAA

[1651] George Iles. 19th Century Actor Autobiographies, volume 1702 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/aauto10.zip.

Hornung:1999:DMT

[1652] E. W. (Ernest William) Hornung. Dead Men Tell No Tales, volume 1703 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/dmtnt10.zip.

Balzac:1999:P

[1653] Honoré de Balzac. Pierrette, volume 1704 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/prrtt10.zip.

Williams:1999:HSV

[1654] Henry Smith Williams. A History of Science, V 1, volume 1705– 1708 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext99/1hsci10.zip   (V.1); ftp://
uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/2hsci10.
zip   (V.2); ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/3hsci10.zip   (V.3); ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/4hsci10.zip   (V.4).

Gissing:1999:NGS

[1655] George Gissing. New Grub Street, volume 1709 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/nwgrb10.zip.

Balzac:1999:GB

[1656] Honoré de Balzac. La Grande Breteche, volume 1710 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/brtch10.zip.

Haggard:1999:CS

[1657] H. Rider Haggard. Child of Storm, volume 1711 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/cstrm10.zip.

Conrad:1999:R

[1658] Joseph Conrad. The Rescue, volume 1712 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999.

Stephenson:1999:LPL

[1659] Nathaniel W. (Nathaniel Wright) Stephenson. Lincoln’s Personal Life, volume 1713 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/lsplf10.zip.

Balzac:1999:ASW

[1660] Honoré de Balzac. Another Study of Woman, volume 1714 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/nswmn10.zip.

Balzac:1999:EG

[1661] Honoré de Balzac. Eugenie Grandet, volume 1715 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/gngnd10.zip.

Freeman:1999:COS

[1662] Mary Eleanor Wilkins Freeman. Copy-Cat and Other Stories, volume 1716 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext99/cpyct10.zip.

Chesterton:1999:WWW

[1663] G. K. (Gilbert Keith) Chesterton. What’s Wrong With The World, volume 1717 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/wwwtw10.zip.

Chesterton:1999:M

[1664] G. K. (Gilbert Keith) Chesterton. Manalive, volume 1718 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/mnalv10.zip.

Chesterton:1999:BWH

[1665] G. K. (Gilbert Keith) Chesterton. The Ballad of the White Horse, volume 1719 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999.

Chesterton:1999:MWKb

[1666] G. K. (Gilbert Keith) Chesterton. The Man Who Knew Too Much, volume 1720 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/mwktm10a.zip.

Chesterton:1999:TPb

[1667] G. K. (Gilbert Keith) Chesterton. The Trees of Pride, volume 1721 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/trprd10.zip.

Luther:1999:MLL

[1668] Martin Luther. Martin Luther’s Large Catechism, Bente and Dau, Trns, volume 1722 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/lrgct10.zip.

Bower:1999:C

[1669] B. M. Bower. Cow-Country, volume 1723 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/cwcnt10.zip.

Haggard:1999:F

[1670] H. Rider Haggard. Finished, volume 1724 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/fnshd10.zip.

Henry:1999:HW

[1671] O. Henry. Heart of the West, volume 1725 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/hrtws10.zip.

Plato:1999:T

[1672] Plato. Theaetetus, volume 1726 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/thtus10.
zip.

Homer:1999:Ob

[1673] Homer. The Odyssey, volume 1727 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/dyssy10.
zip.

Homer:1999:OH

[1674] Homer. The Odyssey of Homer, volume 1728 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/dyssy08a.zip.

Balzac:1999:DW

[1675] Honoré de Balzac. The Deserted Woman, volume 1729 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/dswmn10.zip.

London:1999:MBJ

[1676] Jack London. Michael, Brother of Jerry, volume 1730 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/mcjer10.zip.

Thompson:1999:SS

[1677] Francis Thompson. Sister Songs, volume 1731 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/ssngs10.zip.

Chekhov:1999:SOS

[1678] Anton Pavlovich Chekhov. The Schoolmistress and other stories, volume 1732 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext99/tschm10.zip.  The

Schoolmistress–A Nervous Breakdown–Misery–Champagne –After the Theatre–A Lady’s Story–In Exile–The Cattle-Dealers –Sorrow–On Official Duty–The First-class Passenger –A Tragic Actor–A Transgression– Small Fry–The Requiem –In the Coach-house–Panic Fears–The Bet –The Head-Gardener’s Story–The Beauties –The Shoemaker and the Devil.

Davis:1999:RCG

[1679] Richard Harding Davis. The Red Cross Girl, volume 1733 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/rdcrg10.zip.

Wells:1999:SPH

[1680] H. G. (Herbert George) Wells. Secret Places of the Heart, volume 1734 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/spoth10.zip.

Plato:1999:Sb

[1681] Plato. Sophist, volume 1735 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/sopht10.
zip.

Anonymous:1999:CSA

[1682] Anonymous. Cromwell, Shakespeare Apocrypha, volume 1736 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws4911.zip.

Balzac:1999:FC

[1683] Honoré de Balzac. Facino Cane, volume 1737 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/fcane10.zip.

Plato:1999:Sc

[1684] Plato. Statesman, volume 1738 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/sttsm10.
zip.

Hecker:1999:BDD

[1685] J. F. C. (Justus Friedrich Carl) Hecker. The Black Death, and The Dancing Mania translated by B. G. Babington, volume 1739 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/bdadm10.zip.

Bower:1999:FUL

[1686] B. M. Bower. The Flying U’s Last Stand, volume 1740 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/fusls10.zip.

Packard:1999:WM

[1687] Frank L. (Frank Lucius) Packard. The White Moll, volume 1741 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/wtmol10.zip.

Davis:1999:MC

[1688] Richard Harding Davis. Miss Civilization, volume 1742 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/miscv10.zip.

Wells:1999:TSD

[1689] H. G. (Herbert George) Wells. Twelve Stories and a Dream, volume 1743 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/12sad10.zip.

Plato:1999:Pe

[1690] Plato. Philebus, volume 1744 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/philb10.
zip.

Milton:1999:PW

[1691] John Milton. Poetical Works, volume 1745 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/pmsjm10.zip.

Lang:1999:NCR

[1692] Andrew Lang. New Collected Rhymes, volume 1746 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/nwclr10.zip.

Lincoln:1999:RS

[1693] Natalie Sumner Lincoln. The Red Seal, volume 1747 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/redsl10.zip.

Gaboriau:1999:OPM

[1694] Emile Gaboriau. Other People’s Money, volume 1748 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/opmny10.zip.

Balzac:1999:CB

[1695] Honoré de Balzac. Cousin Betty, volume 1749 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/cbtty10.zip.

Plato:1999:L

[1696] Plato. Laws, volume 1750 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/plaws10.
zip.

Pyle:1999:TL

[1697] Howard Pyle. Twilight Land, volume 1751 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/

etext99/twlnd10.zip. Introduction– The Stool of Fortune– The Talisman of Solomon– Ill-Luck and the Fiddler– Empty Bottles– Good Gifts and a Fool’s Folly– The Good of a Few Words– Woman’s Wit– A Piece of Good Luck– The Fruit of Happiness– Not a Pin to Choose– Much Shall Have More and Little Shall Have Less– Wisdom’s Wages and Folly’s Pay– The Enchanted Island– All Things are as Fate Wills– Where to Lay the Blame– The Salt of Life.

Orczy:1999:D

[1698] Emmuska Orczy Orczy, Baroness. El Dorado, volume 1752 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ldrdo10.zip.

Chekhov:1999:SSI

[1699] Anton Pavlovich Chekhov. Swan Song and Intro, volume 1753 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/swnsg10.zip.

Chekhov:1999:S

[1700] Anton Pavlovich Chekhov. The Sea-Gull, volume 1754 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/cgull10.zip.

Chekhov:1999:I

[1701] Anton Pavlovich Chekhov. Ivanoff, volume 1755 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/vanof10.zip.

Chekhov:1999:UV

[1702] Anton Pavlovich Chekhov. Uncle Vanya, volume 1756 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/vanya10.zip.

Aldrich:1999:CD

[1703] Thomas Bailey Aldrich. The Cruise of the Dolphin, volume 1757 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/dlphn10.zip.

Aldrich:1999:MD

[1704] Thomas Bailey Aldrich. Majorie Daw, volume 1758 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/mjdaw10.zip.

Keith:1999:BB

[1705] Marian Keith. The Black-Bearded Barbarian, volume 1759 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/bbbrb10.zip.

Davis:1999:MWC

[1706] Richard Harding Davis. The Man Who Could Not Lose, volume 1760 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/mwcnl10.zip.

Davis:1999:MBT

[1707] Richard Harding Davis. My Buried Treasure, volume 1761 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/mbtrs10.zip.

Davis:1999:C

[1708] Richard Harding Davis. The Consul, volume 1762 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/tcnsl10.zip.

Davis:1999:NF

[1709] Richard Harding Davis. The Nature Faker, volume 1763 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ntrfk10.zip.

Davis:1999:BBS

[1710] Richard Harding Davis. Billy and the Big Stick, volume 1764 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/bbstk10.zip.

Shakespeare:1999:HVP

[1711] William Shakespeare. Henry VI Part 1, volume 1765 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws0111.zip.

Shakespeare:1999:KRIa

[1712] William Shakespeare. King Richard III, volume 1768 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws0411.zip.

Shakespeare:1999:CE

[1713] William Shakespeare. The Comedy of Errors, volume 1769 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws0611.zip.

Anonymous:1999:KEI

[1714] Anonymous. King Edward III, Shakespeare Apocrypha, volume 1770 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/1ws5010.zip.

Shakespeare:1999:TAa

[1715] William Shakespeare. Titus Andronicus, volume 1771 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws0911.zip.

Shakespeare:1999:TS

[1716] William Shakespeare. The Taming of the Shrew, volume 1772 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws1011.zip.

Shakespeare:1999:TGV

[1717] William Shakespeare. Two Gentlemen of Verona, volume 1773 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws1111.zip.

Shakespeare:1999:LLL

[1718] William Shakespeare. Love’s Labour’s Lost, volume 1774 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws1211.zip.

Shakespeare:1999:KJ

[1719] William Shakespeare. King John, volume 1775 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws1411.zip.

Shakespeare:1999:KRIb

[1720] William Shakespeare. King Richard II, volume 1776 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws1511.zip.

Shakespeare:1999:RJ

[1721] William Shakespeare. Romeo and Juliet, volume 1777 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws1611.zip.

Skakespeare:1999:MND

[1722] William Skakespeare. A Midsummer Night’s Dream, volume 1778 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/1ws1711.zip.

Shakespeare:1999:MV

[1723] William Shakespeare. The Merchant of Venice, volume 1779 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws1811.zip.

Shakespeare:1999:KHIa

[1724] William Shakespeare. King Henry IV, Part 1, volume 1780 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws1911.zip.

Shakespeare:1999:MWW

[1725] William Shakespeare. The Merry Wives of Windsor, volume 1781 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/1ws2011.zip.

Shakespeare:1999:KHIb

[1726] William Shakespeare. King Henry IV, Part 2, volume 1782 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws2111.zip.

Shakespeare:1999:MAA

[1727] William Shakespeare. Much Ado About Nothing, volume 1783 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws2211.zip.

Shakespeare:1999:KHVa

[1728] William Shakespeare. King Henry V, volume 1784 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws2311.zip.

Shakespeare:1999:JC

[1729] William Shakespeare. Julius Caesar, volume 1785 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws2411.zip.

Shakespeare:1999:YLI

[1730] William Shakespeare. As You Like It, volume 1786 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws2511.zip.

Shakespeare:1999:H

[1731] William Shakespeare. Hamlet, volume 1787 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws2611.zip.

Shakespeare:1999:TC

[1732] William Shakespeare. Troilus and Cressida, volume 1790 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws2911.zip.

Shakespeare:1999:AWE

[1733] William Shakespeare. All’s Well That Ends Well, volume 1791 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws3011.zip.

Shakespeare:1999:MM

[1734] William Shakespeare. Measure for Measure, volume 1792 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws3111.zip.

Shakespeare:1999:O

[1735] William Shakespeare. Othello, volume 1793 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws3211.zip.

Shakespeare:1999:KL

[1736] William Shakespeare. King Lear, volume 1794 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws3311.zip.

Shakespeare:1999:M

[1737] William Shakespeare. Macbeth, volume 1795 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws3411.zip.

Shakespeare:1999:AC

[1738] William Shakespeare. Antony and Cleopatra, volume 1796 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws3511.zip.

Shakespeare:1999:Ca

[1739] William Shakespeare. Coriolanus, volume 1797 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws3611.zip.

Shakespeare:1999:TAb

[1740] William Shakespeare. Timon of Athens, volume 1798 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws3711.zip.

Shakespeare:1999:Cb

[1741] William Shakespeare. Cymbeline, volume 1799 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws3911.zip.

Shakespeare:1999:WT

[1742] William Shakespeare. The Winter’s Tale, volume 1800 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws4011.zip.

Shakespeare:1999:T

[1743] William Shakespeare. The Tempest, volume 1801 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/1ws4111.zip.

Shakespeare:1999:KHVb

[1744] William Shakespeare. King Henry VIII, volume 1802 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1ws4211.zip.

Raine:1999:WSO

[1745] William MacLeod Raine. Wyoming, A Story of the Outdoor West, volume 1803 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/wymng10.zip.

Wells:1999:WFI

[1746] H. G. (Herbert George) Wells. War and the Future: Italy, France and Britain at War, volume 1804 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/wrftr10.
zip.

Henry:1999:GG

[1747] O. Henry. The Gentle Grafter, volume 1805 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/

etext99/grftr10.zip. I. The Octopus Marooned – II. Jeff Peters as a Personal Magnet – III. Modern Rural Sports – IV. The Chair of Philanthromathematics – V. The Hand That Riles the World – VI. The Exact Science of Matrimony – VII. A Midsummer Masquerade – VIII. Shearing the Wolf – IX. Innocents of Broadway – X. Conscience in Art – XI. The Man Higher Up – XII. A Tempered Wind – XIII. Hostages to Momus – XIV. The Ethics of Pig.

Davis:1999:F

[1748] Richard Harding Davis. The Frame Up, volume 1806 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/frmup10.zip.

Davis:1999:LH

[1749] Richard Harding Davis. The Lost House, volume 1807 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/lsths10.zip.

Davis:1999:LJP

[1750] Richard Harding Davis. The Log of the Jolly Polly, volume 1808 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/jlply10.zip.

Raine:1999:BOT

[1751] William MacLeod Raine. Bucky O’Connor: A Tale of the Unfenced Border, volume 1809 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext99/bkcnr10.zip.

Balzac:1999:SH

[1752] Honoré de Balzac. A Second Home, volume 1810 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/2ndhm10.zip.

Balzac:1999:MD

[1753] Honoré de Balzac. Massimilla Doni, volume 1811 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/msmdn10.zip.

Balzac:1999:PB

[1754] Honoré de Balzac. A Prince of Bohemia, volume 1812 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/prbhm10.zip.

Balzac:1999:MB

[1755] Honoré de Balzac. A Man of Business, volume 1813 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/mnbus10.zip.

Biggers:1999:AC

[1756] Earl Derr Biggers. The Agony Column, volume 1814 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/gnycl10.zip.

Nicolay:1999:BLA

[1757] Helen Nicolay. The Boys’ Life of Abraham Lincoln, by Helen Nicolay, volume 1815 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/bloal10.zip.

Ogden:1999:T

[1758] Ruth Ogden. Tattine, volume 1816 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp:/

/uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/tttne10.
zip.

Davis:1999:QL

[1759] Richard Harding Davis. A Question of Latitude, volume 1817 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/qlttd10.zip.

Davis:1999:S

[1760] Richard Harding Davis. The Spy, volume 1818 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/thspy10.zip.

Davis:1999:M

[1761] Richard Harding Davis. The Messengers, volume 1819 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/msgrs10.zip.

Davis:1999:WD

[1762] Richard Harding Davis. A Wasted Day, volume 1820 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/wstdy10.zip.

Davis:1999:CL

[1763] Richard Harding Davis. A Charmed Life, volume 1821 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/chmlf10.zip.

Davis:1999:A

[1764] Richard Harding Davis. The Amateur, volume 1822 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/thmtr10.zip.

Davis:1999:MM

[1765] Richard Harding Davis. The Make-Believe Man, volume 1823 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/mbman10.zip.

Davis:1999:PM

[1766] Richard Harding Davis. Peace Manoeuvres, volume 1824 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/pcmnv10.zip.

Burgess:1999:ARF

[1767] Thornton W. Burgess. Adventures of Reddy Fox, volume 1825 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/rdyfx10.zip.

Balzac:1999:S

[1768] Honoré de Balzac. Sarrasine, volume 1826 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/srrsn10.zip.

Gaskell:1999:LCBb

[1769] Elizabeth Cleghorn Gaskell. Life of Charlotte Brontë, V1, volume 1827 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/1locb10.zip.

Scott:1999:CC

[1770] Sir Walter Scott. Chronicles of the Canongate, volume 1828 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/cnngt10.zip.

Mason:1999:MM

[1771] Mary Murdoch Mason. Mae Madden, volume 1829 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/mmmmm10.zip.

Aldrich:1999:WT

[1772] Thomas Bailey Aldrich. Wyndham Towers, volume 1830 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/wndhm10.zip.

Hawthorne:1999:LKL

[1773] Julian Hawthorne. The Lock and Key Library: —bClassic Mystery and Detective Stories - Old Time English, volume 1831 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/lckyl10.zip.

Colbron:1999:CLW

[1774] Grace Isabel Colbron. The Case of The Lamp That Went Out, volume 1832 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext99/tltwo10.zip.

Colbron:1999:CRL

[1775] Grace Isabel Colbron. The Case of The Registered Letter, volume 1833 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/rgstl10.zip.

Colbron:1999:CPD

[1776] Grace Isabel Colbron. The Case of The Pocket Diary Found in the Snow, volume 1834 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/pdfis10.zip.

Colbron:1999:CPB

[1777] Grace Isabel Colbron. The Case of the Pool of Blood in the Pastor’s Study, volume 1835 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.

cso.uiuc.edu/pub/etext/gutenberg/etext99/pbips10.zip.

Colbron:1999:CGB

[1778] Grace Isabel Colbron. The Case of the Golden Bullet, volume 1836 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/cotgb10.zip.

Twain:1999:PP

[1779] Mark Twain. The Prince and the Pauper, volume 1837 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/prppr10.zip.

Lawson:1999:NVC

[1780] John Lawson. A New Voyage to Carolina, volume 1838 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/nvycr10.zip.

Wolf:1999:OTB

[1781] Emma Wolf. Other Things Being Equal, volume 1839 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/otbeq10.zip.

Dreiser:1999:F

[1782] Theodore Dreiser. The Financier, volume 1840 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/tfncr10.zip.

Balzac:1999:ZM

[1783] Honoré de Balzac. Z. Marcas, volume 1841 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/zmrcs10.zip.

Verne:1999:MS

[1784] Jules Verne. Michael Strogoff, volume 1842 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/strgf11.zip.

Davis:1999:VM

[1785] Richard Harding Davis. Vera, The Medium, volume 1843 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/veram10.zip.

Ascham:1999:Sb

[1786] Roger Ascham. The Scholemaster, volume 1844 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/smstr10.zip.

Beerbohm:1999:ZD

[1787] Sir Max Beerbohm. Zuleika Dobson, volume 1845 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/zdbsn10.zip.

Raine:1999:VS

[1788] William MacLeod Raine. The Vision Splendid, volume 1846 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/vspld10.zip.

McNeill:1999:SMS

[1789] John Charles McNeill. Songs, Merry and Sad, volume 1847 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/sngms10.zip.

Haggard:1999:MD

[1790] H. Rider (Henry Rider) Haggard. Montezuma’s Daughter, volume 1848 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/mzdtr10.zip.

Oppenheim:1999:YC

[1791] E. Phillips (Edward Phillips) Oppenheim. The Yellow Crayon, volume 1849 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext99/ycryn10.zip.

Irving:1999:OC

[1792] Washington Irving. Old Christmas, volume 1850 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/oxmas10.zip.

Green:1999:WA

[1793] Anna Katharine Green. The Woman in the Alcove, volume 1851 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/winta10.zip.

Meredith:1999:L

[1794] Owen Meredith. Lucile, volume 1852 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL

ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/etext99/lucil10.
zip.

Beck:1999:NVO

[1795] L. Adams (Lily Adams) Beck. The Ninth Vibration, and other stories, volume 1853 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/9thvb10.zip.

Balzac:1999:CM

[1796] Honoré de Balzac. Catherine de Medici, volume 1854 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ctdmd10.zip.

Lang:1999:BAB

[1797] Andrew Lang. Ban and Arriere Ban - A Rally of Fugitive Rhymes, volume 1855 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.

uiuc.edu/pub/etext/gutenberg/etext99/bnabn10.zip.

Balzac:1999:CP

[1798] Honoré de Balzac. Cousin Pons, volume 1856 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/cspns10.zip.

Green:1999:IO

[1799] Anna Katharine Green. Initials Only, volume 1857 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/ionly10.zip.

Kipling:1999:PTH

[1800] Rudyard Kipling. Plain Tales from the Hills, volume 1858 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/ptfth10.zip.

Beerbohm:1999:WMB

[1801] Sir Max Beerbohm. The Works of Max Beerbohm, volume 1859 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/twomb10.zip.

Kingsley:1999:W

[1802] Charles Kingsley. Westward Ho!, volume 1860 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/wstho10.zip.

Aldrich:1999:OTS

[1803] Thomas Bailey Aldrich. An Old Town By The Sea, volume 1861 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/ldtwn10.zip.

Daudet:1999:TT

[1804] Alphonse Daudet. Tartarin of Tarascon, volume 1862 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/trtrn10.zip.

Thackeray:1999:CGC

[1805] William Makepeace Thackeray. From Cornhill to Grand Cairo, volume 1863 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.

edu/pub/etext/gutenberg/etext99/crhcr10.zip.

Lodge:1999:HTA

[1806] Henry Cabot Lodge. Hero Tales From American History, volume 1864 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/htfah10.zip.

Trollope:1999:NA

[1807] Anthony Trollope. North America, volume 1865–1866 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/1noam10.zip   (Vol.1); ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/etext99/2noam10.zip   (Vol.2).

Wiggin:1999:DGG

[1808] Kate Douglas Smith Wiggin. The Diary of a Goose Girl, volume 1867 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign,

IL 61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/gsgrl10.zip.

Wiggin:1999:PP

[1809] Kate Douglas Smith Wiggin. Penelope’s Postscripts, volume 1868 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL

61825-2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/
etext/gutenberg/etext99/pnlps10.zip.

Rinehart:1999:MLT

[1810] Mary Roberts Rinehart. The Man in Lower Ten, volume 1869 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/tmilt10.zip.

Saki:1999:RRE

[1811] Saki. Reginald in Russia, etc., volume 1870 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/rgrus10.zip.

Balzac:1999:DA

[1812] Honoré de Balzac. The Deputy of Arcis, volume 1871 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/arcis10.zip.

Milne:1999:RHM

[1813] A. A. Milne. The Red House Mystery, volume 1872 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/rdhsm10.zip.

Balzac:1999:G

[1814] Honoré de Balzac. Gambara, volume 1873 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/gmbra10.zip.

Nesbit:1999:RC

[1815] E. (Edith) Nesbit. The Railway Children, volume 1874 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/rlwyc10.zip.

Peattie:1999:PW

[1816] Elia W. Peattie. Painted Windows, volume 1875 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/pwnds10.zip.

Peattie:1999:SF

[1817] Elia W. Peattie. The Shape of Fear, volume 1876 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/tshfr10.zip.

Peattie:1999:MW

[1818] Elia W. Peattie. A Mountain Woman, volume 1877 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/mtwmn10.zip.

Molloy:1999:RR

[1819] J. Fitzgerald Molloy. Royalty Restored, volume 1879 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA,

1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
etext99/rruc210.zip.

Cooper:1999:P

[1820] James Fenimore Cooper. The Pathfinder, volume 1880 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-

2782, USA, 1999. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/etext99/pthfn10.zip.